### v2

# The CFT of SLE loop measures and the Kontsevich–Suhov conjecture

Abstract.
This paper initiates the study of the conformal field theory of the SLEκ loop measure $\nu$ for $\kappa\in(0,4]$, the range where the loop is almost surely simple. First, we construct two commuting representations $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ of the Virasoro algebra with central charge $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}\leq 1$ as (unbounded) first order differential operators on $L^{2}(\nu)$. Second, we introduce highest-weight representations and characterise their structure: in particular, we prove the existence of vanishing singular vectors at arbitrary levels on the Kac table. Third, we prove an integration by parts formula for the SLE loop measure, and use it to define the Shapovalov form of the representation, a non degenerate (but not positive definite) Hermitian form $\mathcal{Q}$ on $L^{2}(\nu)$ with a remarkably simple geometric expression. The fact that $\mathcal{Q}$ differs from the $L^{2}(\nu)$-inner product is a manifestation of non-unitarity. Finally, we write down a spectral resolution of $\mathcal{Q}$ using the joint diagonalisation of $\mathbf{L}_{0}$ and $\bar{\mathbf{L}}_{0}$. As an application of these results, we provide the first proof of the uniqueness of restriction measures, as conjectured by Kontsevich and Suhov.
Our results lay the groundwork for an in-depth study of the CFT of SLE: in forthcoming works, we will define correlation functions on Riemann surfaces, and prove conformal Ward identities, BPZ equations, and conformal bootstrap formulas.
Key words and phrases:
Schramm–Loewner evolutions, Virasoro algebra, conformal field theory.
1991 Mathematics Subject Classification:
Primary: 60J67; 17B68. Secondary: 47L55; 81T40; 30C62.
1. Introduction and main results
1.1. Background and overview
In 2000, Schramm [Sch00] introduced a family of random fractal curves in the plane indexed by a positive real $\kappa$, now known as Schramm–Loewner Evolutions (SLE). These curves are conjectured (and sometimes proved) to describe the scaling limits of interfaces of spin clusters in critical 2D statistical mechanics models. This major breakthrough led to a tremendous increase in the rigorous understanding of the scaling limit of these systems. On the other hand, 2D critical systems have been studied through the lens of conformal field theory (CFT) since the groundbreaking work of Belavin, Polyakov and Zamolodchikov [BPZ84]. The main assumption of this approach is that scaling limits of these models exhibit local conformal invariance, and the algebraic constraints implied by this infinite dimensional symmetry group drastically limit the possible theories existing in the limit. This idea was also present in Schramm’s work: SLE was characterised from the start as the unique family of measures on curves satisfying conformal invariance and a certain domain Markov property. In [LSW03], Lawler, Schramm and Werner studied the “conformal restriction” property of chordal SLE, relating SLE in a domain to SLE in its subdomains. Kontsevich and Suhov then asked whether this property gave another characterisation of the measure; more precisely, they conjectured the existence and uniqueness of a measure on Jordan loops satisfying the restriction property [KS07, Conjecture 1], in the range of parameters corresponding to $\kappa\in(0,4]$. While the existence was settled in previous works [Wer08, BD16, KW16, Zha21], Theorem 1.9 of this article gives a positive answer to the uniqueness part.
In an attempt to reconcile probability and algebra, the CFT aspects of SLE have been studied since the inception of the theory, mostly focusing on the chordal case. The so-called partition functions of SLE can be interpreted as CFT correlation functions [BB03a, BB03b, BBK05, Kyt07, Dub07, Law09, Dub15b, Dub15a, KP16, PW19, FLPW24], see [Pel19] for a thorough review, but the mathematical framework of such a CFT remains only partially defined. As explained in [Pel19], the most unsatisfactory aspect is the absence of mathematically well-defined “fields” whose expectations would correspond to the SLE partition function. For instance, while SLE partition functions satisfy null-vector/BPZ equations, it is not known whether this corresponds to the insertion of a degenerate field in a correlation function. Our Theorem 1.4 gives a first step toward a positive answer to this question (the second step – the conformal Ward identities – will be the topic of a follow-up work).
In this paper, we consider the loop version of SLE [KW16, Zha21] (see also [Wer08, BD16] for the special cases $\kappa=\frac{8}{3},2$), and introduce the main ingredients that will turn the SLE loop measure into a full-fledged CFT. In a nutshell, we construct a structure $(L^{2}(\nu),(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}},\mathcal{Q})$, where $\nu$ is the SLEκ loop measure ($\kappa\in(0,4]$), $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ are two commuting representations of the Virasoro algebra with central charge $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}\leq 1$ acting as (unbounded) operators on $L^{2}(\nu)$, and $\mathcal{Q}$ is a Hermitian form on $L^{2}(\nu)$ such that $\mathbf{L}_{-n}$ is the $\mathcal{Q}$-adjoint of $\mathbf{L}_{n}$ for all $n\in\mathbb{Z}$. In representation theoretic terms, $\mathcal{Q}$ is called the Shapovalov form. It turns out to have a very natural geometric interpretation, consisting in “exchanging” the interior and the exterior of the curve. Our representations are constructed as follows: the Witt algebra defines invariant vector fields on the space of Jordan curves, and the operators are Lie derivatives in these directions. Representations of this flavour (but with substantial differences described in detail in Section 1.3) have been considered in [KY87, KY88, BB03b, Kyt07], seeking an analogue of the Borel-Weil theorem for orbits of the group of diffeomorphisms of the circle. Much closer to our approach are the representations considered in [CP14, GQW24], inspired by the variational formulas of [DS63].
The Hermitian form $\mathcal{Q}$ and the $\mathcal{Q}$-adjoint relations are obtained via an integration by parts formula, which is an infinitesimal version of the conformal restriction property. Interestingly, the integration by parts is precisely the one (formally) satisfied by the path integral with action functional given by the universal Liouville action (a Kähler potential for the Weil-Petersson metric on the universal Teichmüller space [TT06]). This formula gives additional content to the idea – initially motivated by semi-classical considerations [VW20] (see also [CW23]) – that the SLE loop is the correct interpretation of this path integral. The joint spectral theory of $\mathbf{L}_{0}$ and $\bar{\mathbf{L}}_{0}$ allows us to construct highest-weight representations indexed by a scalar $\lambda$ called the conformal weight. We prove that these representations are irreducible (Theorem 1.4), giving important information when the conformal weight lies in the Kac table: in this case, there exist linear relations between certain states (e.g. Corollary 1.5), a fact that will be used in a later work to prove the celebrated BPZ equations. Finally, we provide a spectral resolution of the Shapovalov form (Theorem 1.7), a crucial ingredient in our proof of the uniqueness part of Kontsevich–Suhov conjecture.
One of the most important aspects of our work is that the Shapovalov form is distinct from the $L^{2}(\nu)$-inner product, which is a symptom of the non-unitarity of the theory. In fact, it is a standard result from the representation theory of the Virasoro algebra that the Shapovalov form is not positive definite for $c_{\mathrm{m}}\leq 1$ (except for a discrete set of values). In physics, non-unitary theories and their relations to loop models have been studied quite intensively over the past few years, mostly by algebraic methods and consistency conditions (crossing symmetry, conformal bootstrap, null-vector equations…) [RS15, PRS16, HJS20, JRS23]. However, contrary to the regime of central charge $\geq 25$, there seems to be a zoo of possible theories when the central charge is less than 1. This underlines the importance of studying concrete representations like we do in this paper. In the probabilistic literature, a closely related model has been studied with a different point of view [AS21, ARS22], namely using a coupling of SLE with Liouville CFT [DKRV16, KRV20, GKRV23] inspired by Sheffield’s quantum zipper [She16] (see also the comprehensive book [BP24] for a detailed and gentle introduction to this difficult topic). In fact, the generating function of the “electrical thickness” computed in [AS21] is the partition function of our Shapovalov form. This type of coupling also appears in theoretical physics: [Ebe23, (2.20)] proposes an inner product on the space of Virasoro conformal blocks at central charge $c_{\mathrm{L}}\geq 25$ by reweighting by the partition function of a CFT with central charge $c_{\mathrm{m}}=26-c_{\mathrm{L}}$. The putative CFT behind this partition function is not clearly identified there, but is conjectured to exist for all values of the central charge $c_{\mathrm{m}}\leq 1$, with a continuous spectrum of representations [Ebe23, Conjecture 1]. These two properties are compatible with our results, and make SLE a natural candidate. The inner product of [Ebe23] can also be interpreted in light of the Polyakov path integral [Pol81], where the matter field is an SLEκ-type measure on pairs of pants decomposition, and the ghosts are the trivial Beltrami differentials. We do not make any precise statement at this stage, but view this as an interesting direction of research.
In the rest of this introduction, we will describe precisely our main results. Our CFT results require a precise setup and notations, introduced in Section 1.2. Section 1.7 on uniqueness of MKS measures can be read independently. Future perspectives on the programme initiated in the current paper are described in Section 8.
Acknowledgments We are particularly grateful to E. Peltola for numerous discussions on SLE and a detailed feedback on this manuscript. We thank E. Peltola and B. Wu for bringing the question of uniqueness of MKS measures to our attention, and W. Werner for pointing out the relevance of our results to the reversibility and the duality of SLE. We also thank V. Vargas for discussions on SLE and the conformal bootstrap. Some of this work has been achieved during the fantastic conference “Probability in CFT” organised by J. Aru, C. Guillarmou and R. Rhodes, at the Bernoulli Center, EPFL, whose hospitality is greatly acknowledged. AJ is supported by Eccellenza grant 194648 of the Swiss National Science Foundation and is a member of NCCR SwissMAP.
During the final stages of preparation of the current article, we learnt that M. Gordina, W. Qian and Y. Wang were independently working on a related article [GQW24]. Specifically, they construct representations of the Witt and Virasoro algebras on $L^{2}(\nu)$ which would roughly correspond to our Theorems 1.1 and 1.2 below. They also derive an integration by parts formula and construct the Shapovalov form, which would correspond to Theorem 1.3 below. Their approach partly differs from ours, and we invite the reader to check it out in details. We thank them for bringing the references [DS63, CP14] to our knowledge.
1.2. Setup
Let $\mathcal{J}_{0,\infty}$ be the space of Jordan curves disconnecting 0 from $\infty$, endowed with the Carathéodory topology. Elements of $\mathcal{J}_{0,\infty}$ are viewed either as the image $\eta(\mathbb{S}^{1})\subset\mathbb{C}\setminus\{0\}$ of a continuous injective function $\eta:\mathbb{S}^{1}\to\mathbb{C}\setminus\{0\}$, or as an equivalence class $[\eta]$ of loops up to orientation preserving parametrisation. To ease notations, we will most of the time simply write $\eta$ instead of $[\eta]$. It will be useful to parametrise the space $\mathcal{J}_{0,\infty}$ using conformal maps as follows. Let $\mathcal{E}$ be the space of conformal maps $f:\mathbb{D}\to\mathbb{C}$ extending continuously and injectively to $\bar{\mathbb{D}}$, and such that $f(0)=0$, $f^{\prime}(0)=1$ (and $f(\bar{\mathbb{D}})\subset\mathbb{C}$). Similarly and denoting $\mathbb{D}^{*}=\widehat{\mathbb{C}}\setminus\overline{\mathbb{D}}$, let $\tilde{\mathcal{E}}$ be the space of conformal maps $g:\mathbb{D}^{*}\to\widehat{\mathbb{C}}\setminus\{0\}$ extending continuously and injectively to $\overline{\mathbb{D}^{*}}$, and such that $g(\infty)=\infty$, $g^{\prime}(\infty)=1$ (and $0\not\in g(\overline{\mathbb{D}^{*}})$). Endow $\mathcal{E}$ and $\tilde{\mathcal{E}}$ with the uniform topology on $\bar{\mathbb{D}}$ and $\overline{\mathbb{D}^{*}}$ respectively. Every curve $\eta$ in $\mathcal{J}_{0,\infty}$ can be uniquely represented by a pair $(c_{\eta},f_{\eta})\in\mathbb{R}\times\mathcal{E}$ (resp. a pair $(\tilde{c}_{\eta},g_{\eta})\in\mathbb{R}\times\tilde{\mathcal{E}}$), where $e^{c_{\eta}}f_{\eta}$ uniformises the interior of $\eta$ (resp. $e^{-\tilde{c}_{\eta}}g_{\eta}$ uniformises the exterior of $\eta$). See Figure 1. This identifies $\mathcal{J}_{0,\infty}$ with $\mathbb{R}\times\mathcal{E}$ and $\mathbb{R}\times\tilde{\mathcal{E}}$. By a theorem of Carathéodory [Car52], these identifications are in fact homeomorphisms.
Every $f\in\mathcal{E}$ and $g\in\tilde{\mathcal{E}}$ have series expansions in the neighbourhoods of 00 and $\infty$ respectively:
(1.1) $f(z)=z\Big{(}1+\sum_{m=1}^{\infty}a_{m}z^{m}\Big{)}\qquad\text{and}\qquad g(z)=z\Big{(}1+\sum_{m=1}^{\infty}b_{m}z^{-m}\Big{)}.$
We will think of $(a_{m})_{m\geq 1}$ as (complex analytic) coordinates on $\mathcal{E}$. That is, we think of $c$, $f$, $(a_{m})$, $\tilde{c}$, $g$, $(b_{m})$ as functions defined on $\mathcal{J}_{0,\infty}$, e.g. $c:\eta\in\mathcal{J}_{0,\infty}\mapsto c_{\eta}\in\mathbb{R}$. When the context is clear, we will make the dependence on $\eta$ implicit.
Figure 1. A parametrised curve $\eta:\mathbb{S}^{1}\to\widehat{\mathbb{C}}$ having 0 in its interior and $\infty$ in its exterior. Its equivalence class $[\eta]$ (modulo parametrisation) is uniquely represented by a pair $(c,f)\in\mathbb{R}\times\mathcal{E}$ such that $e^{c}f$ uniformises the interior (or a pair $(\tilde{c},g)\in\mathbb{R}\times\tilde{\mathcal{E}}$ such that $e^{-\tilde{c}}g$ uniformises the exterior). This defines homeomorphisms $\mathcal{J}_{0,\infty}\simeq\mathbb{R}\times\mathcal{E}\simeq\mathbb{R}\times\tilde{\mathcal{E}}$.
We denote by $\nu$ the SLEκ loop measure of [Zha21] restricted to $\mathcal{J}_{0,\infty}$, which is a scale-invariant Borel measure on $\mathcal{J}_{0,\infty}$ (Proposition B.1). In this whole paper, $\kappa$ is fixed and belongs to $(0,4]$, the range corresponding to simple loops. By scale invariance, the marginal law of $\nu$ on the $c$-coordinate is a multiple of Lebesgue measure on $\mathbb{R}$. The multiplicative constant is irrelevant for the current paper, and it will be convenient to fix it at $1/\sqrt{\pi}$. Disintegrating $\nu$, we then have a factorisation
$\nu=\frac{1}{\sqrt{\pi}}\mathrm{d}c\otimes\nu^{\#},$
where $\mathrm{d}c$ is Lebesgue measure on $\mathbb{R}$ and $\nu^{\#}$ is a probability measure on $\mathcal{E}$ called the shape measure (borrowing the terminology of [AS21]).
1.3. Virasoro representations on $L^{2}(\nu)$
We now present our results concerning Witt and Virasoro representations on the space of functions defined on $\mathcal{J}_{0,\infty}$, or more precisely on $L^{2}(\nu)$. These results go beyond the SLE setting and are of general interest. See below Theorem 1.2 for a comparison with earlier works.
The Witt algebra is the space of Laurent polynomial vector fields $\mathbb{C}(z)\partial_{z}$. It is a Lie algebra with bracket
(1.2) $[v\partial_{z},w\partial_{z}]=(vw^{\prime}-v^{\prime}w)\partial_{z}.$
We will use the notation $\mathrm{v}_{n}=-z^{n+1}\partial_{z}$ for all $n\in\mathbb{Z}$. The commutation relations on these generators are $[\mathrm{v}_{n},\mathrm{v}_{m}]=(n-m)\mathrm{v}_{n+m}$. Let $\mathrm{v}=v\partial_{z}\in\mathbb{C}(z)\partial_{z}$. The vector field $\mathrm{v}$ generates a family of conformal transformations $\phi_{t}$, solution to the differential equation (existing for small complex $t$)
$\partial_{t}\phi_{t}=v\circ\phi_{t},\qquad\phi_{0}=\mathrm{id}_{O_{\eta}},$
away from a small neighbourhood of 0 and $\infty$. Hence, for every $\eta\in\mathcal{J}_{0,\infty}$, we can consider the small motion $\phi_{t}\cdot\eta=[\phi_{t}\circ\eta]$. The condition that $v$ is a Laurent polynomial can be weakened to any holomorphic $v$ defined in a neighbourhood of the trace of the curve. In (2.5) and (2.6), we will introduce two dense subspaces $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$ of $L^{2}(\nu)$. In Proposition 2.6, we will show that for any $F\in\mathcal{C}$, there exist $\mathcal{L}_{\mathrm{v}}F,\bar{\mathcal{L}}_{\mathrm{v}}F\in\mathcal{C}$ such that for all $\eta\in\mathcal{J}_{0,\infty}$,
$F(\phi_{t}\cdot\eta)=F(\eta)+t\mathcal{L}_{\mathrm{v}}F(\eta)+\bar{t}\bar{\mathcal{L}}_{\mathrm{v}}F(\eta)+o(t),\qquad\text{as}\quad t\to 0.$
In other words, $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$ are the $(1,0)$- and $(0,1)$-parts of the Lie derivative in direction $\mathrm{v}$, and they are densely defined operators with domain $\mathcal{C}$.
Our first main result concerns the study of these differential operators as operators on $L^{2}(\nu)$.
Theorem 1.1 (Witt representation).
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, the densely defined operators $\mathcal{L}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{v}}:\mathcal{C}\to L^{2}(\nu)$ are closable. On their closed extension, they form two commuting representations of the Witt algebra: for all $\mathrm{v},\mathrm{w}\in\mathbb{C}(z)\partial_{z}$, we have
$[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]=\mathcal{L}_{[\mathrm{v},\mathrm{w}]};\qquad[\bar{\mathcal{L}}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]=\bar{\mathcal{L}}_{[\mathrm{v},\mathrm{w}]};\qquad[\mathcal{L}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]=0.$
In the sequel, we will write $\mathcal{L}_{n}:=\mathcal{L}_{\mathrm{v}_{n}}$, ¯$\bar{\mathcal{L}}_{n}:=\bar{\mathcal{L}}_{\mathrm{v}_{n}}$, $n\in\mathbb{Z}$.
In our next result, we modify the Witt representation to a Virasoro representation with central charge $c_{\mathrm{m}}:=1-6\left(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2}\right)^{2}$. To do so, let us introduce the “differential forms” $\vartheta,\tilde{\vartheta}:\mathcal{J}_{0,\infty}\times\mathbb{C}(z)\partial_{z}\to\mathbb{C}$ defined for $\mathrm{v}=v(z)\partial_{z}\in\mathbb{C}(z)\partial_{z}$, $(c,f)\in\mathbb{R}\times\mathcal{E}$ and $(\tilde{c},g)\in\mathbb{R}\times\tilde{\mathcal{E}}$ by:
(1.3) $\vartheta_{e^{c}f}(\mathrm{v})=\frac{e^{-c}}{2i\pi}\oint\mathcal{S}f^{-1}(z)v(e^{c}z)\mathrm{d}z;\qquad\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})=\frac{e^{\tilde{c}}}{2i\pi}\oint\mathcal{S}g^{-1}(z)v(e^{-\tilde{c}}z)\mathrm{d}z,$
where $\mathcal{S}f=(\frac{f^{\prime\prime}}{f^{\prime}})^{\prime}-\frac{1}{2}(\frac{f^{\prime\prime}}{f^{\prime}})^{2}$ is the Schwarzian derivative. The function $\mathcal{S}f^{-1}$ is viewed as a quadratic differential in the neighbourhood of 0, which is dual to meromorphic vector fields by the residue pairing (contour integral on a small loop around 0). For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, we define
(1.4) $\mathbf{L}_{\mathrm{v}}:=\mathcal{L}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{v})\qquad\text{and}\qquad\bar{\mathbf{L}}_{\mathrm{v}}:=\bar{\mathcal{L}}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\overline{\vartheta(\mathrm{v})}.$
Note that $\mathbf{L}_{\mathrm{v}}=\mathcal{L}_{\mathrm{v}}$ if $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$ has no pole at $z=0$. In the sequel, we will use the notation $\mathbf{L}_{n}=\mathbf{L}_{\mathrm{v}_{n}}$ for all $n\in\mathbb{Z}$.
This definition of $\mathbf{L}_{\mathrm{v}}$ is natural from the point of view of geometric quantisation, and the fact that the Schwarzian derivative is the differential of the Weil–Petersson potential on the universal Teichmüller space [TT06, Chapter 2, Section 3.2]. See the heuristics below Theorem 1.2 for more explanations.
Theorem 1.2 (Virasoro representation).
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, the densely defined operators $\mathbf{L}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{v}}:\mathcal{C}_{\mathrm{comp}}\to L^{2}(\nu)$ are closable. On their closed extension, they form two commuting representations of the Virasoro algebra with central charge $c_{\mathrm{m}}$: for all $\mathrm{v},\mathrm{w}\in\mathbb{C}(z)\partial_{z}$,
$[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=\mathbf{L}_{[\mathrm{v},\mathrm{w}]}+\frac{c_{\mathrm{m}}}{12}\omega(\mathrm{v},\mathrm{w});\qquad[\bar{\mathbf{L}}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]=\bar{\mathbf{L}}_{[\mathrm{v},\mathrm{w}]}+\frac{c_{\mathrm{m}}}{12}\overline{\omega(\mathrm{v},\mathrm{w})};\qquad[\mathbf{L}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]=0,$
where $\omega$ is the Virasoro cocycle
(1.5) $\omega(v\partial_{z},w\partial_{z})=\frac{1}{2i\pi}\oint_{\mathbb{S}^{1}}v^{\prime\prime\prime}(z)w(z)\mathrm{d}z.$
The articles [KY87, KY88, Kir98, AM01] study representations of the Virasoro algebra on the space of functions defined on $\mathcal{J}_{0,\infty}$ with a different approach. As already observed in Section 1.2, the space of Jordan curves $\eta$ separating 00 from $\infty$ and such that $c_{\eta}=0$ can be identified with the space $\mathcal{E}$ of univalent functions $f$ of the unit disc with $f(0)=0$ and $f^{\prime}(0)=1$. It can also be identified with the group $Diff_{+}(\mathbb{S}^{1})/\mathbb{S}^{1}$ of $C^{\infty}$ orientation preserving diffeomorphisms of the unit circle $\mathbb{S}^{1}$, via the welding homeomorphism $f_{\eta}^{-1}\circ(e^{-\tilde{c}_{\eta}}g_{\eta}):\mathbb{S}^{1}\to\mathbb{S}^{1}$ (the quotient by $\mathbb{S}^{1}$ corresponding to a free rotation parameter). The Witt and Virasoro operators considered in these papers are then defined via the action of real vector fields (vector fields of $\mathbb{S}^{1}$ which is the Lie algebra of $Diff_{+}(\mathbb{S}^{1})/\mathbb{S}^{1}$) on the space $\mathcal{E}$ by left composition. As we described in Theorems 1.1 and 1.2, our setup relies instead on the action of complex vector fields on $\mathcal{J}_{0,\infty}$ via the associated flows. In particular, we obtain two commuting families of operators instead of only one. Identifying the right framework was one of the main challenges of this work which was inspired by recent progress on Liouville CFT; see especially [BGK+24].
1.4. Integration by parts and Shapovalov form
Our next result concerns the representation of the Shapovalov form on $L^{2}(\nu)$. Namely, we construct a Hermitian form $\mathcal{Q}$ on $L^{2}(\nu)$ such that $\mathcal{Q}(\mathbf{L}_{-n}F,G)=\mathcal{Q}(F,\mathbf{L}_{n}G)$ for all $n\in\mathbb{Z}$.
Let $\iota:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be the anticonformal involution $\iota(z):=1/\bar{z}$. The conjugation by $\iota$ defined by $g\in\tilde{\mathcal{E}}\mapsto\iota\circ g\circ\iota\in\mathcal{E}$ is a bicontinuous involution. We define a bounded operator $\Theta$ on $L^{2}(\nu)$ by
(1.6) $\Theta F(c_{\eta},f_{\eta}):=F(\tilde{c}_{\eta},\iota\circ g_{\eta}\circ\iota),\qquad\forall F\in L^{2}(\nu),\,\forall\eta\in\mathcal{J}_{0,\infty}.$
This defines an involution of $L^{2}(\nu)$. Moreover, by reflection symmetry of $\nu$ ($\nu$ is invariant by conjugation by $\iota$), $\Theta$ is also self-adjoint. Thus, the spectrum of $\Theta$ is $\{1,-1\}$ and we will think of the corresponding eigenspaces as spaces of even and odd functions. Since $\Theta$ is self-adjoint, the expression
(1.7) $\mathcal{Q}(F,G):=\langle F,\Theta G\rangle_{L^{2}(\nu)}$
defines a Hermitian form (i.e. a symmetric sesquilinear form) on $L^{2}(\nu)$. This form is non-degenerate since $\mathcal{Q}(F,\Theta F)=\left\|F\right\|_{L^{2}(\nu)}^{2}\neq 0$ for all $F\neq 0$, but it is neither positive nor negative: it is positive (resp. negative) on the space of even (resp. odd) functions.
Theorem 1.3.
For all $\mathrm{v}=v\partial_{z}\in\mathbb{C}(z)\partial_{z}$, we have the following hermiticity relations with respect to the $L^{2}(\nu)$-inner product:
$\mathbf{L}_{\mathrm{v}}^{*}=\Theta\circ\mathbf{L}_{\mathrm{v}^{*}}\circ\Theta;\qquad\bar{\mathbf{L}}_{\mathrm{v}}^{*}=\Theta\circ\bar{\mathbf{L}}_{\mathrm{v}^{*}}\circ\Theta,$
where $\mathrm{v}^{*}=v^{*}(z)\partial_{z}:=z^{2}\bar{v}(1/z)\partial_{z}$. In other words, $\mathbf{L}_{\mathrm{v}^{*}}$ is the $\mathcal{Q}$-adjoint of $\mathbf{L}_{\mathrm{v}}$: for all $F,G\in\mathcal{C}_{\mathrm{comp}}$, we have
$\mathcal{Q}(\mathbf{L}_{\mathrm{v}^{*}}F,G)=\mathcal{Q}(F,\mathbf{L}_{\mathrm{v}}G);\qquad\mathcal{Q}(\bar{\mathbf{L}}_{\mathrm{v}^{*}}F,G)=\mathcal{Q}(F,\bar{\mathbf{L}}_{\mathrm{v}}G).$
This result stands out as one of our main contributions and drives some of our proofs. Specifically, it will be a crucial input in the proof of the commutation relations of $\mathbf{L}_{\mathrm{v}}$ and $\mathbf{L}_{\mathrm{w}}$ for $\mathrm{v},\mathrm{w}\in z\mathbb{C}[z^{-1}]\partial_{z}$ and in the proof of the closability of $\mathbf{L}_{\mathrm{v}}$ on $L^{2}(\nu)$.
As we will see, Theorem 1.3 is a consequence of the following integration by parts formula of independent interest (see Theorem 4.1): for all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$,
(1.8) $\mathcal{L}_{\mathrm{v}}^{*}=-\bar{\mathcal{L}}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\overline{\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})}+\frac{c_{\mathrm{m}}}{12}\overline{\vartheta_{e^{c}f}(\mathrm{v})};\qquad\bar{\mathcal{L}}_{\mathrm{v}}^{*}=-\mathcal{L}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})+\frac{c_{\mathrm{m}}}{12}\vartheta_{e^{c}f}(\mathrm{v}),$
where $\vartheta_{e^{c}f}(\mathrm{v})$ and $\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})$ are defined in (1.3).
Heuristics. We now discuss some heuristics that first led us to guess the aforementioned integration by parts formula. The SLE loop measure with parameter $\kappa$ can be thought of as being absolutely continuous w.r.t. the SLE loop measure with parameter $8/3$ (corresponding to a vanishing central charge):
(1.9) $\mathrm{d}\nu_{\kappa}(\eta)=\mathrm{exp}\Big{(}\frac{c_{\mathrm{m}}}{24}I^{L}(\eta)\Big{)}\mathrm{d}\nu_{8/3}(\eta)$
where $I_{L}(\eta)$ is the Loewner energy, or universal Liouville action, of $\eta$. This path integral formulation is especially inspired by [CW23], but, as we are about to argue, is also consolidated by our integration by parts formula (1.8). Making sense of (1.9) would require renormalisations because, for $\nu_{8/3}$-almost every $\eta$, $I^{L}(\eta)=+\infty$ (and, indeed, $\nu_{\kappa}$ is not absolutely continuous w.r.t $\nu_{8/3}$ when $\kappa\neq 8/3$). The SLE8/3-loop measure possesses a greater level of symmetry since it is exactly invariant under restrictions. This can be used to show the integration by parts (1.8) in this case. Making formal computations with (1.9) and using a formula by Takhtajan and Teo [TT06, Chapter 2, Theorem 3.5] relating variations of the Loewner energy to the Schwarzian derivative of $f$ and $g$, we can then derive (1.8).
1.5. Highest-weight representations
The spectral theory of $\mathbf{L}_{0}$ and $\bar{\mathbf{L}}_{0}$ leads to the introduction of a family of operators $(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}^{\lambda}_{n})_{n\in\mathbb{Z}}$ indexed by $\lambda\in\mathbb{C}$, which are densely defined operators on $L^{2}(\nu^{\#})$. They have a simple analytic description (3.4) and can be defined more abstractly by the conjugation of the representations $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ with a certain operator introduced in Section 3.1 (see (3.5)). For each $\lambda\in\mathbb{C}$, the commutation relations are preserved by this conjugation (Lemma 3.3), so we get two commuting Virasoro representations $(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})_{n\in\mathbb{Z}}$ with central charge $c_{\mathrm{m}}$. The relevance of these operators (Proposition 3.3) is that
$\mathbf{L}_{0}^{\lambda}\mathds{1}=\bar{\mathbf{L}}_{0}^{\lambda}\mathds{1}=\lambda\mathds{1}\quad\text{and for each}\quad n\geq 1,\quad\mathbf{L}_{n}^{\lambda}\mathds{1}=\bar{\mathbf{L}}_{n}^{\lambda}\mathds{1}=0.$
In other words, the constant function $\mathds{1}$ is a highest-weight state of weight $\lambda$ for these representations. We are going to use this property to define highest-weight representations, but we first need to introduce some notations.
Integer partitions and polynomials. We can represent an integer partition by a sequence $\mathbf{k}=(k_{m})_{m\in\mathbb{N}^{*}}\in\mathbb{N}^{\mathbb{N}^{*}}$ containing only finitely many non-zero terms, where the number partitioned by $\mathbf{k}$ is $|\mathbf{k}|=\sum_{m=1}^{\infty}mk_{m}$. For $N\geq 0$, we denote by $\mathcal{T}$ (resp. $\mathcal{T}_{N}$) the set of all integer partitions (resp. the set of partitions of $N$). By convention, $\mathcal{T}_{0}=\{\varnothing\}$ where $\varnothing=(0,0,0,\dots)$ denotes the empty partition. For every $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$, we define the monomial
(1.10) $\mathbf{a}^{\mathbf{k}}{\bar{\mathbf{a}}}^{\tilde{\mathbf{k}}}:=\prod_{m=1}^{\infty}a_{m}^{k_{m}}{\bar{a}_{m}}^{\tilde{k}_{m}}\in\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}],$
and we call $|\mathbf{k}|$ (resp. $|\tilde{\mathbf{k}}|$) the left (resp. right) degree of $\mathbf{a}^{\mathbf{k}}{\bar{\mathbf{a}}}^{\tilde{\mathbf{k}}}$. There are #$\#\mathcal{T}_{N}\#\mathcal{T}_{\tilde{N}}$ linearly independent monomials of left degree $N$ and right degree $\tilde{N}$. The total degree of $\mathbf{a}^{\mathbf{k}}\bar{\mathbf{a}}^{\tilde{\mathbf{k}}}$ is $|\mathbf{k}|+|\tilde{\mathbf{k}}|$, and the total degree of an arbitrary polynomial is the maximal total degree over its monomials. We will also use the same shorthand notation for operators:
(1.11) $\mathbf{L}_{-\mathbf{k}}:=\cdots\mathbf{L}_{-3}^{k_{3}}\mathbf{L}_{-2}^{k_{2}}\mathbf{L}_{-1}^{k_{1}};\qquad\mathbf{L}_{\mathbf{k}}:=\mathbf{L}_{1}^{k_{1}}\mathbf{L}_{2}^{k_{2}}\mathbf{L}_{3}^{k_{3}}\cdots.$
Note the reversal of the order in which we read the partition.
With these notations at hand, we can define $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}:=\mathbf{L}_{-\mathbf{k}}^{\lambda}\bar{\mathbf{L}}^{\lambda}_{-\tilde{\mathbf{k}}}\mathds{1}\in\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$, $\Psi_{\lambda,\mathbf{k}}:=\Psi_{\lambda,\mathbf{k},\eset}$, and
(1.12) $\mathcal{V}_{\lambda}:=\mathrm{span}\left\{\Psi_{\lambda,\mathbf{k}}\big{|}\,\mathbf{k}\in\mathcal{T}\right\};\qquad\mathcal{W}_{\lambda}:=\mathrm{span}\left\{\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}\big{|}\,\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}\right\},$
where the linear span is understood in the algebraic sense (finite linear combinations). By definition, $\mathcal{V}_{\lambda}$ is a highest-weight representation of weight $\lambda$. As such, they have a natural grading $\mathcal{V}_{\lambda}=\oplus_{N\in\mathbb{N}}\mathcal{V}_{\lambda}^{N}$ where $\mathcal{V}_{\lambda}^{N}=\mathrm{span}\{\Psi_{\lambda,\mathbf{k}}|\,\mathbf{k}\in\mathcal{T}_{N}\}$, and similarly $\mathcal{W}_{\lambda}^{N,\tilde{N}}=\mathrm{span}\{\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}|\,\mathbf{k}\in\mathcal{T}_{N},\,\tilde{\mathbf{k}}\in\mathcal{T}_{\tilde{N}}\}$. Vectors in $\mathcal{W}_{\lambda}^{N,\tilde{N}}$ are eigenstates of $\mathbf{L}_{0}^{\lambda}$ (resp. $\bar{\mathbf{L}}_{0}^{\lambda}$) with eigenvalue $\lambda+N$ (resp. $\lambda+\tilde{N}$).
The Kac table is the countable set (see e.g. [Pel19, Appendix A])
$kac:=\left\{\lambda_{r,s}:=(r^{2}-1)\frac{\kappa}{16}+(s^{2}-1)\frac{1}{\kappa}+\frac{1}{2}(1-rs)|\,r,s\in\mathbb{N}^{*}\right\}.$
In particular, we have
$\lambda_{1,2}=\frac{6-\kappa}{2\kappa};\qquad\lambda_{2,1}=\frac{3\kappa-8}{16}.$
The next theorem classifies the structure of these representations. The main point is that $\mathcal{V}_{\lambda}$ is irreducible for all $\lambda\in\mathbb{C}$.
Theorem 1.4.
  1. (1)
    Suppose $\lambda\not\in kac$. Then, $\mathcal{V}_{\lambda}$ is the Verma module and, for all $N\in\mathbb{N}$, $\mathcal{V}_{\lambda}^{N}=\mathrm{span}\{\mathbf{a}^{\mathbf{k}}|\,\mathbf{k}\in\mathcal{T}_{N}\}$. In particular, $\mathcal{V}_{\lambda}=\mathbb{C}[(a_{m})_{m\geq 1}]$.
  2. (2)
    Suppose $\lambda\in kac$. Then, $\mathcal{V}_{\lambda}$ is the irreducible quotient of the Verma module by the maximal proper submodule.
  3. (3)
    For all $\lambda\in\mathbb{C}$, $\mathcal{W}_{\lambda}$ is isomorphic to the tensor product of two copies of $\mathcal{V}_{\lambda}$. Moreover, for all $\lambda\not\in kac$ and $N\in\mathbb{N}$, we have $\oplus_{n+\tilde{n}\leq N}\mathcal{W}_{\lambda}^{n,\tilde{n}}=\mathrm{span}\{\mathbf{a}^{\mathbf{k}}\bar{\mathbf{a}}^{\tilde{\mathbf{k}}}|\,|\mathbf{k}|+|\tilde{\mathbf{k}}|\leq N\}$. In particular, $\mathcal{W}_{\lambda}=\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$.
Theorem 1.4 implies that certain linear relations exist between the states at level $rs$ in $\mathcal{V}_{\lambda_{r,s}}$. Here is one instance of such a relation when $rs=2$.
Corollary 1.5.
For $\lambda\in\{\frac{6-\kappa}{2\kappa},\frac{3\kappa-8}{16}\}$, the following equality holds in $\mathbb{C}[(a_{m})_{m\geq 1}]$:
(1.13) $\left((\mathbf{L}_{-1}^{\lambda})^{2}-\frac{2}{3}(2\lambda+1)\mathbf{L}_{-2}^{\lambda}\right)\mathds{1}=0.$
Corollary 1.5 is a consequence of Theorem 1.4, but we will also give a “computational proof” of this result in Section 3.3. This result will be combined with conformal Ward identities in a follow-up work in order to prove BPZ equations for correlation functions involving a degenerate field at level 2. We mention that such BPZ equations are well known in the chordal case (see e.g. the review [Pel19]) and are derived using local martingales of SLE. We stress however that our approach is orthogonal to the literature so far since in our context they are consequences of our CFT framework.
Remark 1.6.
Highest-weight representations of the kind of above have also been studied in Liouville CFT [BGK+24, BW23]. In [BW23], the first author and Wu showed a similar result about the structure of these modules on the Kac table. It would be interesting to see if the coupling of Liouville CFT with SLE could give an alternative proof, following the methods of [AS21].
1.6. Spectral resolution
Our final result is a spectral resolution for the Shapovalov form $\mathcal{Q}$: this consists in decomposing $\mathcal{Q}$ on each module $\mathcal{W}_{ip}\simeq\mathcal{V}_{ip}\otimes\overline{\mathcal{V}}_{ip}$, $p\in\mathbb{R}$, forming the joint spectra of the Hamiltonians $\mathbf{L}_{0},\bar{\mathbf{L}}_{0}$. Interestingly, there exists an exact formula for the partition function of the Shapovalov form (i.e. the norm of the constant function in $\mathcal{W}_{ip}$). To state this precisely, define for $f\in\mathcal{E}$,
(1.14) $\tilde{c}_{0}(f)=\log\text{-conformal radius of }f(\mathbb{S}^{1})\text{ viewed from }\infty;$
namely, the exterior of the curve $f(\mathbb{S}^{1})$ is uniformised by $e^{-\tilde{c}_{0}(f)}g$ for some function $g\in\tilde{\mathcal{E}}$ (depending on $f$). This is a geometrically meaningful nonpositive ([Pom75, Equation (21)]) quantity: it is a Kähler potential for the Velling-Kirillov metric on the universal Teichmüller curve [KY87, TT06], and it is sometimes called the logarithmic (or analytic) capacity [KY87]. In [SSW09, AS21] (originally introduced by Kenyon and Wilson), it goes by the name of electrical thickness (see Lemma 5.1). Remarkably, the moment generating function of this random variable under the probability law $\nu^{\#}$ is explicit [AS21, Theorem 1.3]: for $\mathrm{Re}(\lambda)<\frac{1}{2}(1-\frac{\kappa}{8})$, one has [^1] Our $\lambda$ differs from the one in [AS21] by a factor $-2$ since we choose to parametrise by the conformal weight.
(1.15) $\mathrm{R}(\lambda):=\mathbb{E}^{\#}[e^{-2\lambda\tilde{c}_{0}}]=\frac{\sin(\pi(1-\kappa/4))}{\pi(1-\kappa/4)}\frac{\pi\sqrt{(1-\kappa/4)^{2}+\lambda\kappa}}{\sin(\pi\sqrt{(1-\kappa/4)^{2}+\lambda\kappa})}.$
For $\lambda\in[\frac{1}{2}(1-\frac{\kappa}{8}),\infty)$, the generating function is infinite. Note that the square-roots appear as a ratio, so the choice of branch is not important and $\mathrm{R}(\lambda)$ admits a meromorphic continuation to the whole complex plane, whose poles form a discrete subset of the interval $[\frac{1}{2}(1-\frac{\kappa}{8}),\infty)$. The threshold $\frac{1}{2}(1-\frac{\kappa}{8})$ is easy to explain probabilistically: the probability that $e^{-\tilde{c}_{0}}$ gets larger than some $R>0$ is of the same order as the probability that the SLE gets at distance $R^{-1}$ to a fixed point, which is known to scale like $R^{2-(1+\frac{\kappa}{8})}=R^{\frac{\kappa}{8}-1}$ (the Hausdorff dimension of the curve is $1+\frac{\kappa}{8}$ a.s. [Bef08]).
Let $\lambda\in\mathbb{C}$, and $\mathbf{k},\tilde{\mathbf{k}},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}\in\mathcal{T}$. If this is well defined and finite, we set
(1.16) $\mathcal{Q}_{\lambda}(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}^{\prime}}}):=\mathbb{E}^{\#}\left[e^{-(2\lambda+|\mathbf{k}|\wedge|\mathbf{k}^{\prime}|+|\tilde{\mathbf{k}}|\wedge|\tilde{\mathbf{k}}^{\prime}|)\tilde{c}_{0}}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}\overline{\Theta^{\#}\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}}\right],$
where $\Theta^{\#}P(f)=P(\iota\circ g\circ\iota)$ for all polynomials $P$. From (1.15), we have $|\mathcal{Q}_{\lambda}(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})|<\infty$ when $\mathrm{Re}(\lambda)<\frac{1}{2}(1-|\mathbf{k}|\wedge|\mathbf{k}^{\prime}|-|\tilde{\mathbf{k}}|\wedge|\tilde{\mathbf{k}}^{\prime}|-\frac{\kappa}{8})$. For the statement of the next theorem, we use the following convention for the Fourier transform of $F\in L^{2}(\nu)$:
$\widehat{F}(p,f):=\int_{\mathbb{R}}e^{2ipc}F(c,f)\frac{\mathrm{d}c}{\sqrt{\pi}}.$
It defines a unitary transformation of $L^{2}(\nu)$. For $\lambda\in\mathbb{C}$, we will also denote by $\mathrm{B}_{\lambda}$ the Gram matrix of the Shapovalov form of weight $\lambda$: by definition, for all $\mathbf{k},\mathbf{k}^{\prime}\in\mathcal{T}$, $\mathbf{L}_{\mathbf{k}^{\prime}}^{\lambda}\mathbf{L}_{-\mathbf{k}}^{\lambda}\mathds{1}=\mathrm{B}_{\lambda}(\mathbf{k},\mathbf{k}^{\prime})\mathds{1}$. See Appendix A for more details. The next statement uses the operator $\mathcal{P}_{\lambda}$ acting on $\mathcal{W}_{\lambda}$, and studied in detail in Section 3.1. This operator takes values in $\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ and is characterised by $\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$.
Theorem 1.7.
  1. (1)
    For all $\mathrm{Re}(\lambda)<\frac{1}{2}(1-\frac{\kappa}{8})$, $\mathbf{k},\tilde{\mathbf{k}},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}\in\mathcal{T}$, $F,G\in\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$ (dense subspace of $L^{2}(\mathbb{R})$ defined in (2.4)), we have
    (1.17) $\mathcal{Q}\left(F\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},G\mathcal{P}_{\bar{\lambda}}\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)=\int_{\mathbb{R}}\widehat{F}(p)\overline{\widehat{G}(-p)}\mathrm{R}(\lambda+ip)\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})\frac{\mathrm{d}p}{\sqrt{\pi}}.$
  2. (2)
    For all $F\in\mathcal{C}_{\mathrm{comp}}$ (dense subset of $L^{2}(\nu)$ defined in (2.6)), the following equality holds in $L^{2}(\nu)$:
    (1.18) $F=\int_{\mathbb{R}}\frac{1}{\mathrm{R}(ip)}\sum_{\mathbf{k},\tilde{\mathbf{k}},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\mathcal{Q}(F,\mathcal{P}_{-ip}\Psi_{-ip,\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})\mathrm{B}_{ip}^{-1}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{ip}^{-1}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})e^{-(2ip+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{ip,\mathbf{k},\tilde{\mathbf{k}}}\frac{\mathrm{d}p}{\sqrt{\pi}}.$
  3. (3)
    For all $\lambda\in\mathbb{C}$ and $\mathbf{k},\tilde{\mathbf{k}},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}\in\mathcal{T}$ with $\mathrm{Re}(\lambda)<\frac{1}{2}(1-|\mathbf{k}|\wedge|\mathbf{k}^{\prime}|-|\tilde{\mathbf{k}}|\wedge|\tilde{\mathbf{k}}^{\prime}|-\frac{\kappa}{8})$, we have $|\mathcal{Q}_{\lambda}(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})|<\infty$ and
    (1.19) $\mathcal{Q}_{\lambda}(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})=\mathrm{R}(\lambda)\mathrm{B}_{\lambda}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime}).$
1.7. Uniqueness of restriction measures
Upon reading the proofs of the results stated above, the reader will notice that they use no specific properties of SLE; instead, we rely solely on conformal invariance and the restriction property. In fact, as our final result shows, these two properties are sufficient to completely characterise the measure. To state the result precisely, we first introduce the concepts at play.
For any domain $D\subset\widehat{\mathbb{C}}$, we denote by $\mu^{\mathrm{loop}}_{D}$ the Brownian loop measure in $D$, which is a measure on Brownian-type loops introduced by Lawler–Werner [LW04]; see below Proposition 4.6 for details. For any Borel sets $A,B\subset D$, we will denote by $\Lambda_{D}(A,B)$ the $\mu^{\mathrm{loop}}_{D}$-mass of loops touching both $A$ and $B$. Alternatively, and to make the link with [KS07], $\mathrm{exp}(\Lambda_{D}(A,B))$ can be defined as a ratio of $\zeta$-regularised determinants; see [Dub09b, Proposition 2.1]. When $\partial D$ is non polar and $A$ and $B$ are disjoint closed sets, $\Lambda_{D}(A,B)$ is a finite quantity. If $D=\widehat{\mathbb{C}}$ and $A$ and $B$ are disjoint non polar subsets, this quantity is infinite but can be renormalised as follows (see [FL13, Theorem 1.1]):
(1.20) $\Lambda^{*}(A,B):=\lim_{r\to 0}\Lambda_{O_{r}}(A,B)-\log|\log r|,$
where $O_{r}$ is the Riemann sphere minus a ball of radius $r$.
Definition 1.8 (Restriction measures).
Fix $c_{\mathrm{m}}\leq 1$. Let $\nu$ be a Borel measure on the space $\mathcal{J}$ (resp. $\mathcal{J}_{0,\infty}$) of Jordan curves in the Riemann sphere (resp. separating 00 from $\infty$). For any simply connected domain $D$, denote by
(1.21) $\mathrm{d}\nu_{D}(\eta):=\mathds{1}_{\eta\subset D}e^{\frac{c_{\mathrm{m}}}{2}\Lambda^{*}(\eta,\partial D)}\mathrm{d}\nu(\eta).$
We say that $\nu$ is a restriction measure on $\mathcal{J}$ (resp. a weak restriction measure on $\mathcal{J}_{0,\infty}$), with central charge $c_{\mathrm{m}}$, if for any simply connected domains $D_{1},D_{2}$ (resp. both containing 00 and included in $^{}^\widehat{\mathbb{C}}\setminus\{\infty\}$) and any conformal transformation $\Phi:D_{1}\to D_{2}$, $\nu_{D_{2}}=\Phi^{*}\nu_{D_{1}}$.
Note that, in the above definition of weak restriction measures on $\mathcal{J}_{0,\infty}$, there is no assumption of invariance under conformal maps which exchange the roles of 00 and $\infty$. The word “weak” refers to this weak notion of conformal invariance. This will be used in Section 7.1 in order to give a new proof of reversibility of SLE.
Restriction measures on $\mathcal{J}$, also referred to as Malliavin–Kontsevich–Suhov measures (or simply MKS measures), have a long history. In [AM01], Airault and Malliavin introduced the notion of unitarising measures for the Virasoro algebra from an axiomatic point of view [^2] In our notations, their definition would require $\mathbf{L}_{n}^{*}=\mathbf{L}_{-n}$ on the $L^{2}$-space of the measure, which is impossible due to the lack of unitarity when the central charge is less than 1. The correct notion of unitarising measure is the one studied in this paper where we construct the Shapovalov form as a Hermitian form distinct from the $L^{2}$-product., leaving the construction of such measures open. Interestingly, this notion appeared almost simultaneously as Schramm’s original paper on SLE [Sch00]. In parallel, Lawler, Schramm and Werner showed the restriction property for chordal SLE [LSW03]. Then, Kontsevich and Suhov related the restriction property to unitarising measures and conjectured that, for each central charge $c_{\mathrm{m}}\leq 1$, there exists a unique (up to multiplicative constant) measure on Jordan curves satisfying the properties of Definition 1.8 [KS07, Conjecture 1]. The construction of such loop measures was first done in the special cases $c_{\mathrm{m}}=0$ [Wer08] and $c_{\mathrm{m}}=-2$ [BD16]. A construction for the range $c_{\mathrm{m}}\in(0,1]$ was later provided by Kemppainen and Werner [KW16] based on Conformal Loop Ensemble (counting measure of nested CLE). Finally, Zhan’s construction [Zha21] of the SLE loop measure covers the existence part of Kontsevich–Suhov’s conjecture in the whole region $c_{\mathrm{m}}\leq 1$. In the range $c_{\mathrm{m}}\in(0,1]$, Zhan’s construction is closely related to the argument of Kemppainen–Werner showing reversibility [KW16], and indeed the two measures differ only by a multiplicative constant [AS21, Theorem 2.18]. We also mention that the articles [Ben16, MP24] study more general measures where the conformal restriction anomaly in (1.21) is not necessarily given by the Brownian loop measure. Here, we show the uniqueness part for loops in the sphere.
Theorem 1.9 (Uniqueness of restriction measures).
Let $c_{\mathrm{m}}\leq 1$. Up to a positive multiplicative constant, there exists a unique restriction measure on $\mathcal{J}$ (resp. weak restriction measure on $\mathcal{J}_{0,\infty}$), with central charge $c_{\mathrm{m}}$, that is non-zero and locally finite. This measure is the SLEκ loop measure (resp. the SLEκ loop measure restricted to the space $\mathcal{J}_{0,\infty}$), for the unique $\kappa\in(0,4]$ such that $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}$.
As explained by Kontsevich and Suhov [KS07, Section 2.5], the problem of existence and uniqueness of MKS measures on general Riemann surfaces equipped with a projective structure (see Section 8.4 and [BD16] for more details on this general setting of Riemann surfaces) actually reduces to the case of the Riemann sphere.
We mention that uniqueness was already known in the special case $c_{\mathrm{m}}=0$ ($\kappa=8/3$); see [Wer08, CP14]. We also point out that [BD16] shows the existence of a restriction measure with central charge $c_{\mathrm{m}}=-2$, and uses an independent argument (relying on the convergence of the uniform spanning tree to SLE8 and the duality $\kappa\leftrightarrow\frac{16}{\kappa}$) to prove that this measure is locally absolutely continuous with respect to SLE2. Our theorem implies that this measure is in fact the SLE2 loop measure.
The main difficulty in proving this uniqueness statement is that there is no way a priori to compare two restriction measures (they need not be absolutely continuous with respect to one another). As a matter of comparison, Lawler [Law09, Law11] uses the restriction property to define SLE in multiply connected topologies as absolutely continuous measures with respect to the known SLE measure in simply connected domains. Uniqueness is then much clearer in this type of context.
To prove Theorem 1.9, we show that the $L^{2}$ space of any MKS measure carries a rich algebraic structure (Virasoro representation, Shapovalov form, Verma modules, etc.) which entirely characterises this $L^{2}$ space (via the spectral decomposition) and thus the measure. See Section 6 for details. In fact, our proof shows that the integration by parts formula is sufficient to characterise the measure uniquely; translated in the language of [KS07, Definition 2.4], this means that the “infinitesimally $c_{\mathrm{m}}$-restriction covariant” measures are also unique. The proof can also be understood as harmonic analysis on the group of diffeomorphisms of the unit circle, since $\mathcal{J}_{0,\infty}$ is an orbit of this group [TT06]. However, one of the most remarkable feature of non-unitarity is that the usual Plancherel formula (for the $L^{2}$-inner product) is replaced by our spectral resolution of the Shapovalov form (Theorem 1.7). Thus, we are dealing with an infinite dimensional, non-unitary version of the Peter-Weyl theorem.
This characterisation of SLE in terms of an infinitesimal restriction property could offer a new perspective on cutting/welding random surfaces. We plan to investigate this in a future work and explain this in more details in Section 8.7. We also discuss two consequences of this uniqueness result in Section 7; namely, the reversibility and the duality of SLE loop measures.
1.8. Organisation and notations
The paper is organised as follows:
  • $\bullet$
    Section 2: we construct our representations of the Witt and Virasoro algebras.
  • $\bullet$
    Section 3: we construct and classify the highest-weight representations.
  • $\bullet$
    Section 4: we derive an integration by parts formula and build in this way the Shapovalov form.
  • $\bullet$
    Section 5: we prove the spectral decomposition of the Shapovalov form.
  • $\bullet$
    Section 6: we provide our proof of the uniqueness of MKS measure.
  • $\bullet$
    Section 7: we discuss two consequences of our uniqueness result: the reversibility and the duality of SLE loop measures.
  • $\bullet$
    Section 8: the current paper lays the groundwork for an in-depth study of the CFT of SLE. We explain this programme in some details in this section.
  • $\bullet$
    Appendix A: we review some standard facts of the representation theory of the Virasoro algebra.
  • $\bullet$
    Appendix B: we show that the SLE loop measure is a Borel measure on the space of simple Jordan curves.
Here is a list of the main notations we use in this paper.
  • $\bullet$
    $\mathbb{D}$: the unit disc. $\mathbb{D}^{*}=\widehat{\mathbb{C}}\setminus\bar{\mathbb{D}}$: the exterior disc. $|\mathrm{d}z|^{2}$: the Euclidean volume on $\mathbb{C}$.
  • $\bullet$
    $\mathbb{C}[z]$: Polynomials in $z$. $\mathbb{C}(z)$: Laurent polynomials in $z$.
  • $\bullet$
    $\mathrm{v}_{n},n\in\mathbb{Z}$: Vector field $\mathrm{v}_{n}=-z^{n+1}\partial_{z}$.
  • $\bullet$
    $\bar{f}:z\mapsto\overline{f(\bar{z})}$ and similar notations for $\bar{g}$, $\bar{\mathrm{v}}=\bar{v}(z)\partial_{z}$, etc.
  • $\bullet$
    $\mathcal{J}$ (resp. $\mathcal{J}_{0,\infty}$): Jordan curves in $\widehat{\mathbb{C}}$ (resp. disconnecting 0 from $\infty$).
  • $\bullet$
    $\mathcal{E}$: Space of conformal maps $f:\mathbb{D}\to\mathbb{C}$ with continuous, injective extension to $\bar{\mathbb{D}}$, $f(0)=0$, $f^{\prime}(0)=1$.
  • $\bullet$
    $\tilde{\mathcal{E}}$: Space of conformal maps $g:\mathbb{D}^{*}\to\widehat{\mathbb{C}}$ with continuous, injective extension to $\overline{\mathbb{D}^{*}}$, $g(\infty)=\infty$, $g^{\prime}(\infty)=1$.
  • $\bullet$
    $c:\mathcal{J}_{0,\infty}\to\mathbb{R}$, $f:\mathcal{J}_{0,\infty}\to\mathcal{E}$, $a_{m}:\mathcal{J}_{0,\infty}\to\mathbb{C}$. Coordinates on $\mathcal{J}_{0,\infty}$.
  • $\bullet$
    $\tilde{c}:\mathcal{J}_{0,\infty}\to\mathbb{R}$, $g:\mathcal{J}_{0,\infty}\to\tilde{\mathcal{E}}$, $b_{m}:\mathcal{J}_{0,\infty}\to\mathbb{C}$. Another set of coordinates on $\mathcal{J}_{0,\infty}$.
  • $\bullet$
    $\nu$: SLEκ loop measure restricted to $\mathcal{J}_{0,\infty}$. $\nu^{\#}$: shape measure on $\mathcal{E}$. $\mathbb{E}^{\#}$: expectation wrt $\nu^{\#}$.
  • $\bullet$
    $\Theta$: reflection operator (1.6).
  • $\bullet$
    $\mathcal{Q}$: Shapovalov form (1.7).
  • $\bullet$
    $\mathcal{V}_{\lambda}$: highest-weight representation of weight $\lambda$ (1.12).
  • $\bullet$
    $\mathcal{Q}_{\lambda}$: Shapovalov form of weight $\lambda$ (1.16).
  • $\bullet$
    $\mathcal{T}$ ($\mathcal{T}_{N}$): set of integer partitions (of $N$), see Section 1.5.
  • $\bullet$
    $\vartheta$: Differential of the Weil-Petersson potential (1.3).
  • $\bullet$
    $\varpi$: Differential of the Velling-Kirillov potential (2.8).
  • $\bullet$
    $\omega$: Virasoro cocycle (1.5).
  • $\bullet$
    $O_{\eta}$: for any Jordan curve $\eta\in\mathcal{J}_{0,\infty}$, annulus defined by
    (1.22) $O_{\eta}:=\{z\in\mathbb{C}|\,e^{c}/8<|z|<8e^{-\tilde{c}}\}.$
    By the Koebe quarter theorem, for any $\eta\in\mathcal{J}_{0,\infty}$, we have $\frac{e^{c}}{4}\mathbb{D}\subset\mathrm{int}(\eta)$ and $4e^{-\tilde{c}}\mathbb{D}^{*}\subset\mathrm{ext}(\eta)$, so that $O_{\eta}$ contains $\eta$ in its interior.
2. Witt and Virasoro representations
This section is organised as follows. In Section 2.2, we define two subspaces $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$ of $L^{2}(\nu)$ of “test functions” and show that they are dense. In Section 2.3, we construct two commuting representations of the Witt algebra on $\mathcal{C}$. We will then modify these representations to get two commuting representations of the Virasoro algebra on $\mathcal{C}_{\mathrm{comp}}$ in Section 2.4 (we will later show in Theorem 4.1 that these operators are closable on $L^{2}(\nu)$ as a corollary of the integration by parts formula). These operators are obtained from the infinitesimal action of the group of quasiconformal homeomorphisms of the Riemann sphere on the space of Jordan curves. We will start by recalling some related background in Section 2.1.
2.1. Background on quasiconformal mappings
This section gives some background on quasiconformal homeomorphisms of the Riemann sphere. We refer to [Ahl87] for more details.
A quasiconformal homeomorphism of $\widehat{\mathbb{C}}$ is a homeomorphism $\Phi$ which is differentiable almost everywhere and satisfies
(2.1) $\partial_{\bar{z}}\Phi=\mu\partial_{z}\Phi$
almost everywhere for some $\mu\in L^{\infty}(\mathbb{C})$ with $\left\|\mu\right\|_{L^{\infty}(\mathbb{C})}<1$. The function $\mu$ is called the Beltrami coefficient of $\Phi$. Conversely, for all $\mu\in L^{\infty}(\mathbb{C})$ with $\left\|\mu\right\|_{L^{\infty}(\mathbb{C})}$, the Beltrami equation (2.1) admits a solution in the space of quasiconformal homeomorphisms of $\widehat{\mathbb{C}}$, and this solution is unique modulo Möbius transformations [Ahl87, Chapter V, Theorem 3]. By Weyl’s lemma, the associated quasiconformal map is conformal outside the support of $\mu$. Given the form of the Beltrami equation, a Beltrami coefficient transforms as a $(-1,1)$-tensor: under a conformal change of coordinates $f$, a Beltrami coefficient pulls back to
(2.2) $f^{*}\mu=\mu\circ f\frac{\overline{f^{\prime}}}{f^{\prime}}.$
The quasiconformal group acts naturally on the set of Jordan curves by $\Phi\cdot[\eta]:=[\Phi\circ\eta]$. Of particular interest to us will be those quasiconformal maps which are conformal away from small neighbourhoods of 0 and $\infty$. These can be generated by the Witt algebra, and we will construct the Witt representations as Lie derivatives along the induced vector fields.
We will consider two different types of quasiconformal mappings:
Definition 2.1.
Let $\mu\in L^{\infty}(\mathbb{C})$ with $\left\|\mu\right\|_{L^{\infty}(\mathbb{C})}<1$ and s.t. $\mu$ is compactly supported in $\mathbb{D}$.
  • $\bullet$
    $\Phi_{\mu}$ is the unique solution to the Beltrami equation satisfying $\Phi_{\mu}(0)=0$, and $\Phi_{\mu}(z)=z+O_{z\to\infty}(1)$.
  • $\bullet$
    Extend $\mu$ to the whole complex plane by the formula $\mu(z):=\overline{\mu(1/\bar{z})}\frac{z^{2}}{\bar{z}^{2}}$ for $|z|>1$. Define $\tilde{\Phi}_{\mu}$ to be the unique solution to the Beltrami equation fixing $0,1,\infty$. By reflection symmetry, $\tilde{\Phi}_{\mu}$ fixes the unit circle setwise. Moreover, $\tilde{\Phi}_{\mu}$ is conformal in the neighbourhood of $\mathbb{S}^{1}$ (since $\mu$ vanishes in this neighbourhood), so $\tilde{\Phi}_{\mu}$ restricts to an analytic diffeomorphism of $\mathbb{S}^{1}$.
Next, we recall a variational formula for $\tilde{\Phi}_{\mu}$, i.e. we give a first order expansion of $\tilde{\Phi}_{t\mu}$ for small complex $t$ as found in [Ahl87, Chapter V, Theorem 5].
Proposition 2.2.
For all $\mu\in L^{\infty}(\mathbb{C})$ compactly supported in $\mathbb{D}$, we have
$\tilde{\Phi}_{t\mu}(z)=z-\frac{t}{\pi}\int_{\mathbb{D}}\frac{z(z-1)\mu(\zeta)}{\zeta(\zeta-1)(\zeta-z)}|\mathrm{d}\zeta|^{2}+\frac{\bar{t}}{\pi}\int_{\mathbb{D}}\frac{z(z-1)\overline{\mu(\zeta)}}{\bar{\zeta}(\bar{\zeta}-1)(1-z\bar{\zeta})}|\mathrm{d}\zeta|^{2}+o(t),$
where $o(t)\to 0$ as $t\to 0$ uniformly on compacts.
Proof.
[Ahl87, Chapter V, Theorem 5] and the symmetry of $\mu$ gives
(2.3) $\tilde{\Phi}_{t\mu}(z)=z-\frac{1}{\pi}\int_{\mathbb{D}}\frac{z(z-1)t\mu(\zeta)}{\zeta(\zeta-1)(\zeta-z)}|\mathrm{d}\zeta|^{2}-\frac{1}{\pi}\int_{\mathbb{D}^{*}}\frac{z(z-1)\overline{t\mu(1/\bar{z})}}{\zeta(\zeta-1)(\zeta-z)}|\mathrm{d}\zeta|^{2}+o(t).$
By the change of variable $z\mapsto 1/\bar{z}$, the second integral evaluates to
$\bar{t}\int_{\mathbb{D}}\frac{z(z-1)\overline{\mu(z)}}{\bar{\zeta}^{-1}(\bar{\zeta}^{-1}-1)(\bar{\zeta}^{-1}-z)}\frac{1}{\bar{\zeta}^{4}}|\mathrm{d}\zeta|^{2}=-\bar{t}\int_{\mathbb{D}}\frac{z(z-1)\overline{\mu(z)}}{\bar{\zeta}(\bar{\zeta}-1)(1-z\bar{\zeta})}|\mathrm{d}\zeta|^{2}.$
We conclude with a similar statement for the quasiconformal homeomorphism $\Phi_{t\mu}$. A remarkable feature of this variational formula is that it is analytic in $\mu$, i.e. there is no $\bar{t}$-term.
Lemma 2.3.
Let $r>0$. Let $\mathrm{v}=v\partial_{z}\in z^{-1}\mathbb{C}[z^{-1}]\partial_{z}$ with expansion $v(z)=-\sum_{n=2}^{\infty}v_{n}z^{-n+1}$, and define
$\mu(z):=\sum_{n=2}^{\infty}nv_{n}r^{-2n}\bar{z}^{n-1}z\mathds{1}_{|z|<r}.$
Then, $\Phi_{t\mu}$ is conformal in $r\mathbb{D}^{*}$, and we have $\Phi_{t\mu}(z)=z+tv(z)+o(t)$ uniformly on compacts.
Proof.
The solution theory of the Beltrami equation gives $\Phi_{t\mu}(z)=z+tv_{\mu}(z)+o(t)$ uniformly for $z\in\widehat{\mathbb{C}}$, for some vector field $v_{\mu}$ satisfying $\partial_{\bar{z}}v=\mu$. Moreover, $\Phi_{t\mu}$ is conformal in $r\mathbb{D}^{*}$ and the normalisation of $\Phi_{t\mu}$ is such that $v_{\mu}(z)=O_{z\to\infty}(1)$ and $v_{\mu}(0)=0$. We will show that $v_{\mu}$ is the Cauchy transform of $\mu$.
By linearity, it suffices to work with the generators $\mathrm{v}_{-n}=-z^{-n+1}\partial_{z}$ for $n\geq 2$, so we consider the Beltrami differential $\mu_{n}(z)=n\epsilon^{-2n}\bar{z}^{n-1}z\mathds{1}_{|z|<\epsilon}$. We have
$\frac{1}{\pi}\int_{\widehat{\mathbb{C}}}\frac{\mu_{n}(\zeta)}{\zeta-z}|\mathrm{d}w|^{2}$ $=-\frac{1}{\pi}\int_{r\mathbb{D}}\frac{\bar{\zeta}^{n-1}\zeta}{z(1-\frac{\zeta}{z})}|\mathrm{d}w|^{2}=-\frac{1}{\pi}\int_{r\mathbb{D}}\sum_{m=0}^{\infty}\bar{\zeta}^{n-1}w^{m+1}z^{-m-1}|\mathrm{d}w|^{2}$
$=-\frac{z^{-n+1}}{\pi}\int_{r\mathbb{D}}|\zeta|^{2n-2}|\mathrm{d}\zeta|^{2}=-\frac{r^{2n}}{n}z^{-n+1}.$
Hence, the Cauchy transform of $\mu$ coincides with $v$ in $r\mathbb{D}^{*}$, and we keep denoting by $v$ this Cauchy transform in the whole complex plane. By uniqueness, $v-v_{\mu}$ is a polynomial vector field on $\widehat{\mathbb{C}}$ (i.e. an infinitesimal conformal transformation of $\widehat{\mathbb{C}}$), so it suffices to check that $v$ has the desired normalisation. On the one hand, we have $v(z)=O_{z\to\infty}(1)$, so we have the desired behaviour at $\infty$. On the other hand, the Cauchy transform of $\mu$ at 0 evaluates to 0 since $\int_{r\mathbb{D}}\bar{\zeta}^{n-1}|\mathrm{d}\zeta|^{2}=0$ for all $n\geq 2$. Hence $v(0)=0$ and we deduce that $v=v_{\mu}$. ∎
2.2. Dense subspaces of $L^{2}(\nu)$
In this section, we introduce two subsets $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$ of $L^{2}(\nu)$ and show that they are dense. We start by defining an “exponential Schwartz space” by
(2.4) $\mathcal{S}_{\mathrm{exp}}(\mathbb{R}):=\left\{F\in\mathcal{C}^{\infty}(\mathbb{R})|\,\forall k,l\in\mathbb{N},\,\sup_{c\in\mathbb{R}}\,e^{k|c|}|F^{(l)}(c)|<\infty\right\}.$
This is a Fréchet space with topology induced by the countable family of seminorms $\left\|F\right\|_{k,l}:=\sup_{c\in\mathbb{R}}e^{k|c|}|F^{(l)}(c)|$. We denote its continuous dual by $\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})$. Recalling the coordinates $(a_{m})_{m\geq 1}$ introduced in (1.1), we consider the spaces
(2.5) $\hskip 45.0pt\mathcal{C}:=\mathcal{S}_{\mathrm{exp}}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]\hskip 40.0pt\text{and}$
(2.6) $\mathcal{C}_{\mathrm{comp}}=\left\{F\in\mathcal{C}|\,\exists\epsilon>0,\,\mathrm{supp}(F)\subset\{\eta\subset\{\epsilon<|z|<\epsilon^{-1}\}\}\right\}.$
Importantly, by the Koebe quarter theorem, for all $F\in\mathcal{C}_{\mathrm{comp}}$, there is $\epsilon>0$ such that $c,\tilde{c}>\log\epsilon$ on the support of $F$ (recall that $c$ and $\tilde{c}$ are the log-conformal radius of the interior and exterior of the curve seen from 0 and $\infty$ respectively).
The spaces $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$ will play the role of spaces of “elementary functions”, which will serve as the common domains of our operators (before considering their closure).
Lemma 2.4.
The spaces $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$ are dense subspaces of $L^{2}(\nu)$.
Proof.
We first note that $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]\subset L^{2}(\nu^{\#})$ by de Branges’ theorem. Indeed, for any map $(z\in\mathbb{D}\mapsto z+\sum_{m\geq 1}a_{m}z^{m+1})\in\mathcal{E}$, we can bound $|a_{m}|\leq(m+1)$ and so any polynomial in $(a_{m},\bar{a}_{m})_{m\geq 1}$ is uniformly bounded on $\mathcal{E}$ and in particular belongs to $L^{2}(\nu^{\#})$.
Next, we prove that $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ is dense in $L^{2}(\nu^{\#})$. Let $\mathcal{U}$ be the space of univalent functions $f$ on $\mathbb{D}$, with $f(0)=0$ and $f^{\prime}(0)=1$. This space comes equipped with the topology of uniform convergence of compacts of $\mathbb{D}$. This topology is metrisable, for instance with the distance $d_{\mathcal{U}}(f_{1},f_{2}):=\sum_{n=1}^{\infty}2^{-n}\left\|f_{1}-f_{2}\right\|_{\mathcal{C}^{0}((1-2^{-n})\bar{\mathbb{D}})}\wedge 1$. Of course, we have a continuous inclusion $\mathcal{E}\subset\mathcal{U}$ and every $f\in\mathcal{U}$ has a power series expansion of the form (1.1). Therefore, the space $\mathbb{C}[(a_{m},\bar{a}_{m})_{\geq 1}]$ is well defined as a subspace of $\mathcal{C}^{0}(\mathcal{U})$.
According to [Pom75, Theorem 1.7], $\mathcal{U}$ is compact. Moreover, $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ separates points: for all $f_{1},f_{2}\in\mathcal{U}$, we have $a_{m}(f_{1})=a_{m}(f_{2})$ for all $m\geq 1$ if and only if $f_{1}=f_{2}$. Hence, $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ is dense in $\mathcal{C}^{0}(\mathcal{U})$ by the Stone-Weierstrass theorem.
The spaces $\mathcal{E}$ and $\mathcal{U}$ are Polish (separable, completely metrisable), so they are also Radon spaces. Since $\nu_{\widehat{\mathbb{C}}}$ is a Borel measure on $\mathcal{E}$ (Proposition B.1), it implies that it is a Radon measure. We can also view it as a Radon measure on $\mathcal{U}$ giving full mass to $\mathcal{E}$. It follows that $\mathcal{C}^{0}(\mathcal{U})$ is dense in $L^{2}(\nu^{\#})$, since $\mathcal{U}$ is compact Hausdorff [Rud87, Theorem 3.13].
Wrapping up, let $F\in L^{2}(\mathcal{E},\nu^{\#})\simeq L^{2}(\mathcal{U},\nu^{\#})$ and $\epsilon>0$. By the previous points, there exist $G\in\mathcal{C}^{0}(\mathcal{U})$ and $P\in\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ such that $\left\|F-G\right\|_{L^{2}(\nu^{\#})}\leq\frac{\epsilon}{2}$ and $\left\|P-G\right\|_{\mathcal{C}^{0}(\mathcal{U})}\leq\frac{\epsilon}{2}$. Since $\nu^{\#}$ is a probability measure, we have
$\left\|P-F\right\|_{L^{2}(\nu^{\#})}\leq\left\|P-G\right\|_{L^{2}(\nu^{\#})}+\left\|G-F\right\|_{L^{2}(\nu^{\#})}\leq\left\|P-G\right\|_{\mathcal{C}^{0}(\mathcal{U})}+\frac{\epsilon}{2}\leq\epsilon.$
This proves that $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ is dense in $L^{2}(\nu^{\#})$. Finally, $\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$ is dense in $L^{2}(\mathrm{d}c)$, so $\mathcal{C}=\mathcal{S}_{\mathrm{exp}}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ is dense in $L^{2}(\nu)=L^{2}(\mathrm{d}c)\otimes L^{2}(\nu^{\#})$. The result concerning $\mathcal{C}_{\mathrm{comp}}$ also follows. ∎
2.3. Witt algebra
In this section, we construct and study two commuting representations $(\mathcal{L}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ and (¯$(\bar{\mathcal{L}}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ of the Witt algebra as endomorphisms of the dense subspace $\mathcal{C}$ (2.5) of $L^{2}(\nu)$. The results of this section are similar to [GQW24, Section 3.2] and [CP14, Section 3], which are obtained from the variational formulas of [DS63]. Since we were not aware of this reference, we rederived these formulas from scratch using a different method, namely Ahlfors’ variational formulas for quasiconformal mappings [Ahl87].
The intellectually satisfying aspect of working with quasiconformal maps is that we get an action of a whole group, rather than a mere infinitesimal action (the complex Witt algebra does not exponentiate since there is not complexification of $Diff_{+}(\mathbb{S}^{1})$ [Seg04]). So we can view the Witt algebra as a subalgebra of the Lie algebra of infinitesimal Beltrami differentials, and the representations are obtained as (complex) Lie derivatives along the induced vector fields.
Definition 2.5.
Let $F\in L^{2}(\nu)$ and $\mathrm{v}=v(z)\partial_{z}\in\mathbb{C}(z)\partial_{z}$. We say that $F$ is differentiable at $\eta\in\mathcal{J}_{0,\infty}$ in direction $\mathrm{v}$ if there exist $\mathcal{L}_{\mathrm{v}}F(\eta),\bar{\mathcal{L}}_{\mathrm{v}}(\eta)\in\mathbb{C}$ such that:
(2.7) $F(\phi_{t}\cdot\eta)=F(\eta)+t\mathcal{L}_{\mathrm{v}}F(\eta)+\bar{t}\bar{\mathcal{L}}_{\mathrm{v}}F(\eta)+o(t)$
where $\phi_{t}(z)=z+tv(z)+o(t)$ is the flow of $\mathrm{v}$, defined in $O_{\eta}$ (recall (1.22)) for small complex $t$, as the solution to the differential equation
$\partial_{t}\phi_{t}=v\circ\phi_{t},\qquad\phi_{0}=\mathrm{id}_{O_{\eta}}.$
The function $F$ is said to be differentiable (resp. $\nu$-a.e.) in direction $\mathrm{v}$ if it is differentiable at every (resp. for $\nu$-almost every) $\eta\in\mathcal{J}_{0,\infty}$ in direction $\mathrm{v}$. In this case, the differentials $\mathcal{L}_{\mathrm{v}}F$, ¯$\bar{\mathcal{L}}_{\mathrm{v}}F$ define complex-valued functions on $\mathcal{J}_{0,\infty}$ (resp. defined $\nu$-almost everywhere).
Note that with this definition and for $\mathrm{v}\equiv 0$, $\mathcal{L}_{\mathrm{v}}$ is well defined on $L^{2}(\nu)$ and $\mathcal{L}_{\mathrm{v}}\equiv 0$.
Section 2.3.1 contains the main statements about the Witt representations: Propositions 2.6, 2.8 and 2.9. The respective proofs are given in Sections 2.3.2, 2.3.3 and 2.3.4.
2.3.1. Statements on the Witt algebra
The first main result concerns the construction of the operators $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$, $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, as endomorphisms of $\mathcal{C}$ (2.5) and $\mathcal{C}_{\mathrm{comp}}$ (2.6) as well as the computation of their commutation relations:
Proposition 2.6.
Let $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$. Every $F\in\mathcal{C}$ (resp. $\mathcal{C}_{\mathrm{comp}}$) is differentiable in direction $\mathrm{v}$, and $\mathcal{L}_{\mathrm{v}}F,\bar{\mathcal{L}}_{\mathrm{v}}F\in\mathcal{C}$ (resp. $\mathcal{C}_{\mathrm{comp}}$). The maps $F\mapsto\mathcal{L}_{\mathrm{v}}F$, $F\mapsto\bar{\mathcal{L}}_{\mathrm{v}}F$ define endomorphisms of $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$, and their commutators satisfy
$[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]=\mathcal{L}_{[\mathrm{v},\mathrm{w}]},\qquad[\bar{\mathcal{L}}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]=\bar{\mathcal{L}}_{[\mathrm{v},\mathrm{w}]}\quad\text{and}\quad[\mathcal{L}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]=0,$
where $[\mathrm{v},\mathrm{w}]$ is the Lie bracket of vector fields (1.2).
Recalling that we denote for all $n\in\mathbb{Z}$, $\mathrm{v}_{n}=-z^{n+1}\partial_{z}$, we will simply write $\mathcal{L}_{n}$ for $\mathcal{L}_{\mathrm{v}_{n}}$.
Remark 2.7.
We would like to explain that, when the vector field $\mathrm{v}\in\mathbb{C}_{2}[z]\partial_{z}$ is a quadratic polynomial vector field, one can directly build the operator $\mathcal{L}_{\mathrm{v}}$ on $L^{2}(\nu)$ using conformal perturbations and Stone’s theorem. Recall that $\mathbb{C}_{2}[z]\partial_{z}\simeq\mathfrak{sl}_{2}(\mathbb{C})$ is the Lie algebra of the group $\mathrm{PSL}_{2}(\mathbb{C})$ of Möbius transformations of the Riemann sphere $\widehat{\mathbb{C}}$ which acts on $\mathcal{J}_{0,\infty}$ by $\Phi\cdot\eta:=[\Phi\circ\eta]$, for all $\eta\in\mathcal{J}_{0,\infty}$, $\Phi\in\mathrm{PSL}_{2}(\mathbb{C})$. This action defines a unitary group of operators $(\mathbf{U}_{\Phi})_{\Phi\in\mathrm{PSL}_{2}(\mathbb{C})}$ on $L^{2}(\nu)$:
$\mathbf{U}_{\Phi}F(\eta):=F(\Phi\cdot\eta).$
Using [RS80, Theorem VIII.9], one can easily check that the mapping $\Phi\mapsto\mathbf{U}_{\Phi}$ is a strongly continuous group of unitary operators on $L^{2}(\nu)$. By Stone’s theorem [RS80, Theorem VIII.8], this fact implies that for every $\mathrm{v}\in\mathbb{C}_{2}[z]\partial_{z}\simeq\mathfrak{sl}_{2}(\mathbb{C})$, there exists a (unique) self-adjoint operator $\mathbf{H}_{\mathrm{v}}$ on $L^{2}(\nu)$ such that
$\mathbf{U}_{e^{\mathrm{v}}}=e^{i\mathbf{H}_{\mathrm{v}}}.$
In the above display, $\mathbb{C}_{2}[z]\partial_{z}\simeq\mathfrak{sl}_{2}(\mathbb{C})$ is a six-dimensional real vector space and the exponential of $\mathrm{v}\in\mathbb{C}_{2}[z]\partial_{z}$ corresponds to the matrix exponential in $\mathfrak{sl}_{2}(\mathbb{C})$. The Hamiltonian is the operator
$\mathbf{H}:=\mathbf{H}_{-z\partial_{z}},$
which is the generator of dilations $z\mapsto e^{-t}z$. The operators $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$ can then be obtained by setting
$\mathcal{L}_{\mathrm{v}}:=\frac{i}{2}(\mathbf{H}_{\mathrm{v}}-i\mathbf{H}_{i\mathrm{v}});\qquad\bar{\mathcal{L}}_{\mathrm{v}}:=\frac{i}{2}(\mathbf{H}_{\mathrm{v}}+i\mathbf{H}_{i\mathrm{v}}).$
The next proposition gives the explicit expression for $\mathcal{L}_{0}$, ¯$\bar{\mathcal{L}}_{0}$, $\mathcal{L}_{-1}$ and ¯$\bar{\mathcal{L}}_{-1}$, which will be of use in the sequel. We mention that expressions with a similar flavour have been given for Kirillov’s operators; see for instance [Kir98, (7-0’) and (7-1’)] and recall that the difference with our setup has been described after Theorem 1.2.
Proposition 2.8.
We have
$\mathcal{L}_{0}=-\frac{\partial_{c}}{2}+\frac{1}{2}\sum_{m=1}^{\infty}m(a_{m}\partial_{a_{m}}-\bar{a}_{m}\partial_{\bar{a}_{m}});\qquad\bar{\mathcal{L}}_{0}=-\frac{\partial_{c}}{2}-\frac{1}{2}\sum_{m=1}^{\infty}m(a_{m}\partial_{a_{m}}-\bar{a}_{m}\partial_{\bar{a}_{m}}),$
$\mathcal{L}_{-1}=e^{-c}a_{1}\partial_{c}+e^{-c}\sum_{m=1}^{\infty}(m+2)(a_{m+1}-a_{1}a_{m})\partial_{a_{m}}+m(a_{1}\bar{a}_{m}-\bar{a}_{m-1})\partial_{\bar{a}_{m}},$
¯$\bar{\mathcal{L}}_{-1}=e^{-c}\bar{a}_{1}\partial_{c}+e^{-c}\sum_{m=1}^{\infty}(m+2)(\bar{a}_{m+1}-\bar{a}_{1}\bar{a}_{m})\partial_{\bar{a}_{m}}+m(\bar{a}_{1}a_{m}-a_{m-1})\partial_{a_{m}},$
where we have set $a_{0}=1$.
Finally, we consider the case of the coordinate function $c:\mathcal{J}_{0,\infty}\to\mathbb{R}$ and compute explicitely $\mathcal{L}_{\mathrm{v}}c$ for all vector field $\mathrm{v}$. Before stating the result, we define $\varpi:\mathcal{J}_{0,\infty}\times\mathbb{C}(z)\partial_{z}\to\mathbb{C}$ by
(2.8) $\varpi_{e^{c}f}(v\partial_{z}):=\frac{e^{-c}}{4i\pi}\oint\left(\frac{(f^{-1})^{\prime}(z)}{f^{-1}(z)}\right)^{2}v(e^{c}z)\mathrm{d}z,$
where the contour is a sufficiently small loop around 0.
Proposition 2.9.
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, the coordinate $c:\mathcal{J}_{0,\infty}\to\mathbb{R}$ is differentiable in direction $\mathrm{v}$, and
$\mathcal{L}_{\mathrm{v}}c=\varpi(\mathrm{v}).$
This result has been known for other topologies (smooth Jordan curves [KY87], Weil-Petersson class [TT06]). Our strategy will follow closely the proof of [TT06, Part 2, Theorem 4.5].
The rest of Section 2.3 is dedicated to the proof of Propositions 2.6, 2.8 and 2.9.
2.3.2. Computation of $\mathcal{L}_{0}$, ¯$\bar{\mathcal{L}}_{0}$, $\mathcal{L}_{-1}$ and ¯$\bar{\mathcal{L}}_{-1}$
Proof of Proposition 2.8.
We will compute successively $\mathbf{H}_{\mathrm{v}_{0}}$, $\mathbf{H}_{i\mathrm{v}_{0}}$, $\mathbf{H}_{\mathrm{v}_{-1}}$, and $\mathbf{H}_{i\mathrm{v}_{-1}}$, the existence of which follows from Remark 2.7 (or simply from the arguments below).
Case $\mathbf{H}_{\mathrm{v}_{0}}$. The flow generated by $\mathrm{v}_{0}=-z\partial_{z}$ is $z\mapsto e^{-t}z$. In the coordinates $(c,f)\in\mathbb{R}\times\mathcal{E}$, the motion of the curve is $(c_{t},f_{t})=(c-t,f)$. Hence,
$\mathbf{H}_{\mathrm{v}_{0}}=i\partial_{c}.$
Case $\mathbf{H}_{i\mathrm{v}_{0}}$. The flow generated by $i\mathrm{v}_{0}$ is $z\mapsto e^{-it}z$, creating the motion $(c_{t},f_{t})=(c,e^{-it}f((e^{it}\cdot))$ in the coordinate $(c,f)\in\mathbb{R}\times\mathcal{E}$. For the Taylor coefficients of $f$ as in (1.1), we get $a_{m}(t)=e^{imt}a_{m}$, so that
$\mathbf{H}_{i\mathrm{v}_{0}}=\sum_{m=1}^{\infty}(ma_{m}\partial_{a_{m}}-m\bar{a}_{m}\partial_{\bar{a}_{m}}).$
The formulas for $\mathcal{L}_{0}=\frac{i}{2}(\mathbf{H}_{\mathrm{v}_{0}}-i\mathbf{H}_{i\mathrm{v}_{0}})$ and ¯$\bar{\mathcal{L}}_{0}=\frac{i}{2}(\mathbf{H}_{\mathrm{v}_{0}}+i\mathbf{H}_{i\mathrm{v}_{0}})$ follow.
Case $\mathbf{H}_{\mathrm{v}_{-1}}$. The flow generated by $-\partial_{z}$ is $z\mapsto z-t$. Let $(c_{t},f_{t})\in\mathbb{R}\times\mathcal{E}$ be uniformising the interior of $-t+\eta(\mathbb{S}^{1})$. Since $e^{c}f-t$ is another uniformising map, we have $e^{c_{t}}f_{t}=e^{c}f\circ\phi_{t}-t$ for some Möbius transformation $\phi_{t}$ of the unit disc. The generator of this transformation is of the form
$\phi_{t}(z)=z+t(\alpha-i\theta z-\bar{\alpha}z^{2})+o(t),$
for some $\alpha\in\mathbb{C}$ and $\theta\in\mathbb{R}$. We also have $c_{t}=c+tx+o(t)$ for some $x\in\mathbb{R}$. We have $0=e^{c_{t}}f_{t}(0)=e^{c}f(\phi_{t}(0))-t=t(\alpha e^{c}-1)+o(t)$, so that $\alpha=e^{-c}$. Moreover,
$1=f_{t}^{\prime}(0)=e^{c-c_{t}}\phi_{t}^{\prime}(0)f^{\prime}(\phi_{t}(0))$ $=(1-tx)(1-i\theta t)(1+\alpha tf^{\prime\prime}(0))+o(t)$
$=1+t(e^{-c}f^{\prime\prime}(0)-(x+i\theta))+o(t),$
so that
(2.9) $\alpha=e^{-c};\qquad x=e^{-c}\mathrm{Re}(f^{\prime\prime}(0));\qquad\theta=e^{-c}\mathrm{Im}(f^{\prime\prime}(0)).$
Setting $a_{0}:=1$, we have
$f_{t}(z)$ $=-te^{-c_{t}}+e^{c-c_{t}}\sum_{m=0}^{\infty}a_{m}\phi_{t}(z)^{m+1}$
$=-te^{-c_{t}}+(1-tx)\sum_{m=0}^{\infty}a_{m}\left(z^{m+1}+t(m+1)(\alpha-i\theta z-\bar{\alpha}z^{2})z^{m}\right)+o(t).$
Rearranging according to the power of $z$ and then using (2.9), we get that
$f_{t}(z)$ $=f(z)+t(\alpha-e^{-c})+t(-x+2a_{1}\alpha-i\theta)z$
$+t\sum_{m=1}^{\infty}((m+2)\alpha a_{m+1}-(x+i\theta)a_{m}-im\theta a_{m}-m\bar{\alpha}a_{m-1})z^{m+1}+o(t)$
$=f(z)+te^{-c}\sum_{m=1}^{\infty}((m+2)a_{m+1}-2a_{1}a_{m}-2im\mathrm{Im}(a_{1})a_{m}-ma_{m-1})z^{m+1}+o(t).$
It follows that
$i\mathbf{H}_{\mathrm{v}_{-1}}=2e^{-c}\mathrm{Re}(a_{1})\partial_{c}$ $+e^{-c}\sum_{m=1}^{\infty}\left((m+2)a_{m+1}-2a_{1}a_{m}-2im\mathrm{Im}(a_{1})a_{m}-ma_{m-1}\right)\partial_{a_{m}}$
$+e^{-c}\sum_{m=1}^{\infty}((m+2)\bar{a}_{m+1}-2\bar{a}_{1}\bar{a}_{m}+2im\mathrm{Im}(a_{1})\bar{a}_{m}-m\bar{a}_{m-1})\partial_{\bar{a}_{m}}.$
Case $\mathbf{H}_{i\mathrm{v}_{-1}}$. We apply the same strategy for the flow $z\mapsto z-it$. Using the same parametrisation, we have this time
$\alpha=ie^{-c};\qquad x=-e^{-c}\mathrm{Im}(f^{\prime\prime}(0));\qquad\theta=e^{-c}\mathrm{Re}(f^{\prime\prime}(0)).$
Replacing with these values gives
$i\mathbf{H}_{i\mathrm{v}_{-1}}=-2e^{-c}\mathrm{Im}(a_{1})\partial_{c}$ $+ie^{-c}\sum_{m=1}^{\infty}((m+2)a_{m+1}-2a_{1}a_{m}-2m\mathrm{Re}(a_{1})a_{m}+ma_{m-1})\partial_{a_{m}}$
$-ie^{-c}\sum_{m=1}^{\infty}((m+2)\bar{a}_{m+1}-2\bar{a}_{1}\bar{a}_{m}-2m\mathrm{Re}(a_{1})\bar{a}_{m}+m\bar{a}_{m-1})\partial_{\bar{a}_{m}}.$
The formulas for $\mathcal{L}_{-1}=\frac{i}{2}(\mathbf{H}_{\mathrm{v}_{-1}}-i\mathbf{H}_{i\mathrm{v}_{-1}})$ and ¯$\bar{\mathcal{L}}_{-1}=\frac{i}{2}(\mathbf{H}_{\mathrm{v}_{1}}+i\mathbf{H}_{i\mathrm{v}_{1}})$ follow. ∎
2.3.3. The coordinate function $c:\mathcal{J}_{0,\infty}\to\mathbb{R}$
Proof of Proposition 2.9.
Let $\mathrm{v}=v(z)\partial_{z}\in\mathbb{C}(z)\partial_{z}$ and $\eta\in\mathcal{J}_{0,\infty}$. We are going to show that $c:\mathcal{J}_{0,\infty}\to\mathbb{R}$ is differentiable in the direction $\mathrm{v}$ at $\eta$ and compute the derivative. Let $\phi_{t}(z)=z+tv(z)+o(t)$ be the flow associated to $\mathrm{v}$, defined for small complex $t$, and let $\eta_{t}=\phi_{t}\cdot\eta$ be the perturbed loop. We will denote by $c,f,\tilde{c},g$ and $c_{t},f_{t},\tilde{c}_{t},g_{t}$ the coordinates of $\eta$ and $\eta_{t}$ respectively.
Since $\phi_{t}$ is conformal away from a small neighbourhood of 0, we have $e^{-\tilde{c}_{t}}g_{t}=\phi_{t}\circ(e^{-\tilde{c}}g)$. On the other hand, we can extend $\phi_{t}$ to a quasiconformal homeomorphism $\Phi_{t}$ of $\widehat{\mathbb{C}}$, generated by a Beltrami differential $\mu$ in the way specified by Lemma 2.3 with $r=\min\{\frac{1}{2},\frac{1}{8}e^{c}\}$. We can also impose that $\Phi_{t}$ fixes 0, so that $\Phi_{t}$ is normalised in the sense of Definition 2.1 (first item). Then, the map
$(e^{c_{t}}f_{t})^{-1}\circ\Phi_{t}\circ(e^{c}f)$
is conformal on an annular neighbourhood of $\mathbb{S}^{1}$ in $\mathbb{D}$ and fixes $\mathbb{S}^{1}$ setwise, i.e. it restricts to an analytic diffeomorphism of the unit circle. This map also fixes 0 by construction and extends to a quasiconformal homeomorphism of $\widehat{\mathbb{C}}$ by Schwarz reflection, which is analytic in the neighbourhood of $\mathbb{S}^{1}$. By symmetry, this extension also fixes $\infty$ and the Beltrami differential generating it satisfies the reflection property of Definition 2.1 (second item). So we have shown that there exists $\theta_{t}\in\mathbb{R}$ and a normalised quasiconformal homeomorphism $\tilde{\Phi}_{t}$ (in the sense of Definition 2.1, second item) such that
(2.10) $e^{c_{t}}f_{t}\circ(e^{i\theta_{t}}\tilde{\Phi}_{t})=\Phi_{t}\circ(e^{c}f).$
It is easy to evaluate the Beltrami differential generating $\tilde{\Phi}_{t}$. Indeed, the functions $f_{t}$ and $f$ are holomorphic in $\mathbb{D}$, so the chain rule gives
$\partial_{\bar{z}}\left((e^{c_{t}}f_{t})^{-1}\circ\Phi_{t}\circ(e^{c}f)\right)=((e^{c_{t}}f_{t})^{-1})^{\prime}\circ\Phi_{t}\circ(e^{c}f)\,\partial_{\bar{z}}\Phi_{t}\circ(e^{c}f)\,e^{c}\overline{f^{\prime}};$
$\partial_{z}\left((e^{c_{t}}f_{t})^{-1}\circ\Phi_{t}\circ(e^{c}f)\right)=((e^{c_{t}}f_{t})^{-1})^{\prime}\circ\Phi_{t}\circ(e^{c}f)\,\partial_{z}\Phi_{t}\circ(e^{c}f)\,e^{c}f^{\prime},$
so that
$\frac{\partial_{\bar{z}}\tilde{\Phi}_{t}}{\partial_{z}\tilde{\Phi}_{t}}=\frac{\partial_{\bar{z}}\Phi_{t}\circ(e^{c}f)}{\partial_{z}\Phi_{t}\circ(e^{c}f)}\frac{\overline{f^{\prime}}}{f^{\prime}}=\left(\frac{\partial_{\bar{z}}\Phi_{t}}{\partial_{z}\Phi_{t}}\right)\circ(e^{c}f)\frac{\overline{f^{\prime}}}{f}.$
Hence, the Beltrami differential generating $\tilde{\Phi}_{t}$ coincides with the pullback $(e^{c}f)^{*}\mu=\mu\circ(e^{c}f)\frac{\overline{f^{\prime}}}{f^{\prime}}$ in $\mathbb{D}$ (recall (2.2)), and is extended to the whole sphere by reflection.
Applying $\partial_{z}$ to both sides of (2.10) and specialising to 0 gives
(2.11) $e^{c_{t}+i\theta_{t}-c}=\frac{\partial_{z}\Phi_{t}(0)}{\partial_{z}\tilde{\Phi}_{t}(0)}.$
Now, we use the variational formulas of Proposition 2.2 and Lemma 2.3 in order to expand the RHS. Using that the $\partial_{z}$-derivative of the Cauchy transform is the Beurling transform [Ahl87, Chapter V, Lemma 3], we get
$\partial_{z}\Phi_{t}(z)=1-\frac{t}{\pi}\int_{\mathbb{D}}\frac{\mu(\zeta)}{(z-\zeta)^{2}}|\mathrm{d}\zeta|^{2}+o(t)$
$\partial_{z}\tilde{\Phi}_{t}(z)=1-\frac{t}{\pi}\int_{\mathbb{D}}\frac{(2z-1)(\zeta-z)+z(z-1)}{(\zeta-z)^{2}}\frac{(e^{c}f)^{*}\mu(z)}{\zeta(\zeta-1)}|\mathrm{d}\zeta|^{2}$
$\quad\qquad\qquad+\frac{\bar{t}}{\pi}\int_{\mathbb{D}}\frac{(2z-1)(1-z\bar{\zeta})+z(z-1)\bar{\zeta}}{(1-z\bar{\zeta})}\frac{\overline{(e^{c}f)^{*}\mu(z)}}{\bar{\zeta}(\bar{\zeta}-1)}|\mathrm{d}\zeta|^{2}+o(t),$
where the integrals are in the principal value sense. Evaluating these integrals at $z=0$ gives
$\partial_{z}\Phi_{t}(0)=1-\frac{t}{\pi}\int_{e^{c}f(\mathbb{D})}\mu(\zeta)\frac{|\mathrm{d}\zeta|^{2}}{\zeta^{2}}+o(t);$
$\partial_{z}\tilde{\Phi}_{t}(0)=1+\frac{t}{\pi}\int_{\mathbb{D}}\frac{(e^{c}f)^{*}\mu(\zeta)}{\zeta^{2}(\zeta-1)}|\mathrm{d}\zeta|^{2}-\frac{\bar{t}}{\pi}\int_{\mathbb{D}}\frac{\overline{(e^{c}f)^{*}\mu(z)}}{\bar{\zeta}(\bar{\zeta}-1)}|\mathrm{d}\zeta|^{2}+o(t).$
This shows that the RHS of (2.11) is differentiable in direction $\mathrm{v}$ and computes its differential at the same time. Thus, the LHS of (2.11) is also differentiable in direction $\mathrm{v}$. Expanding both sides of (2.11) at first order in $t$ gives:
$\mathcal{L}_{\mathrm{v}}(c+i\theta)=-\frac{1}{\pi}\int_{e^{c}f(\mathbb{D})}\mu(\zeta)\frac{|\mathrm{d}\zeta|^{2}}{\zeta^{2}}-\frac{1}{\pi}\int_{\mathbb{D}}\frac{(e^{c}f)^{*}\mu(\zeta)}{\zeta^{2}(\zeta-1)}|\mathrm{d}\zeta|^{2};\qquad\bar{\mathcal{L}}_{\mathrm{v}}(c+i\theta)=\frac{1}{\pi}\int_{\mathbb{D}}\frac{\overline{(e^{c}f)^{*}\mu(\zeta)}}{\bar{\zeta}(\bar{\zeta}-1)}|\mathrm{d}\zeta|^{2}.$
To conclude, we have
$\mathcal{L}_{\mathrm{v}}c$ $=\frac{1}{2}(\mathcal{L}_{\mathrm{v}}(c+i\theta)+\mathcal{L}_{\mathrm{v}}(c-i\theta))=\frac{1}{2}\left(\mathcal{L}_{\mathrm{v}}(c+i\theta)+\overline{\bar{\mathcal{L}}_{\mathrm{v}}(c+i\theta)}\right)$
$=-\frac{1}{2\pi}\int_{e^{c}f(\mathbb{D})}\mu(\zeta)\frac{|\mathrm{d}\zeta|^{2}}{\zeta^{2}}-\frac{1}{2\pi}\int_{\mathbb{D}}(e^{c}f)^{*}\mu(\zeta)\frac{|\mathrm{d}\zeta|^{2}}{\zeta^{2}(\zeta-1)}+\frac{1}{2\pi}\int_{\mathbb{D}}(e^{c}f)^{*}\mu(\zeta)\frac{|\mathrm{d}\zeta|^{2}}{\zeta(\zeta-1)}$
$=\frac{e^{-2c}}{2\pi}\int_{e^{c}f(\mathbb{D})}\left(\left(\frac{(f^{-1})^{\prime}(e^{-c}\zeta)}{f^{-1}(e^{-c}\zeta)}\right)^{2}-\frac{1}{\zeta^{2}}\right)\mu(\zeta)|\mathrm{d}\zeta|^{2}=\frac{e^{-c}}{4i\pi}\oint\left(\frac{(f^{-1})^{\prime}(\zeta)}{f^{-1}(\zeta)}\right)^{2}v(e^{c}\zeta)\mathrm{d}\zeta=\varpi(\mathrm{v}).$
In the penultimate equation, we used Stokes’ theorem and we noticed that the integral in $\frac{1}{\zeta^{2}}\mathrm{d}\zeta$ vanishes since $v$ as an expansion in negative powers of $z$. This concludes the proof. ∎
2.3.4. Existence of derivatives on the space $\mathcal{C}$
We now turn to the proof of Proposition 2.6:
Proof of Proposition 2.6.
Let $F\in\mathcal{C}$, $\mathrm{v}=v(z)\partial_{z}\in\mathbb{C}(z)\partial_{z}$ and $\eta\in\mathcal{J}_{0,\infty}$. We are going to show that $F$ is differentiable in the direction $\mathrm{v}$ at $\eta$. Let $\phi_{t}(z)=z+tv(z)+o(t)$ be the flow associated to $\mathrm{v}$, defined for small complex $t$, and let $\eta_{t}=\phi_{t}\cdot\eta$ be the perturbed loop. We will denote by $c,f,\tilde{c},g$ and $c_{t},f_{t},\tilde{c}_{t},g_{t}$ the coordinates of $\eta$ and $\eta_{t}$ respectively. It is enough to consider the case $\mathrm{v}=\mathrm{v}_{n}=-z^{n+1}\partial_{z}$ for some $n\in\mathbb{Z}$. We will treat the cases $n\geq-1$ and $n\leq-2$ separately, the second one being more involved than the first one.
Case $\mathrm{v}_{n}$, $n\geq-1$. We have computed $\mathcal{L}_{n}$ for $n\in\{-1,0\}$, so we will assume $n\geq 1$. Since $\phi_{t}$ is conformal in the neighbourhood of $e^{c}f(\bar{\mathbb{D}})$, we have
(2.12) $e^{c_{t}}f_{t}=\phi_{t}\circ(e^{c}f)=e^{c}(f-te^{nc}f^{n+1})+o(t).$
Let $F\in\mathcal{C}$ and $m_{0}\in\mathbb{N}^{*}$ such that $F$ depends only on $c,a_{1},...,a_{m_{0}}$. We may view $F=F(c,a_{1},...,a_{m_{0}})$ as a smooth function on ×$\mathbb{R}\times\mathbb{C}^{m_{0}}$. From (2.12), we see that $c_{t}=c+o(t)$ (due to $n\geq 1$), and $a_{m}(t)=a_{m}+te^{nc}P_{n,m}+o(t)$, where $P_{n,m}(f)=-\frac{1}{2i\pi}\oint f(z)^{n+1}\frac{\mathrm{d}z}{z^{m+2}}$ is a polynomial in $(a_{m})_{m\geq 1}$. Therefore, we have a (unique) first order expansion
$F(\phi_{t}\cdot\eta)=F(\eta)+t\mathcal{L}_{\mathrm{v}}F(\eta)+\bar{t}\bar{\mathcal{L}}_{\mathrm{v}}F(\eta)+o(t),$
where $\mathcal{L}_{n}=e^{nc}\sum_{m=1}^{\infty}P_{n,m}\partial_{a_{m}}$ and ¯$\bar{\mathcal{L}}_{n}=e^{nc}\sum_{m=1}^{\infty}\bar{P}_{n,m}\partial_{\bar{a}_{m}}$. These reduce to finite sums when acting on $\mathcal{C}$, so we get indeed operators $\mathcal{L}_{n},\bar{\mathcal{L}}_{n}$ preserving $\mathcal{C}$.
Case $\mathrm{v}_{-n},n\geq 2$. Now, we study the deformation $\phi_{t}$ generated by the vector field $\mathrm{v}=v\partial_{z}$ with $v\in\mathbb{C}[z^{-1}]\partial_{z}$. By the Koebe quarter theorem, the interior of $\eta(\mathbb{S}^{1})$ contains $\frac{e^{c}}{4}\mathbb{D}$. Hence, for small enough $t$, we have $\frac{e^{c}}{8}\mathbb{D}\subset\mathrm{int}(\phi_{t}\cdot\eta)$. Let $\mu$ be the associated Beltrami differential given by Lemma 2.3 with $r=\frac{e^{c}}{8}$. So $\Phi_{t}:=\Phi_{t\mu}$ extends $\phi_{t}$ quasiconformally to the whole sphere at first order. It is conformal in $\frac{e^{c}}{8}\mathbb{D}^{*}$ and $\Phi_{t}(z)=z+tv(z)+o(t)$ there. We keep denoting by $\mathrm{v}$ the vector field in $\widehat{\mathbb{C}}$ which generates $\Phi_{t}$. Recall that $\mathrm{v}=\bar{\partial}\mu$ and is given by the variational formula of Lemma 2.2.
Then, $(e^{c_{t}}f_{t})^{-1}\circ\Phi_{t}\circ(e^{c}f)$ is an analytic diffeomorphism of the unit circle, which we may extend to a quasiconformal homeomorphism $\tilde{\Phi}_{t}$ of $\widehat{\mathbb{C}}$. Since $f,f_{t},\Phi_{t}$ all fix 0, $\tilde{\Phi}_{t}$ also fixes 0. This homeomorphism is a solution to the Beltrami equation with the Beltrami differential $(e^{c}f)^{*}\mu$ in $\mathbb{D}$, extended to the whole sphere by reflection. So $\tilde{\Phi}_{t}$ also fixes $\infty$, and it is the normalised solution to the Beltrami equation up to a possible rotation by an angle $\theta_{t}=t\theta+o(t)$. This rotation only gives an additional contribution of $\theta\mathbf{H}_{i\mathrm{v}_{0}}$, so we may assume without loss of generality that $\theta=0$.
Let us write
$\tilde{\Phi}_{t}(z)=z+tw(z)+\bar{t}\tilde{w}(z)+o(t),$
where $w,\tilde{w}$ are given by the variational formulas of Lemma 2.2. By the chain rule, we have in the neighbourhood of $\mathbb{S}^{1}$
(2.13) $e^{c_{t}}f_{t}(z)=\Phi_{t}\circ(e^{c}f)\circ\tilde{\Phi}^{-1}_{t}(z)=e^{c}f(z)+tv(e^{c}f(z))-tw(z)e^{c}f^{\prime}(z)-\bar{t}\tilde{w}(z)e^{c}f^{\prime}(z)+o(t),$
where we have used that $\tilde{\Phi}_{t}^{-1}$ (resp. $\Phi_{t}$) is holomorphic in the neighbourhood of $\mathbb{S}^{1}$ (resp. $e^{c}f(\mathbb{S}^{1})$). By Proposition 2.9, we have $c_{t}=c+t\varpi(\mathrm{v})+\overline{t\varpi(\mathrm{v})}+o(t)$, so we need only focus on the other coefficients. By definition, we have
$a_{m}(t)=\frac{1}{2i\pi}\oint f_{t}(z)z^{-m-2}\mathrm{d}z,$
where we can pick the contour to be $r\mathbb{S}^{1}$ with $r<1$ such that $r\mathbb{D}$ contains the support of $f^{*}\mu$ in its interior. In the neighbourhood of this contour, the three small motions appearing in the expansion (2.13) are each holomorphic for $t$ small enough.
From now on, we specialise to the case $\mathrm{v}=\mathrm{v}_{-n}$ for some $n\geq 2$. Then, we have $\mu(\zeta)=n8^{2n}e^{-2nc}\bar{\zeta}^{n-1}\zeta\mathds{1}_{|\zeta|<\frac{e^{c}}{8}}$, and $(e^{c}f)^{*}\mu(\zeta)=n8^{2n}e^{-nc}\overline{f(\zeta)}^{n-1}f(\zeta)\frac{\overline{f^{\prime}(\zeta)}}{f^{\prime}(\zeta)}\mathds{1}_{|f(\zeta)|<\frac{1}{8}}$. This scaling relation and the explicit formulas for $v,w,\tilde{w}$ show that the RHS of (2.13) depend on $c$ only through a global factor of $e^{-nc}$. Therefore, we will assume $c=0$ in the remainder of the proof. On the one hand, it is easy to see that
$A_{n,m}:=-\frac{1}{2i\pi}\oint f(z)^{-n+1}z^{-m-2}\mathrm{d}z$
is a polynomial in $(a_{m})_{m\geq 1}$. From the explicit formula of $\mu$, we have
(2.14) $B_{n,m}$ $:=\oint w(z)f^{\prime}(z)z^{-m-2}\mathrm{d}z=\oint\int_{\mathbb{D}}\frac{z(z-1)f^{*}\mu(\zeta)}{\zeta(\zeta-1)(\zeta-z)}|\mathrm{d}\zeta|^{2}\frac{f^{\prime}(z)\mathrm{d}z}{z^{m+2}}$
$=n8^{2n}\oint\int_{f^{-1}(\frac{1}{8}\mathbb{D})}\sum_{k,l=0}^{\infty}\zeta^{k+l-1}z^{-k-2-m}(z-1)f^{\prime}(z)\overline{f(\zeta)}^{n-1}f(\zeta)\frac{\overline{f^{\prime}(\zeta)}}{f^{\prime}(\zeta)}|\mathrm{d}\zeta|^{2}\mathrm{d}z.$
By a change of variables and Stokes’ formula, we have for all $k,l\in\mathbb{N}$
(2.15) $\int_{f^{-1}(\frac{1}{8}\mathbb{D})}$ $\overline{f(\zeta)}^{n-1}f(\zeta)\frac{\overline{f^{\prime}(\zeta)}}{f^{\prime}(\zeta)}\zeta^{k+l-1}|\mathrm{d}\zeta|^{2}=\int_{\frac{1}{8}\mathbb{D}}(f^{-1})^{\prime}(\zeta)^{2}f^{-1}(\zeta)^{k+l-1}\bar{\zeta}^{n-1}\zeta|\mathrm{d}\zeta|^{2}$
$=\frac{1}{2in}\oint_{\frac{1}{8}\mathbb{S}^{1}}(f^{-1})^{\prime}(\zeta)^{2}f^{-1}(\zeta)^{k+l-1}\bar{\zeta}^{n}\zeta\mathrm{d}\zeta=\frac{8^{-2n}}{2in}\oint_{\frac{1}{8}\mathbb{S}^{1}}(f^{-1})^{\prime}(\zeta)^{2}f^{-1}(\zeta)^{k+l-1}\zeta^{-n+1}\mathrm{d}\zeta.$
This last expression shows that the integral is a polynomial in $(a_{m})_{m\geq 1}$ (with no dependence on $(\bar{a}_{m})_{m\geq 1}$), and it also vanishes for $k+l$ large enough (depending only on $n$). It follows that the series in (2.14) converges absolutely, so that we can exchange the sum and in the integral to get
$B_{n,m}$ $=\frac{1}{2i}\sum_{k,l=0}^{\infty}\oint f^{\prime}(z)(z-1)z^{-k-m-2}\mathrm{d}z\oint_{\frac{1}{8}\mathbb{S}^{1}}(f^{-1})^{\prime}(\zeta)^{2}f^{-1}(\zeta)^{k+l-1}\zeta^{-n+1}\mathrm{d}\zeta.$
For each fixed $k$, $\oint f^{\prime}(z)(z-1)z^{-k-m-1}\mathrm{d}z$ is a polynomial in the coefficients of $f$. Keeping in mind that the sum evaluates to zero for $k,l$ large enough, we get that $B_{n,m}\in\mathbb{C}[(a_{m})_{m\geq 1}]$.
Finally, we study
$C_{n,m}$ $:=\oint\tilde{w}(z)f^{\prime}(z)z^{-m-2}\mathrm{d}z=\oint\int_{\mathbb{D}}\frac{z(z-1)\overline{f^{*}\mu(\zeta)}}{\bar{\zeta}(\bar{\zeta}-1)(1-\bar{\zeta}z)}|\mathrm{d}\zeta|^{2}f^{\prime}(z)z^{-m-2}\mathrm{d}z$
$=-n\oint\int_{f^{-1}(\frac{1}{8}\mathbb{D})}\sum_{k,l=0}^{\infty}\bar{\zeta}^{k+l-1}z^{k-m-1}(z-1)f^{\prime}(z)f(\zeta)^{n-1}\overline{f(\zeta)}\frac{f^{\prime}(\zeta)}{\overline{f^{\prime}(\zeta)}}|\mathrm{d}\zeta|^{2}\mathrm{d}z.$
For each $k,l\in\mathbb{N}$, the integral over $\frac{e^{c}}{8}\mathbb{D}$ is the complex conjugate of (2.15), so it is a polynomial in $(\bar{a}_{m})_{m\geq 1)}$ and vanishes for $k+l$ large enough (depending only on $n$). It follows that
$C_{n,m}=\sum_{k,l=0}^{\infty}\oint f^{\prime}(z)(z-1)z^{k-m-1}\mathrm{d}z\int_{f^{-1}(\frac{1}{8}\mathbb{D})}f(\zeta)^{n-1}\overline{f(\zeta)}\frac{f^{\prime}(\zeta)}{\overline{f^{\prime}(\zeta)}}\bar{\zeta}^{k+l-1}|\mathrm{d}\zeta|^{2}.$
Each term in this (finite) sum is a polynomial in $(a_{m},\bar{a}_{m})$, so indeed $C_{n,m}$ is a polynomial. Contrary to $A_{n,m}$ and $B_{n,m}$, the polynomial $C_{n,m}$ is in both $(a_{m})$ and $(\bar{a}_{m})_{m\geq 1}$.
Putting things together, (2.13) implies that the operators $\mathcal{L}_{-n}$ and ¯$\bar{\mathcal{L}}_{-n}$ act on $\mathcal{C}$ as
(2.16) $\mathcal{L}_{-n}=\varpi(\mathrm{v}_{-n})\partial_{c}+e^{-nc}\sum_{m=1}^{\infty}A_{n,m}\partial_{a_{m}}-B_{n,m}\partial_{a_{m}}-\overline{C}_{n,m}\partial_{\bar{a}_{m}};$
¯$\bar{\mathcal{L}}_{-n}=\overline{\varpi(\mathrm{v}_{-n})}\partial_{c}+e^{-nc}\sum_{m=1}^{\infty}\overline{A}_{n,m}\partial_{\bar{a}_{m}}-\overline{B}_{n,m}\partial_{\bar{a}_{m}}-C_{n,m}\partial_{a_{m}}.$
As in the proof of the case dealing with $\mathrm{v}_{n},n\geq-1$, it follows from this analysis that $\mathcal{L}_{-n}$, ¯$\bar{\mathcal{L}}_{-n}$ are operators on $\mathcal{C}$ which preserve $\mathcal{C}$, and they are first order differential operators with polynomial coefficients in $(a_{m},\bar{a}_{m})$ and a global prefactor $e^{nc}$. These indeed define endomorphisms of $\mathcal{C}$ and $\mathcal{C}_{\mathrm{comp}}$.
Commutation relations. Now that we have constructed the operators $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$, $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, it remains to prove the commutation relations. These hold by construction since the operators $\mathcal{L}_{\mathrm{v}}$ are Lie derivatives in directions of fundamental vector fields, but we include a proof “by hand” since we work in an infinite dimensional setting. Let $\eta\in\mathcal{J}_{0,\infty}$. Let $\mathrm{v},\mathrm{w}\in\mathbb{C}(z)\partial_{z}$, and let $\phi_{t}(z)=z+tv(z)+o(t)$, $\psi_{s}(z)=z+sw(z)+o(s)$ be there flows, defined for small complex $t,s$ in the region $O_{\eta}$ (1.22). Let $F\in\mathcal{C}$. Since $\mathcal{L}_{\mathrm{v}}F,\mathcal{L}_{\mathrm{w}}F$ exist in $\mathcal{C}$, we have on the one hand,
$F(\psi_{s}\circ\phi_{t}\cdot\eta)$ $=F(\psi_{s}\cdot(\phi_{t}\cdot\eta))=F(\phi_{t}\cdot\eta)+s\mathcal{L}_{\mathrm{w}}F(\phi_{t}\cdot\eta)+\bar{s}\bar{\mathcal{L}}_{\mathrm{w}}F(\phi_{t}\cdot\eta)+o(s)$
$=F(\eta)+t\mathcal{L}_{\mathrm{v}}F(\eta)+\bar{t}\bar{\mathcal{L}}_{\mathrm{v}}F(\eta)+s\mathcal{L}_{\mathrm{w}}F(\eta)+\bar{s}\bar{\mathcal{L}}_{\mathrm{w}}F(\eta)$
$\quad+st\mathcal{L}_{\mathrm{v}}\mathcal{L}_{\mathrm{w}}F(\eta)+s\bar{t}\bar{\mathcal{L}}_{\mathrm{v}}\mathcal{L}_{\mathrm{w}}F(\eta)+\bar{s}t\mathcal{L}_{\mathrm{v}}\bar{\mathcal{L}}_{\mathrm{w}}F(\eta)+\bar{s}\bar{t}\bar{\mathcal{L}}_{\mathrm{v}}\bar{\mathcal{L}}_{\mathrm{w}}F(\eta)+o(ts).$
With a symmetric computation for $F(\phi_{t}\circ\psi_{s}\cdot\eta)$, we deduce that $F(\psi_{s}\circ\phi_{t}\cdot\eta)-F(\phi_{t}\circ\psi_{s}\cdot\eta)$ is equal to
(2.17) $2st[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]F(\eta)+2s\bar{t}[\bar{\mathcal{L}}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]F(\eta)+2\bar{s}t[\mathcal{L}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]F(\eta)+2\bar{s}\bar{t}[\bar{\mathcal{L}}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]F(\eta)+o(ts).$
On the other hand, we have $\psi_{s}\circ\phi_{t}(z)=z+tv(z)+sw(z)+ts(v(z)w^{\prime}(z)-v^{\prime}(z)w(z))+o(ts)$, so (recall $[v\partial_{z},w\partial_{z}]=(vw^{\prime}-v^{\prime}w)\partial_{z}$)
(2.18) $F(\psi_{s}\circ\phi_{t}\cdot\eta)-F(\phi_{t}\circ\psi_{s}\cdot\eta)=2ts\mathcal{L}_{[\mathrm{v},\mathrm{w}]}F(\eta)+2\bar{t}\bar{s}\bar{\mathcal{L}}_{[\mathrm{v},\mathrm{w}]}F(\eta)+o(ts).$
We have noted that the contribution at order 2 vanishes by antisymmetry. Now, identifying the terms in $st$, $\bar{s}\bar{t}$, $\bar{s}t$, $s\bar{t}$ in (2.17) and (2.18) gives
$[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]=\mathcal{L}_{[\mathrm{v},\mathrm{w}]};\qquad[\bar{\mathcal{L}}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]=\bar{\mathcal{L}}_{[\mathrm{v},\mathrm{w}]};\qquad[\mathcal{L}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]=[\bar{\mathcal{L}}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]=0.$
This finishes the proof of Proposition 2.6. ∎
2.4. Virasoro representation
In this section, we modify our representation of the Witt algebra to a representation of the Virasoro algebra. The same modification is considered in [GQW24, Section 4.1]. Recall the definition (1.3) of $\vartheta$ and $\tilde{\vartheta}$ and that, for all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, $\mathbf{L}_{\mathrm{v}}$ and $\bar{\mathbf{L}}_{\mathrm{v}}$ are defined from $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$ by:
$\mathbf{L}_{\mathrm{v}}:=\mathcal{L}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{v}),\qquad\bar{\mathbf{L}}_{\mathrm{v}}:=\bar{\mathcal{L}}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\overline{\vartheta(\mathrm{v})}.$
For ease of future record, we state the following lemma which, in conjunction with Proposition 2.6, implies that $\mathbf{L}_{\mathrm{v}}$ and $\bar{\mathbf{L}}_{\mathrm{v}}$ preserve the subspace $\mathcal{C}_{\mathrm{comp}}$.
Lemma 2.10.
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, the multiplication by $\vartheta(\mathrm{v})$ (resp. by $\tilde{\vartheta}(\mathrm{v})$) preserves $\mathcal{C}_{\mathrm{comp}}$ (2.6).
Proof.
We will consider $\tilde{\vartheta}(\mathrm{v})$. It is enough to treat the case $\mathrm{v}=\mathrm{v}_{n}$ for some $n\in\mathbb{Z}$. We have $\tilde{\vartheta}(\mathrm{v}_{n})=0$ for $n\leq-2$, so we assume $n\geq-1$. Recalling the definition (1.14) of $\tilde{c}_{0}$, we have by scaling that $\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v}_{n})=e^{-n(\tilde{c}_{0}-c)}\tilde{P}_{n}(g)$ for some polynomial $\tilde{P}_{n}$ in the coefficients of the expansion of $g$ at $\infty$. These coefficients are uniformly bounded by de Branges’ theorem, so $\tilde{P}_{n}$ is uniformly bounded. Moreover, by the Koebe quarter theorem, for each $F\in\mathcal{C}_{\mathrm{comp}}$, $\tilde{c}_{0}$ and $c$ are bounded on the support of $F$. Overall, the multiplication by $e^{-n(\tilde{c}_{0}-c)}\tilde{P}_{n}(g)$ preserves $\mathcal{C}_{\mathrm{comp}}$. ∎
Proposition 2.11.
As endomorphisms of $\mathcal{C}_{\mathrm{comp}}$, the following commutation relations are satisfied:
$[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=\mathbf{L}_{[\mathrm{v},\mathrm{w}]}+\frac{c_{\mathrm{m}}}{12}\omega(\mathrm{v},\mathrm{w});\qquad[\bar{\mathbf{L}}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]=\bar{\mathbf{L}}_{[\mathrm{v},\mathrm{w}]}+\frac{c_{\mathrm{m}}}{12}\overline{\omega(\mathrm{v},\mathrm{w})};\qquad[\mathbf{L}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]=0,$
for all $\mathrm{v},\mathrm{w}\in\mathbb{C}(z)\partial_{z}$, where $\omega(\mathrm{v},\mathrm{w})$ is the Virasoro cocycle (1.5).
Proof.
To prove the commutation relations of $(\mathbf{L}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$, it is enough to treat the three cases below by linearity. The first one will follow directly from Proposition 2.6. The second one will require an extra computation related to the Schwarzian derivative. The third one will follow from the first case and a duality argument based on the integration by parts formula of Corollary 4.2.
Case 1. $\mathrm{v},\mathrm{w}\in\mathbb{C}[z]\partial_{z}$. In this case $\mathbf{L}_{\mathrm{v}}=\mathcal{L}_{\mathrm{v}}$ and $\mathbf{L}_{\mathrm{w}}=\mathcal{L}_{\mathrm{w}}$ and by Proposition 2.6 we deduce that $[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]=\mathcal{L}_{[\mathrm{v},\mathrm{w}]}$. Since $[v,w]$ also belongs to $\mathbb{C}[z]\partial_{z}$, $\mathcal{L}_{[\mathrm{v},\mathrm{w}]}$ agrees with $\mathbf{L}_{[\mathrm{v},\mathrm{w}]}$ and we conclude that $[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=\mathbf{L}_{[\mathrm{v},\mathrm{w}]}$.
Case 2. $\mathrm{v}=v\partial_{z}\in z^{2}\mathbb{C}[z]\partial_{z}$ and $\mathrm{w}=w\partial_{z}\in z\mathbb{C}[z^{-1}]\partial_{z}$. In this case $\vartheta(\mathrm{v})=0$ and we only need to deal with $\vartheta(\mathrm{w})$. The latter can be seen as either an element of $\mathcal{C}_{\mathrm{comp}}$ or as an endomorphism of $\mathcal{C}_{\mathrm{comp}}$ acting by multiplication. To clarify, we will denote these two objects in this proof by $\vartheta(\mathrm{w})\in\mathcal{C}_{\mathrm{comp}}$ and $\vartheta(\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}$ respectively. We have
$\hskip 40.0pt\mathcal{L}_{\mathrm{v}}\circ(\vartheta(\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}})=(\mathcal{L}_{\mathrm{v}}\vartheta(\mathrm{w}))\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}+\vartheta(\mathrm{w})\mathcal{L}_{\mathrm{v}}\qquad\qquad\text{and so}$
$[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=\left[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}\right]=[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]-\frac{c_{\mathrm{m}}}{12}(\mathcal{L}_{\mathrm{v}}\vartheta(\mathrm{w}))\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}.$
By Proposition 2.6, we already know that $[\mathcal{L}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}]=\mathcal{L}_{[\mathrm{v},\mathrm{w}]}$. We are going to show that
(2.19) $\mathcal{L}_{\mathrm{v}}\vartheta(\mathrm{w})=-\omega(\mathrm{v},\mathrm{w})+\vartheta([\mathrm{v},\mathrm{w}]).$
We will then deduce that
$[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=\mathcal{L}_{[\mathrm{v},\mathrm{w}]}-\frac{c_{\mathrm{m}}}{12}\vartheta([\mathrm{v},\mathrm{w}])\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}+\frac{c_{\mathrm{m}}}{12}\omega(\mathrm{v},\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}=\mathbf{L}_{[\mathrm{v},\mathrm{w}]}+\frac{c_{\mathrm{m}}}{12}\omega(\mathrm{v},\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}$
as desired.
It remains to prove (2.19). To this end, let $\eta\in\mathcal{J}_{0,\infty}$ parametrised by $(c,f)\in\mathbb{R}\times\mathcal{E}$ and let $\phi_{t}(z)=z+tv(z)+o(t)$ be the flow generated by $\mathrm{v}$, defined for small $t$ in a neighbourhood $\overline{\mathrm{int}(\eta)}$. In the coordinates $(c,f)\in\mathbb{R}\times\mathcal{E}$, the motion $\phi_{t}\cdot\eta$ is represented by a normalised conformal map $e^{c_{t}}f_{t}$. Since $v^{\prime}(0)=0$, $\phi_{t}(e^{c}f)$ is normalised at first order and uniformises the interior of $\phi_{t}\cdot\eta$. Hence, $e^{c_{t}}f_{t}=\phi_{t}(e^{c}f)+o(t)$. By the chain rule for the Schwarzian, we have
$\mathcal{S}((\phi_{t}\circ(e^{c}f))^{-1})=\mathcal{S}((e^{c}f)^{-1}\circ\phi_{t}^{-1})$ $=\mathcal{S}(\phi_{t}^{-1})+((\phi_{t}^{-1})^{\prime})^{2}\mathcal{S}((e^{c}f)^{-1})\circ\phi_{t}^{-1}$
(2.20) $=\mathcal{S}((e^{c}f)^{-1})-t\left(v^{\prime\prime\prime}+2v^{\prime}\mathcal{S}((e^{c}f)^{-1})+v\mathcal{S}((e^{c}f)^{-1})^{\prime}\right)+o(t).$
By integration by parts, $\mathcal{L}_{\mathrm{v}}\vartheta_{e^{c}f}(\mathrm{w})$ is thus equal to
$\underset{t\to 0}{\lim}\,\frac{1}{2i\pi t}\oint\left(\mathcal{S}((\phi_{t}\circ(e^{c}f))^{-1}(z)-\mathcal{S}(e^{c}f)^{-1}(z)\right)w(z)\mathrm{d}z$
$=-\frac{1}{2i\pi}\oint v^{\prime\prime\prime}(z)w(z)\mathrm{d}z-\frac{1}{2i\pi}\oint\left(2v^{\prime}(z)w(z)\mathcal{S}((e^{c}f)^{-1})(z)+v(z)w(z)\mathcal{S}((e^{c}f)^{-1})^{\prime}(z)\right)\mathrm{d}z$
$=-\omega(\mathrm{v},\mathrm{w})+\frac{1}{2i\pi}\oint\left(v(z)w^{\prime}(z)-v^{\prime}(z)w(z)\right)\mathcal{S}((e^{c}f)^{-1})(z)\mathrm{d}z=-\omega(\mathrm{v},\mathrm{w})+\vartheta_{e^{c}f}([\mathrm{v},\mathrm{w}])$
which concludes the proof of (2.19) and thus the proof of Case 2.
Case 3. $\mathrm{v},\mathrm{w}\in z\mathbb{C}[z^{-1}]\partial_{z}$. Recall that we define an antilinear involution of $\mathbb{C}(z)\partial_{z}$ by
(2.21) $\mathrm{v}^{*}=z^{2}\bar{v}(1/z)\partial_{z}.$
On the generators, we have $\mathrm{v}_{n}^{*}=\mathrm{v}_{-n}$. Moreover, for all $n,m\in\mathbb{Z}$,
$[\mathrm{v}_{n},\mathrm{v}_{m}]^{*}=(n-m)\mathrm{v}_{n+m}^{*}=(n-m)\mathrm{v}_{-n-m}=-[\mathrm{v}_{-n},\mathrm{v}_{-m}]=-[\mathrm{v}_{n}^{*},\mathrm{v}_{m}^{*}],$
so that for all $\mathrm{v},\mathrm{w}\in\mathbb{C}(z)\partial_{z}$,
(2.22) $[\mathrm{v},\mathrm{w}]^{*}=-[\mathrm{v}^{*},\mathrm{w}^{*}].$
Let us fix $\mathrm{v},\mathrm{w}\in z\mathbb{C}[z^{-1}]\partial_{z}$. The integration by parts formula of Corollary 4.2 (which is independent of the commutation relations) gives:
(2.23) $[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=[\Theta\circ\mathbf{L}_{\mathrm{v}^{*}}^{*}\circ\Theta,\Theta\circ\mathbf{L}_{\mathrm{w}^{*}}^{*}\circ\Theta]=\Theta\circ[\mathbf{L}_{\mathrm{v}^{*}}^{*},\mathbf{L}_{\mathrm{w}^{*}}^{*}]\circ\Theta=-\Theta\circ[\mathbf{L}_{\mathrm{v}^{*}},\mathbf{L}_{\mathrm{w}^{*}}]^{*}\circ\Theta.$
Now, since $\mathrm{v}^{*},\mathrm{w}^{*}\in z\mathbb{C}[z]\partial_{z}$, we know from Case 1 that $[\mathbf{L}_{\mathrm{v}^{*}},\mathbf{L}_{\mathrm{w}^{*}}]=\mathbf{L}_{[\mathrm{v}^{*},\mathrm{w}^{*}]}$ ($\omega(\mathrm{v}^{*},\mathrm{w}^{*})$ vanishes in this case). Using once more the integration by parts formula, we deduce that
$[\mathbf{L}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=-\Theta\circ\mathbf{L}_{[\mathrm{v}^{*},\mathrm{w}^{*}]}^{*}\circ\Theta=-\mathbf{L}_{[\mathrm{v}^{*},\mathrm{w}^{*}]^{*}}=\mathbf{L}_{[\mathrm{v},\mathrm{w}]}.$
In the last equality we used that $[\mathrm{v}^{*},\mathrm{w}^{*}]^{*}=-[\mathrm{v},\mathrm{w}]$; see (2.22). This is the desired result since $\omega(\mathrm{v},\mathrm{w})=0$ in this case.
This proves the commutation relations for the representation $(\mathbf{L}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$. The proof for the representation $(\bar{\mathbf{L}}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ is identical. The last step is to prove that the two representations commute. If $\mathrm{v},\mathrm{w}\in\mathbb{C}[z]\partial_{z}$, we have $[\mathbf{L}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]=[\mathcal{L}_{\mathrm{v}},\bar{\mathcal{L}}_{\mathrm{w}}]$ which vanishes by Proposition 2.6. In this case, we also have
(2.24) $[\mathbf{L}_{\mathrm{v}^{*}},\bar{\mathbf{L}}_{\mathrm{w}^{*}}]=[\Theta\circ\mathbf{L}_{\mathrm{v}}^{*}\circ\Theta,\Theta\circ\bar{\mathbf{L}}_{\mathrm{w}}^{*}\circ\Theta]=\Theta\circ[\mathbf{L}_{\mathrm{v}}^{*},\bar{\mathbf{L}}_{\mathrm{w}}^{*}]\circ\Theta=-\Theta\circ[\mathbf{L}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]^{*}\circ\Theta=0.$
Finally, suppose $\mathrm{v}\in z^{2}\mathbb{C}[z]\partial_{z}$ and $\mathrm{w}\in z\mathbb{C}[z^{-1}]\partial_{z}$. Then, $[\bar{\mathbf{L}}_{\mathrm{v}},\mathbf{L}_{\mathrm{w}}]=[\bar{\mathcal{L}}_{\mathrm{v}},\mathcal{L}_{\mathrm{w}}-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}]=-\frac{c_{\mathrm{m}}}{12}\bar{\mathcal{L}}_{\mathrm{v}}\vartheta(\mathrm{w})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}$, which vanishes by analyticity of the Schwarzian on $\mathcal{E}$ (there is no $\bar{t}$ term in (2.20)). Similarly, we get that $[\mathbf{L}_{\mathrm{v}},\bar{\mathbf{L}}_{\mathrm{w}}]=0$ for $\mathrm{v}\in z\mathbb{C}[z^{-1}]\partial_{z}$ and $\mathrm{w}\in z^{2}\mathbb{C}[z]\partial_{z}$, so we are done. ∎
Remark 2.12.
We stress that the commutation relations (2.23) and (2.24) would be difficult to obtain without the integration by parts formula of Corollary 4.2.
3. Highest-weight representations
In the previous section, we defined two representation of the Virasoro algebra on $L^{2}(\nu)$. In this section, we do a Laplace transform in the ($\log$) conformal radius, which will allow us to define a family of representations $(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})$ acting on $L^{2}(\nu^{\#})$, indexed by $\lambda\in\mathbb{C}$. The relevance of this family is that it allows us to construct highest-weight representations of weight $\lambda$. The main result of the section is Theorem 1.4 giving a complete classification of these representations.
3.1. Construction
The main idea is that the eigenstates of $\partial_{c}$ will label highest-weight states for the highest-weight representations. We could view the diagonalisation of this operator as an analytic question (consisting in a Fourier/Laplace transform), but since we have in mind algebraic statements (commutations relations etc.) we prefer to take it as such. Namely, we will introduce some operators having the property to conjugate $\partial_{c}$ with a multiplication operator.
Recall that the left and right levels of a polynomial in $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ have been defined below (1.10). Let $\mathbf{N},\overline{\mathbf{N}}$ be the endomorphisms of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ characterised by
$\mathbf{N}(\mathbf{a}^{\mathbf{k}}\bar{\mathbf{a}}^{\tilde{\mathbf{k}}})=|\mathbf{k}|\mathbf{a}^{\mathbf{k}}\bar{\mathbf{a}}^{\tilde{\mathbf{k}}},\qquad\overline{\mathbf{N}}(\mathbf{a}^{\mathbf{k}}\bar{\mathbf{a}}^{\tilde{\mathbf{k}}})=|\tilde{\mathbf{k}}|\mathbf{a}^{\mathbf{k}}\bar{\mathbf{a}}^{\tilde{\mathbf{k}}},\qquad\forall\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}.$
We call these operators the degree operators. They commute and are diagonal with spectrum $\mathbb{N}$. The $N$-eigenspace of $\mathbf{N}$ (resp. $\overline{\mathbf{N}}$) is the space of polynomials with left (resp. right) degree equal to $N$. Observe that $\mathbf{N}=\sum_{m=1}^{\infty}ma_{m}\partial_{a_{m}}$ and $\overline{\mathbf{N}}=\sum_{m=1}^{\infty}m\bar{a}_{m}\partial_{\bar{a}_{m}}$, so that by Proposition 2.8,
$\mathcal{L}_{0}=-\frac{\partial_{c}}{2}+\frac{1}{2}(\mathbf{N}-\overline{\mathbf{N}});\qquad\bar{\mathcal{L}}_{0}=-\frac{\partial_{c}}{2}-\frac{1}{2}(\mathbf{N}-\overline{\mathbf{N}}).$
We have seen in the previous section that $\mathbf{L}_{n}$ is a first order differential operator whose coefficients are polynomials in $a_{m},\bar{a}_{m}$, and $e^{c}$. The first step is to refine this result by giving the degrees of these polynomials. The next lemma uses the standard notation $x_{-}=\max\{-x,0\}$.
Lemma 3.1.
For all $n\in\mathbb{Z}$, we have
$\mathbf{L}_{n}=e^{nc}P_{n,0}\partial_{c}+e^{nc}Q_{n,0}+e^{nc}\sum_{m=1}^{\infty}(P_{n,m}\partial_{a_{m}}+Q_{n,m}\partial_{\bar{a}_{m}}),$
$\bar{\mathbf{L}}_{n}=e^{nc}\tilde{P}_{n,0}\partial_{c}+e^{nc}\tilde{Q}_{n,0}+e^{nc}\sum_{m=1}^{\infty}(\tilde{P}_{n,m}\partial_{\bar{a}_{m}}+\tilde{Q}_{n,m}\partial_{a_{m}}),$
where, for all $m\in\mathbb{Z}$ we have
  • $\bullet$
    $P_{n,m}$ (resp. $\tilde{P}_{n,m}$ has degree $((n-m)_{-},0)$ (resp. $(0,(n-m)_{-})$), for all $n\in\mathbb{Z}$.
  • $\bullet$
    If $n\leq 0$, $Q_{n,m}$ (resp. $\tilde{Q}_{n,m}$) is a linear combination of monomials of degree $(k,\tilde{k})$ with $k-\tilde{k}=(n+m)_{-}$ (resp. $\tilde{k}-k=(n+m)_{-}$). If $n>0$, $Q_{n,m}=\tilde{Q}_{n,m}=0$.
Moreover, recalling the definitions (1.3) of $\vartheta$ and (2.8) of $\varpi$, we have for all $n\in\mathbb{Z}$,
$e^{nc}P_{n,0}=\varpi(\mathrm{v}_{n});\qquad e^{nc}Q_{n,0}=-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{v}_{n});\qquad e^{nc}\tilde{P}_{n,0}=\overline{\varpi(\mathrm{v}_{n})};\qquad e^{nc}\tilde{Q}_{n,0}=-\frac{c_{\mathrm{m}}}{12}\overline{\vartheta(\mathrm{v}_{n})}.$
Remark 3.2.
At this stage, we are not saying anything about the total degree of $Q_{n,m}$ and $\tilde{Q}_{n,m}$. However, a consequence of the proof of Theorem 1.4 is that this degree is bounded by $m-n$. For instance, the reader can check from Proposition 2.8 that $Q_{-1,1}=|a_{1}|^{2}-1$, a non-homogeneous polynomial of total degree 2.
Proof.
The general form of $\mathcal{L}_{n}$ and ¯$\bar{\mathcal{L}}_{n}$ has been given in (2.16), and gives the value of $P_{n,0},\tilde{P}_{n,0}$. The values of $Q_{n,0},\tilde{Q}_{n,0}$ also follow by definition of $\mathbf{L}_{n},\bar{\mathbf{L}}_{n}$. It just remains to prove the statements on the degrees. We only treat the case of $\mathbf{L}_{n}$, the case of $\bar{\mathbf{L}}_{n}$ being identical. The degrees of $P_{n,0}$ and $Q_{n,0}$ are clear by definition of $\varpi$ and $\vartheta$. The proof of Proposition 2.6 also shows that the right degree of $P_{n,m}$ vanishes, and $Q_{n,m}=0$ for $n>0$.
We turn to the left degree of $P_{n,m}$ for $m\geq 1$. We have $P_{n,m}=e^{-nc}\mathbf{L}_{n}(a_{m})-Q_{n,0}a_{m}$. Applying $\bar{\mathbf{L}}_{0}=-\frac{1}{2}(\partial_{c}+\mathbf{N}-\overline{\mathbf{N}})$ on both sides and using $[\bar{\mathbf{L}}_{0},\mathbf{L}_{n}]=0$, we find $\mathbf{N}P_{n,m}=-(n-m)P_{n,m}$, i.e. $P_{n,m}$ has left degree $(n-m)_{-}$.
Finally, we study $Q_{n,m}$ for $n\leq 0$ and $m\geq 1$. We have $Q_{n,m}=e^{-nc}\mathbf{L}_{n}(\bar{a}_{m})-Q_{n,0}\bar{a}_{m}$. Applying again $\bar{\mathbf{L}}_{0}$ to both sides, we find $(\mathbf{N}-\overline{\mathbf{N}})Q_{n,m}=-(n+m)Q_{n,m}$. Hence, $Q_{n,m}$ is a combination of monomials of degree $(k,\tilde{k})$ with $k-\tilde{k}=(n+m)_{-}$. ∎
Since $\partial_{c}$ and the multiplication by $e^{nc}$ are continuous on $\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$, they extend to bounded operators on $\mathcal{S}^{\prime}_{\mathrm{exp}}(\mathbb{R})$ by duality. It is then clear from the expressions of the previous lemma that the operators $\mathbf{L}_{n},\bar{\mathbf{L}}_{n}$ extend to endomorphisms of $\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$, and this extension is unique by density of $\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$ in $\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})$. The level operators can be exponentiated on $\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$: the operator $e^{-\mathbf{N}c}$ acts by multiplication by $e^{-Nc}$ on the space of polynomials with left level $N$. This defines an endomorphism of $\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$.
Let $\lambda\in\mathbb{C}$. The previous lemma gives
(3.1) $\mathbf{L}_{n}(e^{-2\lambda c})=\bar{\mathbf{L}}_{n}(e^{-2\lambda c})=0\quad\text{for all }n\geq 1;\qquad\mathbf{L}_{0}(e^{-2\lambda c})=\bar{\mathbf{L}}_{0}(e^{-2\lambda c})=\lambda e^{-2\lambda c},$
and similarly for the conjugate representation. By definition, $e^{-2\lambda c}$ is a highest-weight state of weight $\lambda$. We will use this to construct highest-weight representations, but first we want to get rid of the exponential prefactors in $c$ in order to view all the states in the space of polynomials. To this end, given $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$, we define
(3.2) $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}:=e^{(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}),$
and set $\Psi_{\lambda,\mathbf{k}}:=\Psi_{\lambda,\mathbf{k},\eset}$, $\Psi_{\lambda}:=\Psi_{\lambda,\eset,\eset}=\mathds{1}$. By Lemma 3.1, we have $\Psi_{\lambda,\mathbf{k}}\in\mathbb{C}[(a_{m})_{m\geq 1}]$ and $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}\in\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$. Next, define
$\mathcal{V}_{\lambda}:=\mathrm{span}\left\{\Psi_{\lambda,\mathbf{k}}\big{|}\mathbf{k}\in\mathcal{T}\right\};\qquad\mathcal{W}_{\lambda}:=\mathrm{span}\left\{\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}\Big{|}\,\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}\right\}.$
We can define the diagonal endomorphisms $\mathbf{L}_{0}^{\lambda},\bar{\mathbf{L}}_{0}^{\lambda}$ of $\mathcal{W}_{\lambda}$ characterised by
(3.3) $\mathbf{L}_{0}^{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=(\lambda+|\mathbf{k}|)\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}};\qquad\bar{\mathbf{L}}_{0}^{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=(\lambda+|\tilde{\mathbf{k}}|)\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}.$
While the states $(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}})_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}$ may not be linearly independent, linear relations may only exist at the same $\lambda$-level by standard representation theory (see Appendix A), i.e. for $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$ and $\Psi_{\lambda,\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}$ with $(|\mathbf{k}|,|\tilde{\mathbf{k}}|)=(|\mathbf{k}^{\prime}|,|\tilde{\mathbf{k}}^{\prime}|)$; so these operators are indeed well defined.
Next, for all $n\in\mathbb{Z}$, we define the endomorphisms $\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda}$ of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ by
(3.4) $\mathbf{L}_{n}^{\lambda}:=-P_{n,0}(\mathbf{L}_{0}^{\lambda}+\bar{\mathbf{L}}_{0}^{\lambda})+Q_{n,0}+\sum_{m=1}^{\infty}(P_{n,m}\partial_{a_{m}}+Q_{n,m}\partial_{\bar{a}_{m}})$
$\bar{\mathbf{L}}_{n}^{\lambda}:=-\tilde{P}_{n,0}(\mathbf{L}_{0}^{\lambda}+\bar{\mathbf{L}}_{0}^{\lambda})+\tilde{Q}_{n,0}+\sum_{m=1}^{\infty}(\tilde{P}_{n,m}\partial_{\bar{a}_{m}}+\tilde{Q}_{n,m}\partial_{a_{m}}).$
Concretely, these operators are just obtained from $\mathbf{L}_{n},\bar{\mathbf{L}}_{n}$ by removing the prefactor $e^{nc}$, and replacing $\partial_{c}$ by the operator $-(\mathbf{L}_{0}^{\lambda}+\bar{\mathbf{L}}_{0}^{\lambda})$. These operators exponentiate to a linear map
(3.5) $\mathcal{P}_{\lambda}:=e^{-(\mathbf{L}_{0}^{\lambda}+\bar{\mathbf{L}}_{0}^{\lambda})c}:\mathcal{W}_{\lambda}\to\mathcal{S}_{\mathrm{exp}}^{\prime}(\mathbb{R})\otimes\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}].$
The next proposition essentially shows that $\mathcal{P}_{\lambda}$ conjugates the operators $\mathbf{L}_{n}^{\lambda}$ and $\mathbf{L}_{n}$.
Proposition 3.3.
For all $n\in\mathbb{Z}$, $\mathbf{L}_{n}^{\lambda}=\mathcal{P}_{\lambda}^{-1}\circ\mathbf{L}_{n}\circ\mathcal{P}_{\lambda}$ and $\bar{\mathbf{L}}_{n}^{\lambda}=\mathcal{P}_{\lambda}^{-1}\circ\bar{\mathbf{L}}_{n}\circ\mathcal{P}_{\lambda}$. In particular, the following commutations relations of endomorphisms of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ holds: for all $n,m\in\mathbb{Z}$,
$[\mathbf{L}_{n}^{\lambda},\mathbf{L}_{m}^{\lambda}]=(n-m)\mathbf{L}_{n+m}^{\lambda}+\frac{c_{\mathrm{m}}}{12}(n^{3}-n)\delta_{n,-m};$
$[\bar{\mathbf{L}}_{n}^{\lambda},\bar{\mathbf{L}}_{m}^{\lambda}]=(n-m)\bar{\mathbf{L}}_{n+m}^{\lambda}+\frac{c_{\mathrm{m}}}{12}(n^{3}-n)\delta_{n,-m};\qquad[\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{m}^{\lambda}]=0.$
Moreover, for all $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$, we have $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=\mathbf{L}_{-\mathbf{k}}^{\lambda}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}^{\lambda}\mathds{1}$. In particular, $\mathbf{L}_{0}^{\lambda}\mathds{1}=\bar{\mathbf{L}}_{0}^{\lambda}\mathds{1}=\lambda\mathds{1}$ and $\mathbf{L}_{n}^{\lambda}\mathds{1}=\bar{\mathbf{L}}_{n}^{\lambda}\mathds{1}=0$ for all $n\geq 1$.
Proof.
Let $n\in\mathbb{Z}$. We verify that $\mathbf{L}_{n}^{\lambda}=\mathcal{P}_{\lambda}^{-1}\circ\mathbf{L}_{n}\circ\mathcal{P}_{\lambda}$ on the generators $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$ of $\mathcal{W}_{\lambda}$. Let $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$. We have
(3.6) $\hskip 50.0pt\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\hskip 60.0pt\text{so}$
$e^{nc}P_{n,0}\partial_{c}(\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}})=-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|-n)c}P_{n,0}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}.$
From this, we get that $\mathbf{L}_{n}(\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}})$ is equal to
$e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|-n)c}\left(-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)P_{n,0}+Q_{n,0}+\sum_{m=1}^{\infty}(P_{n,m}\partial_{a_{m}}+Q_{n,m}\partial_{\bar{a}_{m}})\right)\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$
$=e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|-n)c}\mathbf{L}_{n}^{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=\mathcal{P}_{\lambda}(\mathbf{L}_{n}^{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}).$
The operator $\mathcal{P}_{\lambda}$ is clearly injective, so it has an inverse $\mathcal{P}_{\lambda}^{-1}$ defined on its image, and we have $\mathbf{L}_{n}^{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=\mathcal{P}_{\lambda}^{-1}\circ\mathbf{L}_{n}\circ\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$ as desired. From this, we deduce that the operators $(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})_{n\in\mathbb{Z}}$ satisfy the same commutation relations as the $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$. To check that $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=\mathbf{L}_{-\mathbf{k}}^{\lambda}\bar{\mathbf{L}}^{\lambda}_{-\tilde{\mathbf{k}}}\mathds{1}$, recall first that, by (3.6) applied to $\mathbf{k}=\tilde{\mathbf{k}}=\varnothing$, $\mathcal{P}^{\lambda}\mathds{1}=e^{-2\lambda c}$. From the conjugation relations, we deduce that $\mathbf{L}_{-\mathbf{k}}^{\lambda}\bar{\mathbf{L}}^{\lambda}_{-\tilde{\mathbf{k}}}\mathds{1}=\mathcal{P}_{\lambda}^{-1}\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}e^{-2\lambda c}$. By definition (3.2) of $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$, we then have $\mathbf{L}_{-\mathbf{k}}^{\lambda}\bar{\mathbf{L}}^{\lambda}_{-\tilde{\mathbf{k}}}\mathds{1}=\mathcal{P}_{\lambda}^{-1}e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$ which agrees with $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}$ by (3.6). ∎
3.2. Classification
In the previous section, we defined highest-weight representations $(\mathcal{V}_{\lambda},(\mathbf{L}_{n}^{\lambda})_{n\in\mathbb{Z}}$ and $(\mathcal{W}_{\lambda},(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})_{n\in\mathbb{Z}})$ of weight $\lambda\in\mathbb{C}$. In this section, we characterise the structure of these representations, i.e. we prove Theorem 1.4. This classification shows that the modules $\mathcal{V}_{\lambda}$ are all irreducible, i.e. there are no singular vectors at positive level. Interestingly, we can give two different proofs of this result: a purely algebraic proof and an analytic proof. The existence of an analytic proof relies on the fact that the Virasoro algebra is represented as differential operators on $\mathcal{J}_{0,\infty}$: roughly speaking, a singular vector is a function on $\mathcal{J}_{0,\infty}$ (rather, a section of a line bundle over $\mathcal{E}$) with vanishing differential, so it must be a constant.
Algebraic proof of Theorem 1.4.
Let $M_{c_{\mathrm{m}},\lambda}$ be the (abstract) Verma module of central charge $c_{\mathrm{m}}$ and $V_{c_{\mathrm{m}},\lambda}=M_{c_{\mathrm{m}},\lambda}/I_{c_{\mathrm{m}},\lambda}$ the irreducible quotient of $M_{c_{\mathrm{m}},\lambda}$ by the maximal proper submodule $I_{c_{\mathrm{m}},\lambda}$. We refer the reader to Appendix A for some details on these notions.
If $\lambda\not\in kac$, the Verma module is irreducible, i.e. $I_{c_{\mathrm{m}},\lambda}=\{0\}$. Since $\mathcal{V}_{\lambda}$ is a quotient of the Verma module by a submodule, we have necessarily $\mathcal{V}_{\lambda}\simeq M_{c_{\mathrm{m}},\lambda}$ as highest-weight representations. Hence $\dim\mathcal{V}_{\lambda}^{N}=\#\mathcal{T}_{N}$. On the other hand, $\mathcal{V}_{\lambda}^{N}$ is contained in the space of polynomials at level $(N,0)$, which has dimension #$\#\mathcal{T}_{N}$. We deduce by dimension counting that $\mathcal{V}_{\lambda}^{N}$ is exactly the space of polynomials at level $(N,0)$.
We turn to Item 2. We claim that $\mathbb{C}[(a_{m})_{m\geq 1}]$ equipped with the representation $(\mathbf{L}_{n}^{\lambda})_{n\in\mathbb{Z}}$ is isomorphic to the dual Verma module $M_{c_{\mathrm{m}},\lambda}^{\vee}$ (the space of linear forms on $M_{c_{\mathrm{m}},\lambda}$). Denote by $(e_{\mathbf{k}}=L_{-\mathbf{k}}\cdot e_{\eset})_{\mathbf{k}\in\mathcal{T}}$ a linear basis of $M_{c_{\mathrm{m}},\lambda}$. For all $\mathbf{k},\mathbf{k}^{\prime}\in\mathcal{T}$ with $|\mathbf{k}|=|\mathbf{k}^{\prime}|$, the polynomial $\mathbf{L}_{\mathbf{k}}^{\lambda}(\mathbf{a}^{\mathbf{k}^{\prime}})$ has level 0, i.e. it is proportional to the constant function $\mathds{1}$, and we identify it with this constant. This allows us to define a linear map $\Phi_{\lambda}:\mathbb{C}[(a_{m})_{m\geq 1}]\to M_{c_{\mathrm{m}},\lambda}^{\vee}$ by (recall the notation (1.11))
$\Phi_{\lambda}\mathbf{a}^{\mathbf{k}}(e_{\mathbf{k}^{\prime}}):=\left\{\begin{aligned} &\mathbf{L}_{\mathbf{k}^{\prime}}(\mathbf{a}^{\mathbf{k}}),&&\text{if}|\mathbf{k}|=|\mathbf{k}^{\prime}|,\\ &0,&&\text{otherwise}.\end{aligned}\right.$
This linear map preserves the Virasoro representations by definition. In particular, it preserves the level. Thus, to prove that $\Phi_{\lambda}$ is a linear isomorphism, it suffices to prove that it is injective on each level. Let $\mathbf{k}=(k_{m})_{m\geq 1}\in\mathcal{T}$ and $m_{0}:=\max\{m\geq 1|\,k_{m}>0\}$. By Lemma 3.1 and its proof, for $n>0$, we have $\mathbf{L}_{n}^{\lambda}=-\partial_{a_{n}}+\cdots$, where the dots contain only differential operators in $a_{m}$ with $m>n$. Hence,
$\Phi_{\lambda}\mathbf{a}^{\mathbf{k}}(e_{\mathbf{k}})=(\mathbf{L}_{1}^{\lambda})^{k_{1}}(\mathbf{L}_{2}^{\lambda})^{k_{2}}\cdots(\mathbf{L}_{m_{0}}^{\lambda})^{k_{m_{0}}}\mathbf{a}^{\mathbf{k}}=\prod_{m=1}^{m_{0}}(-1)^{k_{m}}k_{m}!\neq 0.$
This proves that $\Phi_{\lambda}$ is injective, and we deduce that $(\mathbb{C}[(a_{m})_{m\geq 1}],(\mathbf{L}_{n}^{\lambda})_{n\in\mathbb{Z}})\simeq M_{c_{\mathrm{m}},\lambda}^{\vee}$ as required. We conclude from Lemma A.1 that $\mathcal{V}_{\lambda}=\mathrm{span}\{\mathbf{L}_{-\mathbf{k}}^{\lambda}\mathds{1}|\,\mathbf{k}\in\mathcal{T}\}$ is isomorphic to the irreducible quotient $V_{c_{\mathrm{m}},\lambda}$.
Finally, we turn to Item 3. The fact that $\mathcal{W}_{\lambda}$ is the tensor product of two copies of $\mathcal{V}_{\lambda}$ is obvious since the two representations commute. In the sequel, we assume $\lambda\not\in kac$. By dimension counting, we just need to show that $\oplus_{n+\tilde{n}\leq N}\mathcal{W}_{\lambda}^{n,\tilde{n}}$ is a subspace of the space of polynomials of total degree bounded by $N$, for all $N\in\mathbb{N}$. Let $\mathbf{k}_{0},\tilde{\mathbf{k}}_{0}\in\mathcal{T}$ with $|\mathbf{k}_{0}|+|\tilde{\mathbf{k}}_{0}|=N$. On the one hand, the Shapovalov form is non-degenerate, so there exists $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$ with $|\mathbf{k}|+|\tilde{\mathbf{k}}|=N$ and $\mathbf{L}^{\lambda}_{\mathbf{k}}\mathbf{L}^{\lambda}_{\tilde{\mathbf{k}}}\Psi_{\lambda,\mathbf{k}_{0},\tilde{\mathbf{k}}_{0}}\neq 0$, i.e. this polynomial is a non-zero constant. On the other hand, Lemma 3.1 shows that $\mathbf{L}_{\mathbf{k}}^{\lambda}\bar{\mathbf{L}}_{\tilde{\mathbf{k}}}^{\lambda}$ decreases the degree of $\Psi_{\lambda,\mathbf{k}_{0},\tilde{\mathbf{k}}_{0}}$ by $|\mathbf{k}|+|\tilde{\mathbf{k}}|=N$. Hence, the total degree of $\Psi_{\lambda,\mathbf{k}_{0},\tilde{\mathbf{k}}_{0}}$ is exactly $N$. ∎
Analytic proof of Theorem 1.4.
Let $P\in\mathbb{C}[(a_{m})_{m\geq 1}]$. We can view $P$ as a $c$-independent function on $\mathcal{J}_{0,\infty}$, and we note that ¯$\bar{\mathcal{L}}_{n}P=0$ for all $n\geq 1$.
Let $f\in\mathcal{E}$ and consider the path $\mathbb{R}_{+}\ni t\mapsto f_{t}(z):=e^{t}f(e^{-t}z)$ in $\mathcal{E}$, from $f$ to $\mathrm{id}_{\mathbb{D}}$. We have $\partial_{t}f_{t}(z)=e^{t}f_{t}(z)-zf^{\prime}(e^{-t}(z))=v_{t}\circ f_{t}(z)$, with $v_{t}(z)=z-\frac{f_{t}^{-1}(z)}{(f_{t}^{-1})^{\prime}(z)}$. For each $t>0$, the function $f_{t}^{-1}$ extends analytically in a neighbourhood of $f_{t}(\bar{\mathbb{D}})$, so $P$ is differentiable in direction $\mathrm{v}_{t}$ at $f_{t}$. By the fundamental theorem of calculus, we get for all $0<a<b$,
(3.7) $P(f_{b})-P(f_{a})=\int_{a}^{b}(\mathcal{L}_{\mathrm{v}_{t}}+\bar{\mathcal{L}}_{\mathrm{v}_{t}})P(f_{t})\mathrm{d}t=\int_{a}^{b}\mathcal{L}_{\mathrm{v}_{t}}P(f_{t})\mathrm{d}t.$
Since $P$ is continuous on $\mathcal{E}$, the LHS converges to $P(f)-P(\mathrm{id}_{\mathbb{D}})$ as $a\to 0$ and $b\to\infty$.
Now, suppose that $P$ is a singular vector for some representation $(\mathcal{V}_{\lambda},(\mathbf{L}_{n}^{\lambda})_{n\in\mathbb{Z}})$. Given the expression of $\mathbf{L}_{n}^{\lambda}$ for $n\geq 1$, we have $\mathcal{L}_{n}P=\mathbf{L}_{n}^{\lambda}P=0$ for all $n\geq 1$ (with a slight abuse of notation). In the neighbourhood of 0, we have the expansion $\frac{f_{t}^{-1}(z)}{(f_{t}^{-1})^{\prime}(z)}=z+O(z^{2})$, so $v_{t}(z)=O(z^{2})$. Hence, $\mathcal{L}_{\mathrm{v}_{t}}P=0$ for all $t>0$, and the RHS of (3.7) vanishes for all $0<a<b$. Taking the limit $a\to 0$, $b\to\infty$ gives $P(f)-P(\mathrm{id}_{\mathbb{D}})=0$. This shows that $P$ is constant on $\mathcal{E}$, so there is no singular vector in $\mathcal{V}_{\lambda}$ at a strictly positive level, i.e. $\mathcal{V}_{\lambda}$ is irreducible. The rest of the proof of Theorem 1.4 follows as before. ∎
3.3. Computation at level 2
We now give a computational proof of Corollary 1.5.
Computational proof of Corollary 1.5.
In this proof, we will denote by $\mathcal{A}\psi=\psi^{\prime\prime}/\psi^{\prime}$ the pre-Schwarzian derivative of a holomorphic function $\psi$. For a function $f\in\mathcal{E}$, we have in our parametrisation:
$\mathcal{A}f^{-1}(0)=-\mathcal{A}f(0)=-2a_{1},\qquad\mathcal{S}f^{-1}(0)=-\mathcal{S}f(0)=-6(a_{2}-a_{1}^{2}).$
We will slightly abuse notations by identifying $\varpi$ and $\vartheta$ with their restriction to $\mathcal{E}$ (so we evaluate them on $c=0$).
First, we compute $\mathbf{L}_{-1}^{\lambda}\mathds{1}$ and $\mathbf{L}_{-2}^{\lambda}\mathds{1}$. By Lemma 3.1, we have
$\mathbf{L}_{-1}^{\lambda}\mathds{1}=-2\lambda P_{-1,0}+Q_{-1,0}=-2\lambda\varpi(\mathrm{v}_{-1});$
$\mathbf{L}_{-2}^{\lambda}\mathds{1}=-2\lambda P_{-2,0}+Q_{-2,0}=-2\lambda\varpi(\mathrm{v}_{-2})-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{v}_{-2}).$
In order to evaluate $\varpi(\mathrm{v}_{-1})$ and $\varpi(\mathrm{v}_{-2})$, we record the following expansion, where $\psi(z)=z(1+\sum_{m=1}^{\infty}b_{m}z^{m})$ is conformal in the neighbourhood of $z=0$:
$\frac{z^{2}\psi^{\prime}(z)^{2}}{\psi(z)^{2}}$ $=\frac{(1+2b_{1}z+3b_{2}z^{2}+o(z^{2}))^{2}}{(1+b_{1}z+b_{2}z^{2}+o(z^{2}))^{2}}=1+2b_{1}z+(4b_{2}-b_{1}^{2})z^{2}+o(z^{2})$
$=1+\mathcal{A}\psi(0)z+(\frac{2}{3}\mathcal{S}\psi(0)+\frac{3}{4}\mathcal{A}\psi(0)^{2})z^{2}+o(z^{2}).$
Applying this computation to $\psi=f^{-1}$, we find that
$\varpi(\mathrm{v}_{-1})=-\frac{1}{2}\mathcal{A}f^{-1}(0)=a_{1}\quad\text{and}\quad\varpi(\mathrm{v}_{-2})=-\frac{1}{2}\Big{(}\frac{2}{3}\mathcal{S}f^{-1}(0)+\frac{3}{4}\mathcal{A}f^{-1}(0)^{2}\Big{)}=2(a_{2}-a_{1}^{2})-\frac{3}{2}a_{1}^{2}.$
Recalling that $\mathbf{L}_{\mathrm{v}}=\mathcal{L}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{v})$ and that $\vartheta(\mathrm{v}_{-1})=0$ and $\vartheta(\mathrm{v}_{-2})=-\mathcal{S}f^{-1}(0)=6(a_{2}-a_{1})^{2}$, we deduce that
(3.8) $\mathbf{L}_{-1}^{\lambda}\mathds{1}=-2\lambda a_{1}\quad\text{and}\quad\mathbf{L}_{-2}^{\lambda}\mathds{1}=3\lambda a_{1}^{2}-\Big{(}4\lambda+\frac{c_{\mathrm{m}}}{2}\Big{)}(a_{2}-a_{1}^{2}).$
Next, Proposition 2.8 gives $P_{-1,1}=3(a_{2}-a_{1}^{2})$ in the expression of $\mathbf{L}_{-1}^{\lambda}$ given in (3.4), so that
$(\mathbf{L}_{-1}^{\lambda})^{2}\mathds{1}=-2\lambda\mathbf{L}_{-1}^{\lambda}(a_{1})$ $=2\lambda(2\lambda+1)a_{1}^{2}-6\lambda(a_{2}-a_{1}^{2}).$
Wrapping up, we can present the values $(\mathbf{L}_{-1}^{2}\mathds{1},\mathbf{L}_{-2}\mathds{1})$ as a $2\times 2$-linear system:
$\begin{pmatrix}(\mathbf{L}_{-1}^{\lambda})^{2}\mathds{1}\\ \mathbf{L}_{-2}^{\lambda}\mathds{1}\end{pmatrix}=\begin{pmatrix}4\lambda(2\lambda+1)&6\lambda\\ 6\lambda&4\lambda+\frac{c_{\mathrm{m}}}{2}\end{pmatrix}\begin{pmatrix}\frac{1}{2}a_{1}^{2}\\ -(a_{2}-a_{1}^{2})\end{pmatrix}.$
We recover the Gram matrix of the Shapovalov form as it appears in [DFMS97, (7.23)], with the dictionary $\lambda\leftrightarrow h$, $c_{\mathrm{m}}\leftrightarrow c$. The determinant of this system is [DFMS97, (7.25-26)]
$\mathrm{det}\begin{pmatrix}4\lambda(2\lambda+1)&6\lambda\\ 6\lambda&4\lambda+\frac{c_{\mathrm{m}}}{2}\end{pmatrix}=32\lambda(\lambda-\lambda_{1,2})(\lambda-\lambda_{2,1}),\qquad\text{with}$
$\lambda_{1,2}=\frac{1}{16}\left(5-c_{\mathrm{m}}+\sqrt{(c_{\mathrm{m}}-1)(c_{\mathrm{m}}-25)}\right)\quad\text{and}\quad\lambda_{2,1}=\frac{1}{16}\left(5-c_{\mathrm{m}}-\sqrt{(c_{\mathrm{m}}-1)(c_{\mathrm{m}}-25)}\right).$
Elementary algebra gives $\lambda_{1,2}=\frac{6-\kappa}{2\kappa}$, $\lambda_{2,1}=\frac{3\kappa-8}{16}$ [DFMS97, (7.31)]. We deduce that $(\mathbf{L}_{-1}^{\lambda})^{2}\mathds{1}$ and $\mathbf{L}_{-2}^{\lambda}\mathds{1}$ are colinear for these values and these values only. It is an easy check that the corresponding linear combination for these values is $((\mathbf{L}_{-1}^{\lambda})^{2}-\frac{2}{3}(2\lambda+1)\mathbf{L}_{-2}^{\lambda})\mathds{1}=0$. ∎
4. Integration by parts
One of the main goals of this section is to prove the following integration by parts formula (recall the notations (1.3) of $\vartheta$ and $\tilde{\vartheta}$):
Theorem 4.1.
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, the $L^{2}(\nu)$-adjoints $\mathcal{L}_{\mathrm{v}}^{*}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}^{*}$ of $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$, respectively, are densely defined. Hence, $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$ are closable, and their closures satisfy
(4.1) $\mathcal{L}_{\mathrm{v}}^{*}=-\bar{\mathcal{L}}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\overline{\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})}+\frac{c_{\mathrm{m}}}{12}\overline{\vartheta_{e^{c}f}(\mathrm{v})};\qquad\bar{\mathcal{L}}_{\mathrm{v}}^{*}=-\mathcal{L}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})+\frac{c_{\mathrm{m}}}{12}\vartheta_{e^{c}f}(\mathrm{v}).$
We will keep denoting by $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$ the corresponding closures.
This result is essentially equivalent to [GQW24, Corollary 3.11], but the proof methods are different: we relate the differential of the mass of Brownian loops to the Brownian bubble measure, while [GQW24] rely on results from [CW23, SW23].
In Proposition 4.6 we derive an analogous result for the SLE loop measure in the unit disc.
Theorem 4.1 allows us to construct the Shapovalov form $\mathcal{Q}$. To this end, recall that we define an antilinear involution of $\mathbb{C}(z)\partial_{z}$ by $\mathrm{v}^{*}=z^{2}\bar{v}(1/z)\partial_{z}$, an involution $\Theta$ of $L^{2}(\nu)$ (see (1.6)) and a Hermitian $\mathcal{Q}$ on $L^{2}(\nu)$ by $\mathcal{Q}(F,G)=\langle F,\Theta G\rangle_{L^{2}(\nu)}$, $F,G\in L^{2}(\nu)$.
Corollary 4.2.
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, the $\mathcal{Q}$-adjoint of $\mathbf{L}_{\mathrm{v}}$ is densely defined. Hence $\mathbf{L}_{\mathrm{v}}$ is closable and its closure satisfies
(4.2) $\mathbf{L}_{\mathrm{v}}^{*}=\Theta\circ\mathbf{L}_{\mathrm{v}^{*}}\circ\Theta.$
In particular, for all $F\in\mathcal{D}(\mathcal{L}_{\mathrm{v}^{*}}),\,G\in\mathcal{D}(\mathcal{L}_{\mathrm{v}})$, we have $\mathcal{Q}(\mathbf{L}_{\mathrm{v}^{*}}F,G)=\mathcal{Q}(F,\mathbf{L}_{\mathrm{v}}G).$
Remark 4.3.
In Section 3, we defined the modules $\mathcal{W}_{\lambda}$ as algebraic subspaces of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$, and the representations $(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})_{n\in\mathbb{Z}}$ as endomorphisms acting on these spaces. Such definitions are sufficient for the purpose of characterising these modules algebraically. Now, we can also consider the closure of $\mathcal{W}_{\lambda}$ with respect to the $L^{2}(\nu^{\#})$; this is a closed subspace of $L^{2}(\nu^{\#})$, which is proper if and only if $\lambda$ lies on the Kac table. The operators $\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda}$ extend to unbounded operators on the closure, and one can deduce from the above integration by parts formula that they are closable (the adjoints are densely defined).
4.1. Proof of Corollary 4.2, assuming Theorem 4.1
In this section, we obtain Corollary 4.2, assuming Theorem 4.1. We start with a preliminary lemma that will also be useful in the proof of Theorem 4.1.
Lemma 4.4.
For all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$,
(4.3) $\Theta\circ\mathcal{L}_{\mathrm{v}^{*}}\circ\Theta=-\bar{\mathcal{L}}_{\mathrm{v}};\qquad\Theta\circ\bar{\mathcal{L}}_{\mathrm{v}^{*}}\circ\Theta=-\mathcal{L}_{\mathrm{v}};$
(4.4) $\Theta\circ(\vartheta_{e^{c}f}(\mathrm{v}^{*})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}})\circ\Theta=\overline{\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})}\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}};\qquad\Theta\circ(\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v}^{*})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}})\circ\Theta=\overline{\vartheta_{e^{c}f}(\mathrm{v})}\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}.$
Proof.
We start by proving (4.3). From the explicit expressions of $\mathcal{L}_{0}$ and ¯$\bar{\mathcal{L}}_{0}$ from Proposition 2.8, one can directly check (4.3) in this case. Since $\mathrm{v}\mapsto\mathrm{v}^{*}$ maps $z^{2}\mathbb{C}[z]\partial_{z}$ to $\mathbb{C}[z^{-1}]\partial_{z}$, if (4.3) holds for all vector fields in $z^{2}\mathbb{C}[z]\partial_{z}$, by applying $\Theta$ to the left and to the right of (4.3), we see that it also holds for all vector fields in $\mathbb{C}[z^{-1}]\partial_{z}$. By linearity, the vector field $\mathrm{v}_{0}$ covering the case of polynomial vectors fields of degree one in $z$, we would then get (4.3) for all $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$. It is thus sufficient to consider the case $\mathrm{v}\in z^{2}\mathbb{C}[z]\partial_{z}$. Let $\phi_{t}=\mathrm{Id}+tv^{*}+o(t)$ be the flow generated by $\mathrm{v}^{*}$ for small complex $t$. Let $\eta_{t}:=\phi_{t}\circ\eta$ and $(c_{\eta_{t}},f_{\eta_{t}})\in\mathbb{R}\times\mathcal{E}$, $(\tilde{c}_{\eta_{t}},g_{\eta_{t}})\in\mathbb{R}\times\tilde{\mathcal{E}}$ be the uniformising maps of the interior and exterior of $\eta$. Since $\mathrm{v}^{*}\in\mathbb{C}[z^{-1}]\partial_{z}$, we have $e^{-\tilde{c}_{\eta_{t}}}g_{\eta_{t}}=\phi_{t}\circ(e^{-\tilde{c}_{\eta}}g_{\eta})$. A small computation then shows that $\iota\circ g_{\eta_{t}}\circ\iota=\iota\circ g_{\eta}\circ\iota-\bar{t}v\circ(\iota\circ g_{\eta}\circ\iota)+o(t)$. By definition of $\mathcal{L}_{\mathrm{v}}$ and ¯$\bar{\mathcal{L}}_{\mathrm{v}}$, we then get that $(\Theta F)(c_{\eta_{t}},f_{\eta_{t}})$ is equal to
$F(\tilde{c}_{\eta_{t}},\iota\circ g_{\eta_{t}}\circ\iota)=F(\tilde{c}_{\eta},\iota\circ g_{\eta}\circ\iota)-t\bar{\mathcal{L}}_{\mathrm{v}}F(\tilde{c}_{\eta},\iota\circ g_{\eta}\circ\iota)-\bar{t}\mathcal{L}_{\mathrm{v}}F(\tilde{c}_{\eta},\iota\circ g_{\eta}\circ\iota)+o(t).$
Thus,
$\mathcal{L}_{\mathrm{v}^{*}}\circ\Theta F(c_{\eta},f_{\eta})=-\bar{\mathcal{L}}_{\mathrm{v}}F(\tilde{c}_{\eta},\iota\circ g_{\eta}\circ\iota)\quad\text{and}\quad\bar{\mathcal{L}}_{\mathrm{v}^{*}}\circ\Theta F(c_{\eta},f_{\eta})=-\mathcal{L}_{\mathrm{v}}F(\tilde{c}_{\eta},\iota\circ g_{\eta}\circ\iota)$
Applying $\Theta$ again gives (4.3).
We now prove (4.4). It is enough to prove the first equality. Let $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$. Note first that
$\Theta\circ(\vartheta_{e^{c}f}(\mathrm{v}^{*})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}})\circ\Theta=\Theta(\vartheta_{e^{c}f}(\mathrm{v}^{*}))\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}.$
By definition of $\Theta$ and $\vartheta$, we have
$\Theta(\vartheta_{e^{c}f}(\mathrm{v}^{*}))=\vartheta_{e^{\tilde{c}}\iota\circ g\circ\iota}(\mathrm{v}^{*})=\frac{e^{-\tilde{c}}}{2i\pi}\oint\mathcal{S}(\iota\circ g^{-1}\circ\iota)(z)v^{*}(e^{\tilde{c}}z)\mathrm{d}z.$
By the chain rule of the Schwarzian, $\mathcal{S}(\iota\circ g^{-1}\circ\iota)=\frac{1}{z^{4}}\overline{\mathcal{S}g^{-1}}(1/z).$ Plugging $v^{*}(z)=z^{2}\bar{v}(1/z)$ and performing the change of variables $z\mapsto 1/\bar{z}$, we deduce that
$\Theta(\vartheta_{e^{c}f}(\mathrm{v}^{*}))=-\frac{e^{\tilde{c}}}{2i\pi}\oint\overline{\mathcal{S}g^{-1}}(\bar{z})\bar{v}(e^{\tilde{c}}\bar{z})\mathrm{d}\bar{z}=\overline{\frac{e^{\tilde{c}}}{2i\pi}\oint\mathcal{S}g^{-1}(z)v(e^{\tilde{c}}z)\mathrm{d}z}=\overline{\vartheta_{e^{-\tilde{c}}g}(\mathrm{v})}$
as desired. ∎
Proof of Corollary 4.2, assuming Theorem 4.1.
Let $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$. By definition (1.4) of $\mathbf{L}_{\mathrm{v}^{*}}$ in terms of $\mathcal{L}_{\mathrm{v}^{*}}$ and combining (4.3) and (4.4), we have
$\Theta\circ\mathbf{L}_{\mathrm{v}^{*}}\circ\Theta=\Theta\circ(\mathcal{L}_{\mathrm{v}^{*}}-\frac{c_{\mathrm{m}}}{12}\vartheta_{e^{c}f}(\mathrm{v}^{*})\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}})\circ\Theta=-\bar{\mathcal{L}}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\overline{\tilde{\vartheta}_{e^{-\tilde{c}}g}(\mathrm{v})}\mathrm{Id}_{\mathcal{C}_{\mathrm{comp}}}.$
By Theorem 4.1, the right-hand side agrees with $\mathbf{L}_{\mathrm{v}}^{*}$ showing the desired relation $\mathbf{L}_{\mathrm{v}}^{*}=\Theta\circ\mathbf{L}_{\mathrm{v}^{*}}\circ\Theta.$ The last statement of Corollary 4.2 is immediate since then
$\mathcal{Q}(\mathbf{L}_{\mathrm{v}^{*}}F,G)=\langle\mathbf{L}_{\mathrm{v}^{*}}F,\Theta G\rangle_{L^{2}(\nu)}=\langle F,\Theta\circ\mathbf{L}_{\mathrm{v}}G\rangle_{L^{2}(\nu)}=\mathcal{Q}(F,\mathbf{L}_{\mathrm{v}}G).$
4.2. Integration by parts for the SLE loop measure in the unit disc
In this section, we study the SLE loop measure $\nu_{\mathbb{D}}$ in the unit disc $\mathbb{D}$. We start by introducing a few notations:
Notation 4.5.
For $q\in(0,1)$, we will denote by $\mathbb{A}_{q}$ the annulus $\mathbb{D}\setminus\overline{q\mathbb{D}}$. Let $\eta$ be a Jordan curve in $\mathbb{D}$. Let $\mathbb{A}_{\eta}:=\mathbb{D}\setminus\overline{\mathrm{int}(\eta)}$ be the annulus-type domain encircled by $\mathbb{S}^{1}$ and $\eta$. There exists a unique $q(\eta)\in(0,1)$ and a conformal map $\psi_{\eta}:\mathbb{A}_{\eta}\to\mathbb{A}_{q(\eta)}$ with $\psi_{\eta}(1)=1$.
Note that for all Jordan curve $\eta$ in $\mathbb{D}$, $\psi_{\eta}$ extends to a neighbourhood of $\mathbb{S}^{1}$ by Schwarz reflection, and in particular $\mathcal{S}\psi_{\eta}$ is holomorphic in a neighbourhood of $\mathbb{S}^{1}$. Given $\mathrm{v}\in\mathbb{C}(z)\partial_{z}$, we will write $(\mathcal{S}\psi_{\eta},\mathrm{v}):=\frac{1}{2i\pi}\oint_{\mathbb{S}^{1}}\mathcal{S}\psi_{\eta}(z)v(z)\mathrm{d}z$ the residue pairing and similarly for $(\psi_{\eta}^{-2}{\psi_{\eta}^{\prime}}^{2},\mathrm{v})$.
The main result of this section is the following integration by parts formula for $\nu_{\mathbb{D}}$:
Proposition 4.6.
For all $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$ and $F,G\in\mathcal{C}$, we have
(4.5) $(\mathcal{L}_{\mathrm{v}}F,G)_{L^{2}(\nu_{\mathbb{D}})}=\Big{(}F,-\bar{\mathcal{L}}_{\mathrm{v}}G-\Big{(}\frac{c_{\mathrm{m}}}{12}\overline{(\mathcal{S}\psi_{\eta},\mathrm{v})}+\frac{c_{\mathrm{m}}}{2}U(q(\eta))\overline{(\psi_{\eta}^{-2}{\psi_{\eta}^{\prime}}^{2},\mathrm{v})}\hskip 1.0pt\Big{)}G\Big{)}_{L^{2}(\nu_{\mathbb{D}})},$
(4.6) $\text{where}\qquad U(q):=\frac{1}{12}+\frac{\pi^{2}}{12|\log q|^{2}}-\frac{\pi^{2}}{2|\log q|^{2}}\sum_{n\geq 1}\sinh^{-2}\Big{(}\frac{n\pi^{2}}{|\log q|}\Big{)},\qquad q\in(0,1).$
Notice that $\lim_{q\to 0^{+}}U(q)=0$. As a consequence, the term featuring $U(q(\eta))$ disappears when passing to the limit to get the integration by parts formula in the sphere as stated in Theorem 4.1. This term arises in computations with the Brownian bubble measure on the annulus $\mathbb{A}_{q}$, and is related to the boundary excursion kernel (see Section 4.3 for details). The limit $U(q)\to 0$ can be interpreted probabilistically as the fact that $\{0\}$ is polar for the bubble measure in $\mathbb{D}$.
Brownian loop measure and Brownian bubble measure The proof of Proposition 4.6 will make use of two families of measures on Brownian paths (measures on continuous paths of finite duration). As introduced by Lawler–Werner [LW04], the whole-plane Brownian loop measure is defined by
$\mu^{\mathrm{loop}}_{\mathbb{C}}=\int_{0}^{\infty}\frac{\mathrm{d}t}{t}\int_{\mathbb{C}}|\mathrm{d}z|^{2}\frac{1}{2\pi t}\mathbb{P}_{D}^{t,z,z}$
where $\mathbb{P}_{D}^{t,z,z}$ is the Brownian bridge measure from $z$ to $z$ with duration $t$ (in our convention, the associated generator is $\frac{1}{2}\Delta$). For a domain $D\subset\mathbb{C}$, it can be defined by restricting the whole plane measure to $D$:
$\mu^{\mathrm{loop}}_{D}(\mathrm{d}\mathcal{P})=\mathbf{1}_{\mathcal{P}\subset D}\mu^{\mathrm{loop}}_{\mathbb{C}}(\mathrm{d}\mathcal{P}).$
On the other hand, for a smooth bounded simply connected domain $D$ (we will only consider discs) and a boundary point $z\in\partial D$, the Brownian bubble measure in $D$ rooted at $z$ is a $\sigma$-finite measure on Brownian excursions from $z$ to $z$ in $D$; see [LW04, Section 3.4].
Given a subdomain $D\subset\mathbb{D}$ of the unit disc such that $\mathbb{C}\setminus D$ is non polar, the Brownian loop measure naturally appears in the Radon–Nikodym derivative of $\nu_{D}$ with respect to $\nu_{\mathbb{D}}$ (see [Zha21, Equation (5.3)]):
(4.7) $\mathrm{d}\nu_{D}(\eta)=e^{\frac{c_{\mathrm{m}}}{2}\Lambda_{\mathbb{D}}(\eta,\mathbb{D}\setminus D)}\mathbf{1}_{\eta\subset D}\mathrm{d}\nu_{\mathbb{D}}(\eta),\quad\text{where}\quad\Lambda_{\mathbb{D}}(\eta,\mathbb{D}\setminus D)=\mu^{\mathrm{loop}}_{\mathbb{D}}(\text{hit both }\eta\text{ and }\mathbb{D}\setminus D).$
Here, “hit both $\text{hit both }\eta\text{ and }\mathbb{D}\setminus D$” is the event that the sample of $\mu^{\mathrm{loop}}_{D}$ intersects both the trace of $\eta$ and the boundary of $\mathbb{D}\setminus D$. This transparent notation will be used throughout the proof. In Definition 1.8, the conformal restriction property of the SLE loop measure was phrased slightly differently since we started with a measure on the sphere. These two restriction properties agree since for any simply connected domains $D_{1},D_{2}$ with $\overline{D_{2}}\subset D_{1}$ and for any Jordan curve $\eta$ in $D_{2}$,
(4.8) $\Lambda_{D_{1}}(\eta,D_{2}^{c})=\Lambda^{*}(\eta,D_{2}^{c})-\Lambda^{*}(\eta,D_{1}^{c}).$
This last relation in turn simply follows from the restriction property of the Brownian loop measure.
Proof of Proposition 4.6.
We will prove (4.5) for so-called Markovian vector fields $\mathrm{v}=v\partial_{z}\in\mathbb{C}[z]\partial_{z}$ (following the terminology of [BGK+24]), i.e. vector fields satisfying $\mathrm{Re}(\bar{z}v(z))<0$ for all $z\in\mathbb{S}^{1}$. This will be enough since they linearly generate $\mathbb{C}[z]\partial_{z}$. So consider a Markovian vector field $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$ and the associated flow $\phi_{t}(z)=z+tv(z)+o(t)$ defined for small complex values of $t$. Because $\mathrm{v}$ is Markovian, when $t$ belongs to a cone $\{\arg(t)\in(-\delta,\delta)\}$ that is narrow enough, the maps $\phi_{t}$ are well defined for all $t$ in this cone and contract the unit disc, i.e. $\phi_{t}(\bar{\mathbb{D}})\subset\mathbb{D}$ [BGK+24]. In this proof we will always restrict ourselves to such values of $t$. Let $F,G\in\mathcal{C}$. By dominated convergence theorem, $(\mathcal{L}_{\mathrm{v}}F,G)_{L^{2}(\nu_{\mathbb{D}})}$ corresponds to the $t$-coefficient in the expansion of
$\int(F(\phi_{t}\cdot\eta)-F(\eta))\overline{G(\eta)}\mathrm{d}\nu_{\mathbb{D}}(\eta).$
By conformal invariance of the SLE loop measure and then by (4.7) (using that $\phi_{t}(\bar{\mathbb{D}})\subset\mathbb{D}$), the above display equals
(4.9) $\int F(\eta)\overline{G(\phi_{t}^{-1}\cdot\eta)}\mathrm{d}\nu_{\phi_{t}(\mathbb{D})}(\eta)-\int F(\eta)\overline{G(\eta)}\mathrm{d}\nu_{\mathbb{D}}(\eta)$
$=\int F(\eta)\Big{(}\overline{G(\phi_{t}^{-1}\cdot\eta)}e^{\frac{c_{\mathrm{m}}}{2}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))}\mathbf{1}_{\eta\subset\phi_{t}(\mathbb{D})}-\overline{G(\eta)}\Big{)}\mathrm{d}\nu_{\mathbb{D}}(\eta).$
In Lemmas 4.7 and 4.8 below, we will show that
(4.10) $\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))=-t\Big{(}\frac{1}{6}(\mathcal{S}\psi_{\eta},\mathrm{v})+U(q(\eta))(\psi_{\eta}^{-2}{\psi_{\eta}^{\prime}}^{2},\mathrm{v})\Big{)}-\bar{t}\Big{(}\frac{1}{6}(\mathcal{S}\psi_{\eta},\mathrm{v}^{*})+U(q(\eta))(\psi_{\eta}^{-2}{\psi_{\eta}^{\prime}}^{2},\mathrm{v}^{*})\Big{)}+o(t).$
The $t$-coefficient in the expansion of the term in parenthesis on the right hand side of (4.9) is thus equal to
¯$-\overline{\bar{\mathcal{L}}_{\mathrm{v}}G(\eta)}-\frac{c_{\mathrm{m}}}{12}\overline{G(\eta)}(\mathcal{S}\psi_{\eta},\mathrm{v})-\frac{c_{\mathrm{m}}}{2}U(q(\eta))\overline{G(\eta)}(\psi_{\eta}^{-2}{\psi_{\eta}^{\prime}}^{2},\mathrm{v}).$
By dominated convergence theorem we then get (4.5).
The rest of the section is dedicated to the proof of (4.10). Up to considering the vector field $e^{i\theta}\mathrm{v}$ for $|\theta|<\delta$, it suffices to consider the case of small positive values of $t$. We decompose the proof into two steps, corresponding to Lemmas 4.7 and 4.8 respectively.
Lemma 4.7.
The following derivative estimate holds:
(4.11) $\lim_{\begin{subarray}{c}t\to 0\\ t>0\end{subarray}}\frac{1}{t}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))=-\frac{1}{\pi}\int_{0}^{2\pi}\mathrm{d}\theta\leavevmode\nobreak\ \mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)\mathrm{Re}(e^{-i\theta}v(e^{i\theta})).$
Proof.
We start by recalling the following decomposition of the loop measure which amounts to rooting the loop at the point of maximal modulus (see [LW04, Proposition 8]):
$\mu^{\mathrm{loop}}_{\mathbb{C}}=\frac{1}{\pi}\int_{\mathbb{C}}|\mathrm{d}z|^{2}\leavevmode\nobreak\ \mu^{\mathrm{bub}}_{|z|\mathbb{D},z}.$
This leads to
(4.12) $\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))=\frac{1}{\pi}\int_{0}^{1}\mathrm{d}r\leavevmode\nobreak\ r\int_{0}^{2\pi}\mathrm{d}\theta\leavevmode\nobreak\ \mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ both\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1})\text{ and }\eta).$
We also record the following estimate: for any $\theta\in[0,2\pi]$, $\phi_{t}(e^{i\theta})=e^{i\theta}+tv(e^{i\theta})+o(t)$ implying that
(4.13) $|\phi_{t}(e^{i\theta})|=1+\mathrm{Re}(e^{-i\theta}v(e^{i\theta}))t+o(t).$
We now provide an upper bound for $\limsup_{t\to 0}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))/t$. Let $\delta>0$ be small but fixed and $C:=\max_{0\leq\theta\leq 2\pi}\{-\mathrm{Re}(e^{-i\theta}v(e^{i\theta}))\}$. Let $\theta\in[0,2\pi]$. By (4.13), if $r\leq 1-(C+\delta)t$, then $\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1}))$ vanishes for $t$ small enough. Moreover, we claim that if $1-(C+\delta)t\leq r\leq|\phi_{t}(e^{i\theta})|-\delta t$, then
(4.14) $\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1}))\to 0\quad\text{as}\quad t\to 0.$
Indeed, in this case and by (4.13) and continuity of $v$, there exists $c=c(\delta)>0$ such that for all $\theta^{\prime}\in[\theta-c,\theta+c]$, $\phi_{t}(e^{i\theta^{\prime}})\notin r\mathbb{D}$ and for $\theta^{\prime}$ outside this neighbourhood, $|\phi_{t}(e^{i\theta^{\prime}})|\geq 1-Ct+o(t)$. So to get a chance of hitting $\phi_{t}(\mathbb{S}^{1})$, the loop rooted at $re^{i\theta}$ would have to at least travel a macroscopic distance and hit a small region located near the boundary of $r\mathbb{D}$ (in the case where $\phi_{t}(\mathbb{S}^{1})\cap r\mathbb{D}\neq\varnothing$). This shows (4.14). Since the length of the interval $[1-(C+\delta)t,|\phi_{t}(e^{i\theta})|-\delta t]$ is of order $t$, it shows that
(4.15) $\lim_{t\to 0}\frac{1}{t}\int_{0}^{|\phi_{t}(e^{i\theta})|-\delta t}\mathrm{d}r\leavevmode\nobreak\ r\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ both\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1})\text{ and }\eta)=0.$
Consider now the case $r\geq|\phi_{t}(e^{i\theta})|-\delta t$. We can bound
$\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ both\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1})\text{ and }\eta)\leq\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ }\eta).$
The right hand side converges to $\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit\leavevmode\nobreak\ }\eta)$ as $r\to 1$. Together with (4.13) and (4.15), we obtain that
$\limsup_{t\to 0}\frac{1}{t}\int_{0}^{1}\mathrm{d}r\leavevmode\nobreak\ r\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ both\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1})\text{ and }\eta)\leq(-\mathrm{Re}(e^{-i\theta}v(e^{i\theta}))+\delta)\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit\leavevmode\nobreak\ }\eta).$
Since the left hand side does not depend on $\delta$, and going back to (4.12), this shows that
$\limsup_{t\to 0}\frac{1}{t}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))\leq-\frac{1}{\pi}\int_{0}^{2\pi}\mathrm{d}\theta\leavevmode\nobreak\ \mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)\mathrm{Re}(e^{-i\theta}v(e^{i\theta}))$
as desired.
To get a lower bound, we simply bound
$\liminf_{t\to 0}\frac{1}{t}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))\geq\frac{1}{\pi}\liminf_{t\to 0}\frac{1}{t}\int_{0}^{2\pi}\mathrm{d}\theta\int_{|\phi_{t}(e^{i\theta})|+\delta t}^{1}\mathrm{d}r\leavevmode\nobreak\ r\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ both\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1})\text{ and }\eta)$
where $\delta>0$ is small but fixed. If $r\geq|\phi_{t}(e^{i\theta})|+\delta t$ and $t$ is small enough, $re^{i\theta}\in\phi_{t}(\mathbb{D})^{c}$ and
$\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ both\leavevmode\nobreak\ }\phi_{t}(\mathbb{S}^{1})\text{ and }\eta)=\mu^{\mathrm{bub}}_{r\mathbb{D},re^{i\theta}}(\text{hit\leavevmode\nobreak\ }\eta)=\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit\leavevmode\nobreak\ }\eta)-o(1)$
so that
$\liminf_{t\to 0}\frac{1}{t}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))\geq-\frac{1}{\pi}\int_{0}^{2\pi}\mathrm{d}\theta\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit\leavevmode\nobreak\ }\eta)(\mathrm{Re}(e^{-i\theta}v(e^{i\theta}))+\delta).$
Sending $\delta\to 0$ finishes the proof. ∎
We now relate $\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)$ to the quantity appearing in the statement of Proposition 4.6. Recall that $q(\eta)$ and $\psi_{\eta}$ are introduced in Notation 4.5 and that $U(q)$ is defined in (4.6).
Lemma 4.8.
Let $\eta$ be a Jordan curve in $\mathbb{D}$. Then, for any $\theta\in[0,2\pi]$, $e^{2i\theta}\mathcal{S}\psi_{\eta}(e^{i\theta})\in\mathbb{R}$ and $e^{i\theta}\psi_{\eta}^{\prime}(e^{i\theta})/\psi_{\eta}(e^{i\theta})\in\mathbb{R}$ are real numbers and
(4.16) $\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)=\frac{1}{6}e^{2i\theta}\mathcal{S}\psi_{\eta}(e^{i\theta})+e^{2i\theta}\frac{\psi_{\eta}^{\prime}(e^{i\theta})^{2}}{\psi_{\eta}(e^{i\theta})^{2}}U(q(\eta)).$
Remark 4.9.
This formula can be seen as an extension of [LW04, Equation (13)] to a non simply connected setup.
Proof.
We start by checking that $e^{2i\theta}\mathcal{S}\psi_{\eta}(e^{i\theta})\in\mathbb{R}$. This follows from the following more general result. Let $\chi$ be an analytic diffeomorphism of $\mathbb{S}^{1}$. Then
(4.17) for $\text{for all }z\in\mathbb{S}^{1},\quad z^{2}\mathcal{S}\chi(z)\in\mathbb{R}.$
To prove (4.17), we write locally $\chi(e^{i\theta})=e^{ih(\theta)}$ for some analytic function $h$ defined in a complex neighbourhood of the real line, with $h(\theta)\in\mathbb{R}$ if $\theta\in\mathbb{R}$. In particular, $\mathcal{S}h\in\mathbb{R}$ on $\mathbb{R}$ and $h^{\prime}\in\mathbb{R}$ on $\mathbb{R}$. Let us denote $e_{\alpha}(z):=e^{\alpha z}$ and observe that $\mathcal{S}e_{\alpha}(z)=-\frac{\alpha^{2}}{2}$. By the chain rule for the Schwarzian, we get
$\frac{1}{2}-e^{2i\theta}\mathcal{S}\chi(e^{i\theta})=\mathcal{S}h(\theta)+\frac{1}{2}h^{\prime}(\theta)^{2}.$
Hence, $e^{2i\theta}\mathcal{S}\chi(e^{i\theta})=-\mathcal{S}h(\theta)+\frac{1}{2}(1-h^{\prime}(\theta)^{2})$ is real for any $\theta\in\mathbb{R}$. This proves (4.17). The proof that
(4.18) $e^{i\theta}\frac{\psi_{\eta}^{\prime}(e^{i\theta})}{\psi_{\eta}(e^{i\theta})}=|\psi_{\eta}^{\prime}(e^{i\theta})|$
proceeds similarly and we omit the details.
We now turn to the proof of (4.16). Let $\theta\in[0,2\pi]$. We start by writing
$\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)=\pi\lim_{\theta^{\prime}\to\theta}\mu_{\mathbb{D},e^{i\theta},e^{i\theta^{\prime}}}^{\mathrm{exc}}(\text{hit }\eta)$
where $\mu_{\mathbb{D},e^{i\theta},e^{i\theta^{\prime}}}^{\mathrm{exc}}$ is a measure on Brownian excursions from $e^{i\theta}$ to $e^{i\theta^{\prime}}$ whose total mass equals the boundary Poisson kernel $H_{\mathbb{D}}(e^{i\theta},e^{i\theta^{\prime}})$. See [LW04, Sections 3.3 and 3.4]. By the restriction property, we further have
$\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)=\pi\lim_{\theta^{\prime}\to\theta}(|\mu_{\mathbb{D},e^{i\theta},e^{i\theta^{\prime}}}^{\mathrm{exc}}|-|\mu_{\mathbb{A}_{\eta},e^{i\theta},e^{i\theta^{\prime}}}^{\mathrm{exc}}|)=\pi\lim_{\theta^{\prime}\to\theta}(H_{\mathbb{D}}(e^{i\theta},e^{i\theta^{\prime}})-H_{\mathbb{A}_{\eta}}(e^{i\theta},e^{i\theta^{\prime}})).$
The rest of the proof is then concerned with Poisson kernel computations. The boundary Poisson kernel in the unit disc is explicit and given by $H_{\mathbb{D}}(z,w)=\frac{1}{\pi|z-w|^{2}}$, for all $z,w\in\mathbb{S}^{1}$. To deal with the Poisson kernel in $\mathbb{A}_{\eta}$, we first notice that, by conformal covariance of the boundary Poisson kernel [LW04, Section 3.3],
$H_{\mathbb{A}_{\eta}}(e^{i\theta},e^{i\theta^{\prime}})=|\psi_{\eta}^{\prime}(e^{i\theta})||\psi_{\eta}^{\prime}(e^{i\theta^{\prime}})|H_{\mathbb{A}_{q(\eta)}}(\psi_{\eta}(e^{i\theta}),\psi_{\eta}(e^{i\theta^{\prime}})).$
The multivalued conformal map $z\mapsto\log z/|\log q|$ maps the annulus $\mathbb{A}_{q}$ to the vertical strip $V:=\{z\in\mathbb{C}:-1<\mathrm{Re}(z)<0\}$, so
$H_{\mathbb{A}_{q}}(e^{i\theta},e^{i\theta^{\prime}})=\frac{1}{|\log q|^{2}}\sum_{n\in\mathbb{Z}}H_{V}\Big{(}i\frac{\theta}{|\log q|},i\frac{\theta^{\prime}+2n\pi}{|\log q|}\Big{)},$
the factor $1/|\log q|^{2}$ coming from the derivative of the conformal map. The boundary Poisson kernel in the strip $V$ is explicit (this boils down to finding an explicit conformal map from $V$ to the unit disc; see e.g. [Wid61]): for all $y,y^{\prime}\in\mathbb{R}$,
$H_{V}(iy,iy^{\prime})=\frac{\pi}{4}\sinh^{-2}\Big{(}\frac{\pi}{2}(y-y^{\prime})\Big{)}.$
This shows that
$H_{\mathbb{A}_{q}}(e^{i\theta},e^{i\theta^{\prime}})=\frac{\pi}{4}\frac{1}{|\log q|^{2}}\sum_{n\in\mathbb{Z}}\sinh^{-2}\Big{(}\frac{\pi}{2}\frac{\theta^{\prime}-\theta+2n\pi}{|\log q|}\Big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\frac{1}{\pi|e^{i\theta}-e^{i\theta^{\prime}}|^{2}}-\frac{1}{12\pi}-\frac{\pi}{12|\log q|^{2}}+\frac{\pi}{2|\log q|^{2}}\sum_{n\geq 1}\sinh^{-2}\Big{(}\frac{n\pi^{2}}{|\log q|}\Big{)}+o(1)\qquad\text{as }\theta^{\prime}\to\theta.$
Putting things together and using the notation $U(q)$ (4.6),
(4.19) $\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)$ $=\lim_{\theta^{\prime}\to\theta}\Big{(}\frac{1}{|e^{i\theta}-e^{i\theta^{\prime}}|^{2}}-\frac{|\psi_{\eta}^{\prime}(e^{i\theta})\psi_{\eta}^{\prime}(e^{i\theta^{\prime}})|}{|\psi_{\eta}(e^{i\theta^{\prime}})-\psi_{\eta}(e^{i\theta})|^{2}}\Big{)}+|\psi_{\eta}^{\prime}(e^{i\theta})|^{2}U(q(\eta)).$
To compute the above limit, one can do some tedious Taylor expansion computations and find the result. We proceed differently by noticing that for any distinct $z,w\in\mathbb{S}^{1}$,
$\frac{1}{|z-w|^{2}}=-\mathrm{Re}\Big{\{}\frac{zw}{(z-w)^{2}}\Big{\}}.$
Combining this with (4.18), we deduce that
$\lim_{\theta^{\prime}\to\theta}\Big{(}\frac{1}{|e^{i\theta}-e^{i\theta^{\prime}}|^{2}}-\frac{|\psi_{\eta}^{\prime}(e^{i\theta})\psi_{\eta}^{\prime}(e^{i\theta^{\prime}})|}{|\psi_{\eta}(e^{i\theta^{\prime}})-\psi_{\eta}(e^{i\theta})|^{2}}\Big{)}=\lim_{\theta^{\prime}\to\theta}\mathrm{Re}\Big{\{}e^{i\theta}e^{i\theta^{\prime}}\Big{(}\frac{\psi_{\eta}^{\prime}(e^{i\theta})\psi_{\eta}^{\prime}(e^{i\theta^{\prime}})}{(\psi_{\eta}(e^{i\theta})-\psi_{\eta}(e^{i\theta^{\prime}}))^{2}}-\frac{1}{(e^{i\theta}-e^{i\theta^{\prime}})^{2}}\Big{)}\Big{\}}.$
The term in parenthesis on the right hand side converges as $\theta^{\prime}\to\theta$ to
$\frac{\partial^{2}}{\partial z\partial w}\log\Big{(}\frac{\psi_{\eta}(z)-\psi_{\eta}(w)}{z-w}\Big{)}\Big{|}_{z=w=e^{i\theta}}$
which agrees with $\mathcal{S}\psi_{\eta}(e^{i\theta})/6$ by an elementary property of the Schwarzian derivative. Going back to (4.19), recalling that $e^{2i\theta}\mathcal{S}\psi_{\eta}(e^{i\theta})\in\mathbb{R}$ and using (4.18) once more, we have obtained (4.16) as desired. ∎
Combining Lemmas 4.7 and 4.8, we obtain
$\lim_{\begin{subarray}{c}t\to 0\\ t>0\end{subarray}}\frac{1}{t}\Lambda_{\mathbb{D}}(\eta,\phi_{t}(\mathbb{S}^{1}))$ $=-\mathrm{Re}\Big{\{}\frac{1}{\pi}\int_{0}^{2\pi}\mathrm{d}\theta\Big{(}\frac{1}{6}e^{2i\theta}\mathcal{S}\psi_{\eta}(e^{i\theta})+e^{2i\theta}\frac{\psi_{\eta}^{\prime}(e^{i\theta})^{2}}{\psi_{\eta}(e^{i\theta})^{2}}U(q(\eta))\Big{)}e^{-i\theta}v(e^{i\theta})\Big{\}}$
$=-2\mathrm{Re}\Big{\{}\frac{1}{6}(\mathcal{S}\psi_{\eta},\mathrm{v})+U(q(\eta))(\psi_{\eta}^{-2}{\psi_{\eta}^{\prime}}^{2},\mathrm{v})\Big{\}}$
which corresponds to the asymptotic behaviour (4.10) when $t$ is real. This concludes the proof of Proposition 4.6. ∎
4.3. Proof of Theorem 4.1
In this section, we will use a limiting procedure together with Proposition 4.6 to prove Theorem 4.1. We start with two lemmas.
Lemma 4.10.
We have $(\log R)^{\frac{c_{\mathrm{m}}}{2}}\nu_{R\mathbb{D}}\to\nu_{\widehat{\mathbb{C}}}$ weakly as $R\to\infty$.
Proof.
By [Zha21, p385], $|\log\epsilon|^{\frac{c_{\mathrm{m}}}{2}}\nu_{\epsilon\mathbb{D}^{*}}\to\nu_{\widehat{\mathbb{C}}}$ weakly as $\epsilon\to 0$. The result then follows by taking an inversion and by conformal invariance of $\nu_{\widehat{\mathbb{C}}}$. ∎
Lemma 4.11.
Let $\eta\in\mathcal{J}_{0,\infty}$ and $e^{-\tilde{c}}g:\mathbb{D}^{*}\to\mathbb{C}\setminus\overline{\mathrm{int}(\eta)}$ be a conformal map uniformising the exterior of $\eta$. Let $R_{0}>0$ be such that $\overline{\mathrm{int}(\eta)}\subset R_{0}\mathbb{D}$. For $R\geq R_{0}$, let $\psi_{\eta,R}:R\mathbb{D}\setminus\overline{\mathrm{int}(\eta)}\to\mathbb{A}_{q(\eta,R)}$, for some $q(\eta,R)\in(0,1)$, be the conformal map with $q(\eta,R)^{-1}\psi_{\eta,R}\circ(e^{-\tilde{c}}g)(1)=1$. Then, $q(\eta,R)^{-1}\psi_{\eta,R}\circ(e^{-\tilde{c}}g)\to\mathrm{Id}_{\mathbb{D}^{*}}$ uniformly on $\overline{\mathbb{D}^{*}}$ as $R\to\infty$. More precisely, $q(\eta,R)^{-1}\psi_{\eta,R}\circ(e^{-\tilde{c}}g)$ extends to a quasiconformal diffeomorphism of $\overline{\mathbb{D}^{*}}$ that converges uniformly.
Proof.
By conjugating with an inversion, it is equivalent to proving the following statement:
Let $\eta$ be a bounded Jordan curve containing 0 in its interior, and $f:\mathbb{D}\to\mathrm{int}(\eta)$ a conformal map. Let $r_{0}$ such that $r_{0}\bar{\mathbb{D}}\subset\mathrm{int}(\eta)$ and $\psi_{r}:\mathrm{int}(\eta)\setminus r\bar{\mathbb{D}}\to\mathbb{A}_{q(r)}$ be the conformal such that $\psi_{r}\circ f(1)=1$, for all $0<r<r_{0}$. Then, $\psi_{r}\circ f\to\mathrm{id}_{\mathbb{D}}$ uniformly on $\bar{\mathbb{D}}$ as $r\to 0$.
Assume without loss of generality that $f\in\mathcal{E}$. The map $\phi_{r}:=\psi_{r}\circ f$ restricts to an analytic diffeomorphism of $\mathbb{S}^{1}$, so it extends analytically in a neighbourhood of $\mathbb{S}^{1}$ in $\mathbb{D}^{*}$ by Schwarz reflection. Pick any smooth quasiconformal extension of $\phi_{r}$ to $\bar{\mathbb{D}}$ and keep denoting it the same way. We can choose this quasiconformal extension so that it fixes 0.
First, we give a (very rough) bound on $q(r)$. Let $C:=\max\{|f(z)|\leavevmode\nobreak\ |\,z\in\mathbb{S}^{1}\}$. By the Koebe $1/4$-theorem, we have $\frac{1}{4}\in\mathrm{int}(\eta)$. Fix $c>0$ such that $\psi_{r}(1/4)\geq c$ for all $r>0$. On the one hand, the harmonic measure of $r\mathbb{S}^{1}$ viewed from $\frac{1}{4}$ in $\mathrm{int}(\eta)$ is upper-bounded by $\log(1/4C)/\log(r/C)$. On the other hand, the harmonic measure of $q(r)\mathbb{S}^{1}$ viewed from $c$ in $\mathbb{A}_{q(r)}$ is $\log c/\log q(r)$. Hence, by conformal invariance of harmonic measure
(4.20) $q(r)\leq\tilde{C}r^{\alpha},$
with $\tilde{C}=C^{\frac{\log c}{\log(4C)}}$ and $\alpha=-\frac{\log c}{\log(4C)}>0$.
Second, we adapt the proof of Montel’s theorem in order to extract a subsequence converging uniformly on compacts of $\mathbb{D}\setminus\{0\}$. Let $(z_{m})_{m\in\mathbb{N}}$ be a dense sequence in $\mathbb{D}\setminus\{0\}$. Since $\phi_{r}$ is uniformly bounded by 1, a standard diagonal argument shows that there exists a sequence $(\phi(z_{m}))_{m\in\mathbb{N}}$ and a a subsequence $(\phi_{r_{n}})$ such that $\phi_{r_{n}}(z_{m})\to\phi(z_{m})$ simultaneously for all $m$ as $n\to\infty$.
By Koebe $1/4$-theorem, $r\mathbb{D}\subset f(4\mathbb{D})$ so $\phi_{r}$ is defined on $\mathbb{A}_{4r}$. By Cauchy’s integral formula, we have for all $z,w\in\mathbb{A}_{8r}$,
$|\phi_{r}(z)-\phi_{r}(w)|\leq C^{\prime}r^{-2}|z-w|,$
for some $C^{\prime}>0$. This implies that $\phi$ extends uniquely to a continuous function on $\mathbb{D}\setminus\{0\}$, and the convergence $\phi_{r_{n}}\to\phi$ is uniform on compacts of $\mathbb{D}\setminus\{0\}$. In particular, $\phi$ is holomorphic on $\mathbb{D}\setminus\{0\}$. Moreover, the first step gives $\phi(z)\to 0$ as $z\to 0$, so $\phi$ extends analytically on $\mathbb{D}$.
Since $\phi$ is not constant, it restricts to an analytic diffeomorphism of $\mathbb{S}^{1}$ as a uniform limit of analytic diffeomorphisms. Moreover, $\phi$ extends to the whole plane by Schwarz reflection, with a unique pole at $\infty$. Hence, $\phi$ is affine. Since $\phi$ fixes 1 and 0, we have $\phi=\mathrm{id}_{\mathbb{C}}$.
Finally, by uniqueness of the subsequential limit, we deduce that $\psi_{\epsilon}\circ f$ converges uniformly on $\bar{\mathbb{D}}$ to the identity as $\epsilon\to 0$. ∎
We now have all the ingredients to prove Theorem 4.1.
Proof of Theorem 4.1.
Let us first assume that (4.1) holds for all holomorphic vector fields $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$. Let $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$. Applying $\Theta$ to the left and to the right of the first identity in (4.1) and using Lemma 4.4, we obtain that
¯$-\bar{\mathcal{L}}_{\mathrm{v}^{*}}^{*}=\mathcal{L}_{\mathrm{v}^{*}}-\frac{c_{\mathrm{m}}}{12}\vartheta(\mathrm{v}^{*}).$
This is the second identity in (4.1) for $\mathrm{v}^{*}$. Since $\mathrm{v}\mapsto\mathrm{v}^{*}$ maps $\mathbb{C}[z]\partial_{z}$ to $z^{2}\mathbb{C}[1/z]\partial_{z}$ and by linearity, this shows that the second identity in (4.1) holds for all vector fields in $\mathbb{C}(z)\partial_{z}$. The case of the first identity of (4.1) is symmetric and it remains to prove (4.1) for $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$. We focus on the first identity.
Let $\mathrm{v}\in\mathbb{C}[z]\partial_{z}$, $F,G\in\mathcal{C}$ and $R>0$. For any Jordan curve $\eta$ contained in $R\mathbb{D}$, we will denote by $\psi_{\eta,R}$ the conformal map introduced in Lemma 4.11 which sends $R\mathbb{D}\setminus\overline{\mathrm{int}(\eta)}$ to $\mathbb{A}_{q(\eta,R)}$. By scaling and Proposition 4.6,
$(\mathcal{L}_{\mathrm{v}}F,G)_{L^{2}(\nu_{R\mathbb{D}})}$ $=\Big{(}F,-\bar{\mathcal{L}}_{\mathrm{v}}G-\Big{(}\frac{c_{\mathrm{m}}}{12}\overline{(\mathcal{S}\psi_{\eta,R},\mathrm{v})}+\frac{c_{\mathrm{m}}}{2}U(q(\eta,R))\overline{(\psi_{\eta,R}^{-2}{\psi_{\eta,R}^{\prime}}^{2},\mathrm{v})}\hskip 1.0pt\Big{)}G\Big{)}_{L^{2}(\nu_{R\mathbb{D}})}.$
Since $\mathcal{S}\psi_{\eta,R}=\mathcal{S}(q(\eta,R)^{-1}\psi_{\eta,R})$, Lemma 4.11 implies that $(\mathcal{S}\psi_{\eta,R},\mathrm{v})\to(\mathcal{S}((e^{-\tilde{c}}g_{\eta})^{-1}),\mathrm{v})$ in $\mathcal{C}^{0}(\mathcal{J}_{0,\infty})$. By Lemma 4.10, $(\log R)^{\frac{c}{2}}\nu_{R\mathbb{D}}\to\nu_{\widehat{\mathbb{C}}}$ weakly. So $(\mathcal{S}\psi_{\eta,R},\mathrm{v})(\log R)^{\frac{c_{\mathrm{m}}}{2}}\nu_{R\mathbb{D}}\to(\mathcal{S}((e^{-\tilde{c}}g_{\eta})^{-1}),\mathrm{v})$ weakly. Concerning the term featuring $U(q(\eta,R))$, and because $\lim_{q\to 0^{+}}U(q)=0$, we similarly have the weak convergence $U(q(\eta,R))(\psi_{\eta,R}^{-2}{\psi_{\eta,R}^{\prime}}^{2},\mathrm{v})(\log R)^{\frac{c_{\mathrm{m}}}{2}}\nu_{R\mathbb{D}}\to 0$. Altogether, this shows that
$(\log R)^{\frac{c_{\mathrm{m}}}{2}}(\mathcal{L}_{\mathrm{v}}F,G)_{L^{2}(\nu_{R\mathbb{D}})}\xrightarrow[R\to\infty]{}\Big{(}F,-\bar{\mathcal{L}}_{\mathrm{v}}G-\frac{c_{\mathrm{m}}}{12}\overline{(\mathcal{S}(e^{-\tilde{c}}g_{\eta})^{-1},\mathrm{v})}G\Big{)}_{L^{2}(\nu_{\widehat{\mathbb{C}}})}.$
This concludes the proof that
$(\mathcal{L}_{\mathrm{v}}F,G)_{L^{2}(\nu_{\widehat{\mathbb{C}}})}=\Big{(}F,-\bar{\mathcal{L}}_{\mathrm{v}}G-\frac{c_{\mathrm{m}}}{12}\overline{(\mathcal{S}(e^{-\tilde{c}}g_{\eta})^{-1},\mathrm{v})}G\Big{)}_{L^{2}(\nu_{\widehat{\mathbb{C}}})}.$
Finally, by Proposition 2.6 and Lemma 2.10, $\mathcal{L}_{\mathrm{v}}^{*}$ and ¯$-\bar{\mathcal{L}}_{\mathrm{v}}-\frac{c_{\mathrm{m}}}{12}\overline{\tilde{\vartheta}(\mathrm{v})}$ are endomorphisms of $\mathcal{C}_{\mathrm{comp}}$. Since $\mathcal{C}_{\mathrm{comp}}$ is dense in $L^{2}(\nu)$ (see Lemma 2.4), this implies the closability of $\mathcal{L}_{\mathrm{v}}$ [RS80, Theorem VIII.1.(b)]. ∎
5. Spectral resolution
This section is devoted to the spectral resolution of the Shapovalov form. We will in particular prove Theorem 1.7. We first record the elementary relation between the electrical thickness and the conformal radius viewed from $\infty$.
Lemma 5.1.
The coordinate $-\tilde{c}_{0}$ on $\mathcal{E}$ coincides with the electrical thickness of [AS21].
Proof.
The electrical thickness is defined on $\mathcal{E}$ as follows. Let $f\in\mathcal{E}$ and fix the unique $\delta>0$ such that $e^{\delta}f(\mathbb{S}^{1})\subset\overline{\mathbb{D}^{*}}$ and $f(\mathbb{S}^{1})\cap\mathbb{S}^{1}\neq\eset$. Then, the interior of the Jordan curve $e^{-\delta}/f(\mathbb{S}^{1})$ is uniformised by $z\mapsto e^{\tilde{c}_{0}-\delta}/g(1/z)$. In [AS21], the electrical thickness is defined to be $-\delta-(\tilde{c}_{0}-\delta)=-\tilde{c}_{0}$. ∎
Proof of Theorem 1.7.
Let $F,G\in\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$, $\mathbf{k},\mathbf{k}^{\prime},\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime}\in\mathcal{T}$ with $|\mathbf{k}|=|\mathbf{k}^{\prime}|$ and $|\tilde{\mathbf{k}}|=|\tilde{\mathbf{k}}^{\prime}|$ and $\lambda\in\mathbb{C}$ such that $02\mathrm{Re}(\lambda)+|\mathbf{k}|+|\tilde{\mathbf{k}}|+\kappa/8-1<$ 0. Let us first show that
(5.1) $\mathcal{Q}\left(e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}F\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},e^{-(2\bar{\lambda}+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}G\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)=\frac{1}{\sqrt{\pi}}\int_{\mathbb{R}}\widehat{F}(p)\overline{\widehat{G}(-p)}\mathcal{Q}_{\lambda+ip}\left(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\lambda,\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)\mathrm{d}p.$
We have $e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}F\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},e^{-(2\bar{\lambda}+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}G\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\in\mathcal{C}$, so the Hermitian form $\mathcal{Q}$ evaluated at these two elements is well defined and finite. By definition (1.7) of $\mathcal{Q}$ and decomposing $\nu=\pi^{-1/2}\mathrm{d}c\otimes\nu^{\#}$, we have
$\mathcal{Q}\left(e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}F\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},e^{-(2\bar{\lambda}+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}G\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)$
$=\int_{\mathbb{R}}e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}F(c)\mathbb{E}^{\#}\left[e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)\tilde{c}}\overline{G(\tilde{c})}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}(f)\overline{\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}(\iota\circ g\circ\iota)}\right]\frac{\mathrm{d}c}{\sqrt{\pi}}.$
Using that $\tilde{c}=\tilde{c}_{0}-c$ and a change of variables, this is further equal to
$\mathbb{E}^{\#}\left[\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}(f)\overline{\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}(\iota\circ g\circ\iota)}\int_{\mathbb{R}}F(c)\overline{G(\tilde{c}_{0}-c)}e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)\tilde{c}_{0}}\frac{\mathrm{d}c}{\sqrt{\pi}}\right]$
$=\mathbb{E}^{\#}\left[\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}(f)\overline{\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}(\iota\circ g\circ\iota)}\int_{\mathbb{R}}\widehat{F}(p)\overline{\widehat{G}(-p)}e^{-(2ip+2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)\tilde{c}_{0}}\frac{\mathrm{d}p}{\sqrt{\pi}}\right].$
Now, since $02\mathrm{Re}(\lambda)+|\mathbf{k}|+|\tilde{\mathbf{k}}|+\kappa/8-1<$ 0, (1.15) guarantees that the last line in the above display is integrable. Hence, we can apply Fubini to get (5.1).
We are now going to prove (1.19) by computing
(5.2) $\mathcal{Q}\left(\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F),\mathbf{L}_{-\mathbf{k}^{\prime}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}^{\prime}}(e^{-2\bar{\lambda}c}G)\right)$
in two ways. First, since $\mathcal{P}_{\lambda+ip}\mathds{1}=e^{-2(\lambda+ip)c}$ for all $p\in\mathbb{R}$, we can write in $L^{2}(\mathbb{R},\mathrm{d}c)$ by Fourier inversion:
$e^{-2\lambda c}F(c)=\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\leavevmode\nobreak\ \widehat{F}(p)e^{-2(ip+\lambda)c}=\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\leavevmode\nobreak\ \widehat{F}(p)\mathcal{P}_{\lambda+ip}\mathds{1}.$
Using that $\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}\mathcal{P}_{\lambda+ip}=\mathcal{P}_{\lambda+ip}\mathbf{L}_{-\mathbf{k}}^{\lambda+ip}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}^{\lambda+ip}$ (consequence of the conjugation relation of Proposition 3.3) and dominated convergence theorem, we infer that
(5.3) $\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F)=\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\leavevmode\nobreak\ \widehat{F}(p)\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}\mathcal{P}_{\lambda+ip}\mathds{1}=\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\leavevmode\nobreak\ \widehat{F}(p)\mathcal{P}_{\lambda+ip}\mathbf{L}^{\lambda+ip}_{-\mathbf{k}}\bar{\mathbf{L}}^{\lambda+ip}_{-\tilde{\mathbf{k}}}\mathds{1}.$
By Proposition 3.3, $\mathbf{L}^{\lambda+ip}_{-\mathbf{k}}\bar{\mathbf{L}}^{\lambda+ip}_{-\tilde{\mathbf{k}}}\mathds{1}=\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}}$. By definition of $\mathcal{P}_{\lambda+ip}=e^{-(\mathbf{L}_{0}^{\lambda+ip}+\bar{\mathbf{L}}_{0}^{\lambda+ip})}$ and by (3.3), $\mathcal{P}_{\lambda+ip}\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}}=e^{-(2\lambda+2ip+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}}$. Altogether, this shows that
$\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F)=e^{-(2\lambda+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\leavevmode\nobreak\ \widehat{F}(p)e^{-2ipc}\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}}.$
Since the Fourier transform of
$c\mapsto\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\leavevmode\nobreak\ \widehat{F}(p)e^{-2ipc}\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}}\qquad\text{is given by}\qquad p\mapsto\widehat{F}(p)\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}},$
an interpolation of (5.1) implies that (5.2) is equal to
(5.4) $\int_{\mathbb{R}}\widehat{F}(p)\overline{\widehat{G}(-p)}\mathcal{Q}_{\lambda+ip}\left(\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda}+ip,\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)\frac{\mathrm{d}p}{\sqrt{\pi}}.$
To get a second expression for (5.2), we first use the integration by parts of Theorem 1.3:
$\mathcal{Q}\left(\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F),\mathbf{L}_{-\mathbf{k}^{\prime}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}^{\prime}}(e^{-2\bar{\lambda}c}G)\right)=\mathcal{Q}\left(\mathbf{L}_{\mathbf{k}^{\prime}}\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{\tilde{\mathbf{k}}^{\prime}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F),e^{-2\bar{\lambda}c}G\right).$
Similarly to (5.3), we have
$\mathbf{L}_{\mathbf{k}^{\prime}}\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{\tilde{\mathbf{k}}^{\prime}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F)=\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\,\widehat{F}(p)\mathcal{P}_{\lambda+ip}\mathbf{L}^{\lambda+ip}_{\mathbf{k}^{\prime}}\mathbf{L}^{\lambda+ip}_{-\mathbf{k}}\bar{\mathbf{L}}^{\lambda+ip}_{\tilde{\mathbf{k}}^{\prime}}\bar{\mathbf{L}}^{\lambda+ip}_{-\tilde{\mathbf{k}}}\mathds{1}.$
By definition of the Gram matrix, $\mathbf{L}^{\lambda+ip}_{\mathbf{k}^{\prime}}\mathbf{L}^{\lambda+ip}_{-\mathbf{k}}\bar{\mathbf{L}}^{\lambda+ip}_{\tilde{\mathbf{k}}^{\prime}}\bar{\mathbf{L}}^{\lambda+ip}_{-\tilde{\mathbf{k}}}\mathds{1}=\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})\mathds{1}$ and, because $\mathcal{P}_{\lambda+ip}\mathds{1}=e^{-2(\lambda+ip)c}$, we find that
$\mathbf{L}_{\mathbf{k}^{\prime}}\mathbf{L}_{-\mathbf{k}}\bar{\mathbf{L}}_{\tilde{\mathbf{k}}^{\prime}}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}(e^{-2\lambda c}F)=e^{-2\lambda c}\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\,\widehat{F}(p)e^{-2ipc}\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime}).$
Noting that the Fourier transform of
$c\mapsto\int_{\mathbb{R}}\frac{\mathrm{d}p}{\sqrt{\pi}}\,\widehat{F}(p)e^{-2ipc}\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})\quad\text{is given by}\quad p\mapsto\widehat{F}(p)\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime}),$
we can conclude as before, using (5.1), that (5.2) equals
(5.5) $\int_{\mathbb{R}}\widehat{F}(p)\overline{\widehat{G}(-p)}\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})\mathcal{Q}_{\lambda+ip}(\mathds{1},\mathds{1})\frac{\mathrm{d}p}{\sqrt{\pi}}.$
Recalling that $\mathrm{R}(\lambda+ip)=\mathcal{Q}_{\lambda+ip}(\mathds{1},\mathds{1})$ and because (5.4) and (5.5) agree for all $F,G\in\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$, we get for Lebesgue-almost every $p\in\mathbb{R}$,
$\mathcal{Q}_{\lambda+ip}(\Psi_{\lambda+ip,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda}-ip,\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})=\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})\mathrm{R}(\lambda+ip).$
Finally, both functions are continuous (in fact, analytic) in $p\in\mathbb{R}$ since we are in the region of absolute convergence. Thus, the equality holds for all $p\in\mathbb{R}$. In the case where the $\lambda$-levels $(|\mathbf{k}|,|\tilde{\mathbf{k}}|)$ and $(|\mathbf{k}^{\prime}|,|\tilde{\mathbf{k}}^{\prime}|)$ are different, the proof is by all means identical and we find zero. This concludes the proof of (1.19).
Concerning the proof of (1.17), combining (5.1) and (1.19) we see that (1.17) holds for all $\lambda\in\mathbb{C}$ with $02\mathrm{Re}(\lambda)+|\mathbf{k}|+|\tilde{\mathbf{k}}|+\kappa/8-1<$ 0, i.e. that for such values of $\lambda$,
$\mathcal{Q}\left(F\mathcal{P}_{\lambda}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},G\mathcal{P}_{\bar{\lambda}}\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)=\int_{\mathbb{R}}\widehat{F}(p)\overline{\widehat{G}(-p)}\mathrm{R}(\lambda+ip)\mathrm{B}_{\lambda+ip}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda+ip}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime})\frac{\mathrm{d}p}{\sqrt{\pi}}.$
We conclude by noticing that the LHS is analytic in the whole $\lambda$-plane, and the RHS is analytic in the region $\mathrm{Re}(\lambda)<\frac{1}{2}(1-\frac{\kappa}{8})$, so both sides coincide in this region. (In particular, the RHS analytically continues to the whole $\lambda$-plane too, but the analytic continuation of this integral includes some additional contributions when the poles of $\mathrm{R}(\lambda)$ cross the integration line.)
It remains to prove (1.18). The Fourier transform of any $F\in\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$ extends analytically to the whole $p$-plane. Moreover, $\widehat{F}$ decays faster than any polynomials as $|p|\to\infty$. Now, let $F\in\mathcal{C}_{\mathrm{comp}}$, and recall that $(\Psi_{ip,\mathbf{k},\tilde{\mathbf{k}}})_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}$ is a basis of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ for all $p\in\mathbb{R}$ (Theorem 1.4). Hence, the Fourier transform $\widehat{F}(p,\cdot)$ decomposes on this basis for each $p\in\mathbb{R}$, i.e. we can write
$F=\int_{\mathbb{R}}\sum_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}\alpha_{F}(p,\mathbf{k},\tilde{\mathbf{k}})e^{-2ipc}\Psi_{ip,\mathbf{k},\tilde{\mathbf{k}}}\frac{\mathrm{d}p}{\sqrt{\pi}},$
where $p\mapsto\alpha_{F}(p,\mathbf{k},\tilde{\mathbf{k}})$ is analytic for all $\mathbf{k},\mathbf{k}^{\prime}\in\mathcal{T}$ and decays faster than any polynomial as $|p|\to\infty$. Moreover, $\alpha_{F}(\,\cdot,\mathbf{k},\mathbf{k}^{\prime})=0$ for $|\mathbf{k}|+|\tilde{\mathbf{k}}|$ large enough ($F(p,\cdot)$ is a polynomial). Let $\mathcal{P}:\mathcal{C}_{\mathrm{comp}}\to\mathcal{C}_{\mathrm{comp}}$ be the operator defined by
$\mathcal{P}F:=\int_{\mathbb{R}}\sum_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}\alpha_{F}(p,\mathbf{k},\tilde{\mathbf{k}})e^{-(2ip+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{ip,\mathbf{k},\tilde{\mathbf{k}}}\frac{\mathrm{d}p}{\sqrt{\pi}}.$
(Formally, we have $\mathcal{P}=\int_{\mathbb{R}}\mathcal{P}_{ip}\frac{\mathrm{d}p}{\sqrt{\pi}}$.) This operator is clearly invertible, and by standard properties of the Laplace transform, we have $\alpha_{\mathcal{P}^{-1}F}(p,\mathbf{k},\tilde{\mathbf{k}})=\alpha_{F}(p-\frac{i}{2}(|\mathbf{k}|+|\tilde{\mathbf{k}}|),\mathbf{k},\tilde{\mathbf{k}})=:\tilde{\alpha}_{F}(p,\mathbf{k},\tilde{\mathbf{k}})$. Hence, we have in $L^{2}(\nu)$
$F=\int_{\mathbb{R}}\sum_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}\tilde{\alpha}_{F}(p,\mathbf{k},\tilde{\mathbf{k}})e^{-(2ip+|\mathbf{k}|+|\tilde{\mathbf{k}}|)c}\Psi_{ip,\mathbf{k},\tilde{\mathbf{k}}}\frac{\mathrm{d}p}{\sqrt{\pi}}.$
It just remains to express $\tilde{\alpha}(p,\mathbf{k},\tilde{\mathbf{k}})$ in terms of the pairings $\mathcal{Q}(F,e^{(2ip-|\mathbf{k}|-|\tilde{\mathbf{k}}|)c}\Psi_{-ip,\mathbf{k},\tilde{\mathbf{k}}})$. Note that these pairings are well defined since $F\in\mathcal{C}_{\mathrm{comp}}$. Moreover, for $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$ fixed and $\mathrm{Re}(\lambda)<\frac{1}{2}(1-|\mathbf{k}|-\tilde{\mathbf{k}}-\frac{\kappa}{8})$, we have by (1.17)
$\mathcal{Q}(F,\mathcal{P}_{\bar{\lambda}}\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})$ $=\sum_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}\tilde{\alpha}_{F}(-i\lambda,\mathbf{k},\tilde{\mathbf{k}})\mathcal{Q}_{\lambda}(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\Psi_{\bar{\lambda},\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}})$
$=\sum_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}\tilde{\alpha}_{F}(-i\lambda,\mathbf{k},\tilde{\mathbf{k}})\mathrm{R}(\lambda)\mathrm{B}_{\lambda}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{\lambda}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime}).$
[…]
$\tilde{\alpha}_{F}(p,\mathbf{k},\tilde{\mathbf{k}})=\frac{1}{\mathrm{R}(ip)}\sum_{\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}\in\mathcal{T}}\mathcal{Q}\left(F,\mathcal{P}_{-ip}\Psi_{-ip,\mathbf{k}^{\prime},\tilde{\mathbf{k}}^{\prime}}\right)\mathrm{B}_{ip}^{-1}(\mathbf{k},\mathbf{k}^{\prime})\mathrm{B}_{ip}^{-1}(\tilde{\mathbf{k}},\tilde{\mathbf{k}}^{\prime}),$
concluding the proof. ∎
6. Uniqueness of restriction measures
In this section, we prove the uniqueness of restriction measures as stated in Theorem 1.9. Before moving to the actual proof, we would like to remind the reader that the representations $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ and the map $\Theta$ are universal, and so are the representations $(\mathcal{W}_{\lambda},(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})_{n\in\mathbb{Z}})$. What may depend on the measure are the properties determined by the associated $L^{2}$-product (and the integration by parts formula), in particular the adjoint relations $\Theta^{*}=\Theta$, $\mathbf{L}_{n}^{*}=\Theta\circ\mathbf{L}_{-n}\circ\Theta$ etc.
Proof of Theorem 1.9.
The existence part is already known [Zha21], so we only need to show uniqueness. A conformal invariant Borel measure on the space $\mathcal{J}$ is completely characterised by its restriction to $\mathcal{J}_{0,\infty}$. Since the restriction to $\mathcal{J}_{0,\infty}$ of a restriction measure on $\mathcal{J}$ (in the sense of Definition 1.8) is a weak restriction measure on $\mathcal{J}_{0,\infty}$, it is enough to prove uniqueness of weak restriction measures on $\mathcal{J}_{0,\infty}$. Let $\tilde{\nu}$ be such a Borel measure on $\mathcal{J}_{0,\infty}$. We first claim that $\tilde{\nu}$ is scale invariant, i.e. $\tilde{\nu}(\lambda\cdot)=\tilde{\nu}$ for $\lambda>0$. It is enough to check that for any $\lambda,R>0$, $\tilde{\nu}(\lambda\cdot)\mathds{1}_{\cdot\subset R\mathbb{D}}=\tilde{\nu}(\cdot)\mathds{1}_{\cdot\subset R\mathbb{D}}$. This equality is a consequence of the fact that $\tilde{\nu}$ is a weak restriction measure and the scale invariance of $\Lambda^{*}$, proving the desired claim. Now, scale invariance of $\tilde{\nu}$ implies that (up to rescaling by a multiplicative constant) $\tilde{\nu}=\pi^{-1/2}\mathrm{d}c\otimes\tilde{\nu}^{\#}$ for some probability measure $\tilde{\nu}^{\#}$ on $\mathcal{E}$. Let $\tilde{\mathbb{E}}^{\#}$ be the expectation w.r.t. $\tilde{\nu}^{\#}$. At this stage, we do not know that $\nu$ is invariant under $z\mapsto 1/\bar{z}$. So we do not know yet that $\Theta$ is self-adjoint w.r.t. $L^{2}(\tilde{\nu})$ and that $\tilde{\mathcal{Q}}(F,G):=\langle F,\Theta G\rangle_{L^{2}(\tilde{\nu})}$ is a Hermitian form on $L^{2}(\tilde{\nu})$. This will be a consequence of the uniqueness we are currently proving.
By definition, the family of measures $\mathrm{d}\tilde{\nu}_{D}(\eta):=\mathds{1}_{\eta\subset D}e^{\frac{c_{\mathrm{m}}}{2}\Lambda^{*}(\eta,\partial D)}\mathrm{d}\tilde{\nu}(\eta)$ indexed by simply connected domains $D\subset\mathbb{C}$ containing 00 is conformally invariant. Now, the conclusion of Proposition 4.6 remains true if we replace $\nu_{\mathbb{D}}$ by $\tilde{\nu}_{\mathbb{D}}$ (the proof only uses this conformal restriction property), so $\tilde{\nu}_{\mathbb{D}}$ satisfies the integration by parts formula of Proposition 4.6. To get an integration by parts in the sphere, we proceed as before by noting that $|\log\epsilon|^{\frac{c_{\mathrm{m}}}{2}}\tilde{\nu}_{\epsilon\mathbb{D}^{*}}\to\tilde{\nu}$ weakly as $\epsilon\to 0$. Indeed, by the restriction property of $\tilde{\nu}$,
$\tilde{\nu}=\lim_{\epsilon\to 0}\mathds{1}_{\eta\subset\epsilon\mathbb{D}^{*}}\mathrm{d}\tilde{\nu}(\eta)=\lim_{\epsilon\to 0}e^{-\frac{c_{\mathrm{m}}}{2}\Lambda^{*}(\eta,\epsilon\mathbb{S}^{1})}\mathrm{d}\tilde{\nu}_{\epsilon\mathbb{D}^{*}}(\eta)$
and $\Lambda^{*}(\eta,\epsilon\mathbb{S}^{1})=-\log|\log\epsilon|+o(1)$ (this term does not depend on the specific choice of weak restriction measure on $\mathcal{J}_{0,\infty}$ and is thus the same as the one in [Zha21, page 385]). It follows that $\tilde{\nu}$ also satisfies the integration by parts formula of Theorem 4.1, i.e. we have the relations $\mathbf{L}_{n}^{*}=\Theta\circ\mathbf{L}_{-n}\circ\Theta$ and $\bar{\mathbf{L}}_{n}^{*}=\Theta\circ\bar{\mathbf{L}}_{-n}\circ\Theta$ for all $n\in\mathbb{Z}$, where the adjoint is w.r.t. to the $L^{2}(\tilde{\nu})$-inner product. In particular, $\tilde{\mathcal{Q}}$ satisfies the conclusions of Theorem 1.7 with $\tilde{\mathrm{R}}(\lambda):=\tilde{\mathbb{E}}^{\#}[e^{-2\lambda\tilde{c}_{0}}]$ in place of $\mathrm{R}(\lambda)$.
It is a simple exercise to check that the Kac table is bounded from below by $-\frac{(\kappa-4)^{2}}{16\kappa}$. In the sequel, we fix some $\lambda<-\frac{(\kappa-4)^{2}}{16\kappa}<0$, and note that $\tilde{\mathrm{R}}(\lambda):=\tilde{\mathbb{E}}^{\#}[e^{-2\lambda\tilde{c}_{0}}]<\infty$ since $\tilde{c}_{0}\leq 0$ a.s. (see e.g. [Pom75, Equation (21)]). Since $\lambda\not\in kac$, the space $\mathcal{W}_{\lambda}$ contains all polynomials, i.e. $(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}})_{\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}}$ is a linear basis of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$. Then, by (1.19) in Theorem 1.7, for all $\mathbf{k},\tilde{\mathbf{k}}\in\mathcal{T}$ with $|\mathbf{k}|+|\tilde{\mathbf{k}}|\geq 1$,
(6.1) $\tilde{\mathbb{E}}^{\#}[e^{-2\lambda\tilde{c}_{0}}\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}]=\tilde{\mathcal{Q}}_{\lambda}(\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}},\mathds{1})=\tilde{\mathrm{R}}(\lambda)\mathrm{B}_{\lambda}(\mathbf{k},\eset)\mathrm{B}_{\lambda}(\tilde{\mathbf{k}},\eset)=\tilde{\mathrm{R}}(\lambda)\delta_{\mathbf{k},\eset}\delta_{\tilde{\mathbf{k}},\eset}.$
The space $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$ is dense in $\mathcal{C}^{0}(\mathcal{E})$ (see the proof of Lemma 2.4) and $\mathcal{E}$ is a Radon space, so we can conclude by the Riesz–Markov representation theorem (e.g. [Rud87, Theorem 2.14]) that (6.1) characterises uniquely the measure $e^{-2\lambda\tilde{c}_{0}}\tilde{\nu}^{\#}$ in the space of Borel measures on $\mathcal{E}$, up to the global multiplicative constant $\tilde{\mathrm{R}}(\lambda)$. Namely, $e^{-2\lambda\tilde{c}_{0}}\tilde{\nu}^{\#}=Ce^{-2\lambda\tilde{c}_{0}}\nu^{\#}$ for some $C>0$. It follows that $\tilde{\nu}^{\#}=C\nu^{\#}$ and $C=1$ since both are probability measures. ∎
Remark 6.1.
The operators $\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda}$ (3.4) are well defined for all complex values of $c_{\mathrm{m}}$ and form two commuting representations of the Virasoro algebra with central charge $c_{\mathrm{m}}$. From these operators and as before, we can then define $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}=\mathbf{L}_{-\mathbf{k}}^{\lambda}\bar{\mathbf{L}}_{-\tilde{\mathbf{k}}}^{\lambda}\mathds{1}$. For $\lambda$ not in the Kac table, the Riesz-Markov representation theorem shows that the assignment $\Psi_{\lambda,\mathbf{k},\tilde{\mathbf{k}}}\to\delta_{(\mathbf{k},\tilde{\mathbf{k}})=(\eset,\eset)}$ extends uniquely to a complex Borel measure $\nu^{\#}_{c_{\mathrm{m}}}$ on $\mathcal{E}$ (recall (6.1)). On the other hand, Definition 1.8 also makes sense for all complex $c_{\mathrm{m}}$, and following the same route as in our proof of uniqueness, such a measure (rather, its shape) is determined by the above measure $\nu^{\#}_{c_{\mathrm{m}}}$. So, if such a measure exists, it is unique. With some additional effort, it should be possible to prove the existence as well, i.e. that $\nu^{\#}_{c_{\mathrm{m}}}$ satisfies the properties of Definition 1.8. Of course, the existence of the SLE loop measure grants for free that the existence and positivity for $c_{\mathrm{m}}\leq 1$. For non-real $c_{\mathrm{m}}$, it is clear that a restriction measure cannot be positive. Finally, we believe that positivity fails for $c_{\mathrm{m}}>1$ (due to positivity properties of the Shapovalov form in this region, it is very plausible that an odd function of $\Theta$ would have negative “$L^{2}$-norm”).
7. Applications
We now discuss two consequences of the characterisation of the SLE loop measure in terms of restriction property.
7.1. Reversibility
The reversibility of the SLE loop measure is its invariance under the map $\iota:z\mapsto 1/\bar{z}$, which is present in Zhan’s axiomatisation. This property is not obvious a priori: in the chordal case for instance, Schramm’s definition of SLE only uses scale invariance [Sch00], but not the full Möbius invariance. In fact, the key property of invariance under inversion $z\mapsto 1/z$ was not part of the axioms and remained to be proved. This proof came only several years later, in [Zha08b] for the simple case $\kappa\in(0,4]$, and in [MS16c] for the non-simple case $\kappa\in(4,8)$.
Now, we explain how Theorem 1.9 provides an alternative and simpler approach to the question of reversibility. We remind the reader that our proof of uniqueness uses only the scale invariance of the SLE loop measure (which allows to write $\nu=\mathrm{d}c\otimes\nu^{\#}$), and not the full Möbius invariance (hence the argument is not recursive). By invariance of the Brownian loop measure under complex conjugation and Möbius inversion, it follows immediately from Definition 1.8 that $\iota^{*}(\mathrm{d}c\otimes\nu^{\#})$ is also a restriction measure, hence $\iota^{*}(\mathrm{d}c\otimes\nu^{\#})=K\mathrm{d}c\otimes\nu^{\#}$ for some $K>0$. To see that $K=1$, observe that for all $F\in\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$ with $\int_{\mathbb{R}}F(c)\mathrm{d}c=1$, we have
$1=\int_{\mathbb{R}}F(c)\mathrm{d}c=K\mathbb{E}^{\#}[\int F(\tilde{c})\mathrm{d}c]=K\mathbb{E}^{\#}[\int_{\mathbb{R}}F(c-\tilde{c}_{0})\mathrm{d}c]=K\mathbb{E}^{\#}[\int_{\mathbb{R}}F(c)\mathrm{d}c]=K\int_{\mathbb{R}}F(c)\mathrm{d}c=K.$
As mentioned to us by W. Werner, and before the proof of Zhan [Zha08b] (see also [MS16b] and [LY21]), such an approach for proving reversibility of SLE had been considered, but a characterisation of SLE in terms of restriction property was missing; see for instance [Wer05, Section 8.1]. Theorem 1.9 is this missing piece, at least in the loop context.
7.2. Duality
Our second application concerns the duality $\kappa\leftrightarrow 16/\kappa$ for SLE. The duality of chordal SLE was first conjectured by Duplantier (see also Dubédat [Dub05]) and later proved by Zhan [Zha08a], Dubédat [Dub09a] and Miller–Sheffield [MS16a]. Duality for the SLE loop, as stated in Theorem 7.1 below, appears to be new. We mention that the paper in preparation [ACSW] gives an independent proof of this result, based on conformal welding of quantum discs and Lévy excursion theory (private communication with the authors).
[…]
Theorem 7.1 (Duality).
[…]
Theorem 7.1 does not address the self-dual case $\kappa^{\prime}=\kappa=4$. In this case, the curve is almost surely simple, and the duality for SLE4 can be understood as the reversibility proved in Section 7.1.
By conformal restriction, Theorem 7.1 implies a similar result in any simply connected domain $D$, where one looks at the boundary of the connected component which contains the boundary $\partial D$.
Proof.
[…]
$\mathds{1}_{\eta^{\prime}\subset D}=\mathds{1}_{\eta\subset D}\quad\text{and}\quad\Lambda^{*}(\eta^{\prime},D^{c})=\Lambda^{*}(\eta,D^{c}).$
[…]
We have just proved that the shape measure of the outer boundary of $\eta^{\prime}$ is the SLEκ shape measure $\nu_{\kappa}^{\#}$. This implies that $\mu_{\kappa}=K\nu_{\kappa}$ for some global multiplicative constant $K>0$. Our choice of normalisation ensures that the distribution of the $\log$-conformal radius of $\eta$ viewed from $\infty$ is $\frac{\mathrm{d}\tilde{c}}{\sqrt{\pi}}$. This coincides with our choice of normalisation for $\nu_{\kappa}$, so indeed $K=1$. ∎
8. Future work and open questions
In this paper, we have studied the local/algebraic aspects of the CFT of SLE (Hilbert space, representation, Shapovalov form). In forthcoming work, we will study the global/geometric objects associated with this CFT. We use this section to outline this programme.
8.1. Correlation functions
The first step is to define $N$-point correlation functions on the Riemann sphere. Given $\mathbf{z}=(z_{1},...,z_{N})\in\widehat{\mathbb{C}}$ distinct marked points, one can use the Brownian loop measure in order to define a measure $\mathrm{d}\nu_{\widehat{\mathbb{C}},\mathbf{z}}=\mathcal{A}_{\widehat{\mathbb{C}},\mathbf{z}}\mathrm{d}\nu^{\otimes N}$ on $N$ pairwise disjoint SLEκ-type loops $(\eta_{1},...,\eta_{N})$ where $\eta_{j}$ is homotopic to $z_{j}$. The Radon-Nikodym derivative $\mathcal{A}_{\widehat{\mathbb{C}},\mathbf{z}}$ is the exponential of a normalised mass of Brownian loops. We call it the amplitude of (^$(\widehat{\mathbb{C}},\mathbf{z})$, following the terminology of [GKRV21].
Each loop $\eta_{j}$ can be represented by a pair $(c_{j},f_{j})\in\mathbb{R}\times\mathcal{E}$ as we did in this paper. Then, the correlation function is $\langle\prod_{j=1}^{N}\Psi_{\lambda_{j}}(z_{j})\rangle_{\widehat{\mathbb{C}},\mathbf{z}}:=\int\prod_{j=1}^{N}e^{-2\lambda_{j}c_{j}}\mathrm{d}\nu_{\widehat{\mathbb{C}},\mathbf{z}}$, which is finite for suitable values of $\lambda_{j}$. These correlation functions are conformally covariant. In particular, the three-point correlation function is determined by choosing $(z_{1},z_{2},z_{3})=(0,1,\infty)$ and we set $C_{\mathrm{SLE}}(\lambda_{1},\lambda_{2},\lambda_{3}):=\langle\Psi_{\lambda_{1}}(0)\Psi_{\lambda_{2}}(1)\Psi_{\lambda_{3}}(\infty)\rangle_{\widehat{\mathbb{C}},(0,1,\infty)}$.
The correlation functions will not converge in one of the two following cases: lack of integrability in the range of small conformal radius, or blow-up of the mass of Brownian loops when two loops get close together. An important question is the meromorphic continuation of these correlation functions, which is tightly linked to the distribution of Brownian loops.
8.2. Conformal Ward identities and BPZ equations
Using the result of this paper, the space $L^{2}(\nu^{\otimes N})$ carries $2N$ commuting representations of the Virasoro algebra $((\mathbf{L}_{n}^{(j)},\bar{\mathbf{L}}_{n}^{(j)})_{n\in\mathbb{Z}})_{1\leq j\leq N}$, where $\mathbf{L}_{n}^{(j)}$ acts as $\mathbf{L}_{n}$ on the $j^{\text{th}}$ tensor product and is the identity elsewhere. These operators are obtained using quasiconformal deformations generated by Beltrami differentials supported in a small neighbourhood of the points. In our follow-up work, we will prove the conformal Ward identities: for all meromorphic vector fields $\mathrm{v}$ on the Riemann sphere with poles only allowed at the marked points, we have
$\sum_{j=1}^{N}\mathbf{L}_{\mathrm{v}}^{(j)}\mathcal{A}_{\widehat{\mathbb{C}},\mathbf{z}}=0.$
As usual in CFT, this identity allows us to write the correlator of a descendant state as a partial differential operator (in the variables $(z_{1},...,z_{N})$) applied to the $\langle\prod_{j=1}^{N}\Psi_{\lambda_{j}}(z_{j})\rangle_{\widehat{\mathbb{C}},\mathbf{z}}$. In particular, considering the degenerate weight $\lambda_{r,s}$, the combination of Ward identities with Theorem 1.4 lead to the BPZ equations. In Liouville CFT, these equations (at level 2) are key in the proof of the DOZZ formula [KRV20]. It would be interesting to try to implement the same strategy for this model, as already suggested in [AS21, Quetion 1.4].
One key input for the implementation of this strategy is the fusion rules (a.k.a. operator product expansion), which is a structure of associative algebra on the Hilbert space (a structure expected in any CFT). The three-point amplitude $\mathcal{A}_{\widehat{\mathbb{C}},(0,1,\infty)}$ achieves this role (since a trilinear form can be viewed as a product). However, in order to exploit the information of degenerate fields, it is crucial to perform the analytic continuation of this product away from the spectrum, a point already mentioned in the previous subsection.
8.3. Conformal bootstrap
Suppose now that $N=4$ and $\mathbf{z}=(0,z,1,\infty)$. Conditionally on $(\eta_{j})_{1\leq j\leq 4}$, one can sample an SLEκ-loop $\eta$ in the domain ^$\widehat{\mathbb{C}}\setminus\cup_{j=1}^{4}\overline{\mathrm{int}(\eta_{j})}$, conditioned on the topological event that $\eta$ disconnects $(0,z)$ from $(1,\infty)$. The law of $\eta$ is also defined thanks to conformal restriction and the Brownian loop measure. On the other hand, conditionally on $\eta$, the law of the loops surrounding 0 and $z$ is that of a pair of SLEκ-loops in $\mathrm{int}(\eta)$, encoded in the amplitude $\mathcal{A}_{\mathrm{int}(\eta),(0,z)}$. A similar statement holds for the conditional law of the loops around 1 and $\infty$. Then, one can write
$\langle\prod_{j=1}^{4}\Psi_{\lambda_{j}}(z_{j})\rangle_{\widehat{\mathbb{C}},(0,z,1,\infty)}=\mathcal{Q}\left(\mathcal{A}_{\mathrm{int}(\cdot),(0,z)},\mathcal{A}_{\mathrm{int}(\cdot),(0,1)}\right).$
One can then use Theorem 1.7 to express the RHS on the Virasoro basis: this is the conformal bootstrap, which allows us to decompose the four-point function as a product of three-point functions and conformal blocks. This generalises easily to any $N\geq 4$.
8.4. Higher genus
The Brownian loop measure can also be used to define measures on loops embedded in a Riemann surface $\Sigma$, following a standard procedure [Law11, Wer08, Zha21]. Correlation functions are defined by considering loops homotopic to marked points. Note that we need a local coordinate at the marked point in order to define the Taylor series expansion of the embedding. In particular, the conformal radius depends on a 1-jet of local coordinate, and this dependence can be encoded in a line bundle over Teichmüller space (involving the conformal weight). Moreover, since the integration by parts formula involves a Schwarzian derivative, we see that the measure will depend on a choice of projective structure [^3] A projective structure is a choice of atlas such that transition functions are Möbius transformations (see e.g. [Dum09]). On such an atlas, the (locally defined) Schwarzian derivatives glue to a global object.. In fact, this dependence is explicit and can be encoded in a line bundle over Teichmüller space (involving the central charge). Thus, correlation functions are defined as sections of a line bundle over Teichmüller space, the construction of which is by all means similar to [BGKR].
In particular, for $N=1$ loop surrounding a single marked point $p\in\Sigma$, the measure has the same structure as the SLE loop measure in $\widehat{\mathbb{C}}$: it is characterised by the (universal) Shapovalov form and its partition function/reflection coefficient $\mathrm{R}_{\Sigma,p}(\lambda)$. In other words, the reflection coefficient carries all the information about the complex structure of the surface. One can then ask about the variation of this coefficient with respect to the complex structure.
More generally, one can decompose correlation functions by sampling SLEκ-type curves in a non-trivial homotopy class. Similarly to the case of the sphere, this leads to writing a correlation function as the Shapovalov inner product of several amplitudes. Applying the spectral theorem yields a conformal bootstrap formula. Again, conformal blocks are defined as (holomorphic) sections of the line bundle introduced above.
8.5. Open curves
We have studied the loop version of SLE, and it would be interesting to obtain a similar picture for open loops, e.g. Jordan curves from 0 to $\infty$. It is more difficult to find a good set of coordinates in this case, since we do not have a power series expansion of the function around a boundary point. One possible solution is to pass to the double cover and generate the deformations on the other sheet. Another approach is to observe that the open model is equivalent to the limit of the closed loop where both conformal radii are sent to 0.
8.6. Minimal models
Theorem 1.4 gives a concrete construction of the fully degenerate modules $(\mathcal{V}_{\lambda},\mathcal{Q}_{\lambda})$ for any $c_{\mathrm{m}}\leq 1$ and any conformal weight in the Kac table. If $c_{\mathrm{m}}=1-6\frac{(p-q)^{2}}{pq}$ with $p,q\geq 2$ coprime, the (analytic continuation of the) fusion product described above is expected to close on a direct sum of representations indexed by (a subset of) the Kac table. If $q=p+1$, the Shapovalov form is positive definite, so one gets an honest Hilbert space, and Theorem 1.4 gives a probabilistic construction of this Hilbert space (even though the inner-product is not an $L^{2}$-product a priori). In order to construct the minimal model in full, we need to understand the fusion rules for these representations.
8.7. Couplings with the Gaussian free field
Following Sheffield’s seminal article [She16], many couplings have been proved between SLE and the Gaussian free field (GFF). This includes a recent proof that the independent pair of an SLE loop with a quantum sphere (roughly speaking, the GFF on the sphere) can be obtained as the conformal welding of two independent quantum discs conditioned on their quantum length [AHS24].
On the other hand, this paper has shown that SLE/restriction measures are characterised by an integration by parts formula, which is of course also true of the Gaussian measure (including in infinite dimensions). This suggests an alternative approach to prove such couplings: upon cutting/welding, the integration by parts formula for one side should translate to an integration by parts formula on the other side. This should be sufficient to characterise the laws of the different objects. We will address this question in an upcoming work.
Appendix A Representation theory of the Virasoro algebra
This appendix reviews some standard facts from the representation theory of the Virasoro algebra. The main reference is [KRR13].
The Virasoro algebra is the infinite dimensional Lie algebra with generators $((L_{n})_{n\in\mathbb{Z}},\mathbf{c})$ and relations
$[L_{n},L_{m}]=(n-m)L_{n+m}+\frac{n^{3}-n}{12}\delta_{n,-m}\mathbf{c},\qquad[\mathbf{c},L_{n}]=0.$
In all representations, we assume that $\mathbf{c}$ acts by multiplication by a scalar, and we slightly abuse notation by identifying $\mathbf{c}$ with the operator $\mathbf{c}\mathrm{Id}$ where $\mathbf{c}\in\mathbb{C}$ called the central charge. [^4] In the main text, the central charge is denoted $c_{\mathrm{m}}$, but we want to distinguish between the abstract modules and the representation under study. The bold font is used to avoid any confusion with the coordinate $c$ on $\mathcal{J}_{0,\infty}$.
A highest-weight representation is a Virasoro module $V$ such that $V=\mathrm{span}\{e_{\mathbf{k}}=L_{-\mathbf{k}}\cdot e_{\eset}|\,\mathbf{k}\in\mathcal{T}\}$ for some vector $e=e_{\eset}\in V$ called the highest-weight state. If $L_{0}\cdot e=\lambda e$ for some $\lambda\in\mathbb{C}$, the representation is called of weight $\lambda$. Every highest-weight representation has a grading $V=\oplus_{N\in\mathbb{N}}V^{N}$, with $V^{N}=\mathrm{span}\{L_{-\mathbf{k}}\cdot e|\,\mathbf{k}\in\mathcal{T}_{N}\}$.
A Verma module is a highest-weight representation such that all vectors $L_{-\mathbf{k}}\cdot e$ are linearly independent. For each $\mathbf{c}\in\mathbb{C}$ and $\lambda\in\mathbb{C}$, there exists a unique Verma module of central charge $c$ and weight $\lambda$, which we denote by $M_{\mathbf{c},\lambda}$. Vectors in $M_{\mathbf{c},\lambda}^{N}$ for $N>0$ are called the descendants of $e_{\eset}$. Every highest-weight representation of central charge $\mathbf{c}$ and weight $\lambda$ is a quotient of $M_{\mathbf{c},\lambda}$ by a submodule. The Verma module $M_{\mathbf{c},\lambda}$ has a unique maximal proper (possibly trivial) submodule $I_{\mathbf{c},\lambda}$, and the quotient
$V_{\mathbf{c},\lambda}:=M_{\mathbf{c},\lambda}/I_{\mathbf{c},\lambda}$
is irreducible. It is called the irreducible quotient (of $M_{\mathbf{c},\lambda}$ by the maximal proper submodule). A vector $\chi\in M_{\mathbf{c},\lambda}$ is singular if $L_{n}\cdot\chi=0$ for all $n\geq 1$, and $\chi$ is a descendant of $e_{\eset}$. A singular vector at some level $N$ generates its own highest-weight representation (isomorphic to $M_{\mathbf{c},\lambda+N}$); in particular $M_{\mathbf{c},\lambda}$ is reducible if it has at least one singular vector. Conversely, every proper submodule is generated by singular vectors (and their descendants). In particular, $I_{\mathbf{c},\lambda}$ is the space of all singular vectors in $M_{\mathbf{c},\lambda}$ and their descendants.
The Shapovalov form is the unique Hermitian form $\langle\cdot,\cdot\rangle_{c,\lambda}$ on $M_{\mathbf{c},\lambda}$ such that $\langle e,e\rangle_{\mathbf{c},\lambda}=1$, and
$\langle L_{n}\cdot v,w\rangle_{\mathbf{c},\lambda}=\langle v,L_{-n}\cdot w\rangle_{\mathbf{c},\lambda},\qquad\forall v,w\in M_{\mathbf{c},\lambda},\,\forall n\in\mathbb{Z}.$
The coefficients of the Gram matrix are denoted $\mathrm{B}_{\lambda}(\mathbf{k},\mathbf{k}^{\prime})=\langle L_{-\mathbf{k}}\cdot e_{\eset},L_{-\mathbf{k}^{\prime}}\cdot e_{\eset}\rangle_{\mathbf{c},\lambda}$, for all $\mathbf{k},\mathbf{k}^{\prime}\in\mathcal{T}$. Equivalently, one has $L_{\mathbf{k}^{\prime}}L_{-\mathbf{k}}\cdot e_{\eset}=\mathrm{B}_{\lambda}(\mathbf{k},\mathbf{k}^{\prime})e_{\eset}$ (recall notation (1.11)). These coefficients are universal: they depend only on the central charge and the conformal weight. The maximal proper submodule $I_{\mathbf{c},\lambda}$ is the kernel of the Shapovalov form, i.e.
$I_{\mathbf{c},\lambda}=\{v\in M_{\mathbf{c},\lambda}|\,\forall w\in M_{\mathbf{c},\lambda},\,\langle v,w\rangle_{\mathbf{c},\lambda}=0\}.$
The dual Verma module $M^{\vee}_{\mathbf{c},\lambda}$ is the space of linear forms on $M_{\mathbf{c},\lambda}$ equipped with the dual representation $(L_{n}^{\vee})_{n\in\mathbb{Z}}$, i.e. [^5] In principle, the dual representation should rather be $L_{n}^{\vee}\cdot\alpha(v):=-\alpha(L_{n}\cdot v)$ but this is equivalent due to the automorphism $L_{n}\mapsto-L_{-n}$.
$L_{n}^{\vee}\cdot\alpha(v):=\alpha(L_{-n}\cdot v),\qquad\forall\alpha\in M_{\mathbf{c},\lambda}^{\vee},\,\forall v\in M_{\mathbf{c},\lambda},\,\forall n\in\mathbb{Z}.$
The Verma module and its dual are not isomorphic as Virasoro modules if $M_{\mathbf{c},\lambda}$ is reducible ($M_{\mathbf{c},\lambda}^{\vee}$ is not even a highest-weight representation in this case), but they are always isomorphic as graded vector spaces. Let $V_{\mathbf{c},\lambda}^{\vee}:=\mathrm{span}\{L_{-\mathbf{k}}^{\vee}e^{\vee}|\,\mathbf{k}\in\mathcal{T}\}$. By definition, $(V_{\mathbf{c},\lambda}^{\vee},(L_{n}^{\vee})_{n\in\mathbb{Z}})$ is a highest-weight representation of weight $\lambda$.
Lemma A.1.
We have $V_{\mathbf{c},\lambda}\simeq V_{\mathbf{c},\lambda}^{\vee}$ as highest-weight representations.
Proof.
Since $V_{\mathbf{c},\lambda}=M_{\mathbf{c},\lambda}/I_{\mathbf{c},\lambda}$, we have a linear isomorphism
$V_{\mathbf{c},\lambda}\simeq\{\alpha\in M_{\mathbf{c},\lambda}^{\vee}|\,\forall v\in I_{\mathbf{c},\lambda},\,\alpha(v)=0.\}$
In the rest of the proof, we abuse notations by writing $V_{\mathbf{c},\lambda}$ for this subspace of $M_{\mathbf{c},\lambda}^{\vee}$.
For all non-empty partitions $\mathbf{k}\in\mathcal{T}$ and all $\chi\in I_{\mathbf{c},\lambda}$, we have $L_{-\mathbf{k}}^{\vee}e^{\vee}(\chi)=e^{\vee}(L_{\mathbf{k}}\chi)=0$, where the last equality is an easy consequence of the fact that the kernel of the Shapovalov form is generated by singular vectors and their descendants. This proves that $V_{\mathbf{c},\lambda}^{\vee}\subset V_{\mathbf{c},\lambda}$.
On the other hand, $V_{\mathbf{c},\lambda}^{\vee}$ is a highest-weight representation of weight $\lambda$, we have $\dim V_{\mathbf{c},\lambda}^{N}\leq\dim V_{\mathbf{c},\lambda}^{\vee N}$ for all $N\in\mathbb{N}$. It follows that $\dim V_{\mathbf{c},\lambda}^{\vee N}=\dim V_{\mathbf{c},\lambda}^{N}$ for all $N\in\mathbb{N}$, so $V_{\mathbf{c},\lambda}$ is isomorphic to the irreducible quotient by uniqueness of the maximal proper submodule. ∎
Appendix B The SLE loop measure is Borel
The following elementary result is certainly known, but we were unable to locate a proof.
Proposition B.1.
The SLE loop measure is a Borel measure on the space of simple Jordan curves equipped with the Carathéodory topology.
To prove this proposition, we will first start by proving a similar statement for chordal SLE. For the reader’s convenience, we start by a short reminder on this theory. We refer to [RS05] for a more complete introduction, and to [Pom75, Chapter 6] for some background on the original analytic (deterministic) theory due to Loewner.
A Jordan slit is an (unparametrised) simple Jordan arc in $\mathbb{H}$ with an endpoint at $0\in\partial\mathbb{H}$. We denote by $\mathcal{K}$ the space of Jordan slits, equipped with the Carathéodory topology. Let also $\overline{\mathcal{K}}$ be the space of all simple Jordan arcs with an endpoint at 0 and an endpoint in ${}\mathbb{H}\cup\{\infty\}$. This space is also equipped with the Carathéodory topology, and $\mathcal{K}$ is an open subset of $\overline{\mathcal{K}}$.
Each Jordan slit has a unique parametrisation $\eta$ on some interval $[0,T]$ such that the following holds. For all $t\in[0,T]$, let $f_{t}:\mathbb{H}\setminus\eta[0,t]\to\mathbb{H}$ be the conformal map with the hydrodynamical normalisation (i.e. $f_{t}(z)=z+O(1/z)$ as $z\to\infty$). Then, $f_{t}(z)=z+\frac{2t}{z}+o(1/z)$ as $z\to\infty$. One says that $\eta$ is parametrised by half-plane capacity. In this case, $f_{t}$ satisfies the differential equation $\partial_{t}f_{t}=\frac{2}{f_{t}-W_{t}}$, where $W\in\mathcal{C}^{0}([0,T];\mathbb{R})$ is called the driving function. Conversely, every driving function $W\in\mathcal{C}^{0}([0,T];\mathbb{R})$ generates a family of compact hulls in the upper-half plane (but not necessarily a curve). We denote by $\mathfrak{L}:\mathcal{C}^{0}(\mathbb{R}_{+})\to\overline{\mathcal{K}}$ the Loewner map sending driving functions to Jordan slits.
Chordal SLEκ (run until capacity $T>0$) is the probability measure on $\mathcal{K}$ where the driving function is a Brownian motion in $[0,T]$ with speed $\kappa$. We denote by $\mathrm{SLE}^{T}_{\kappa}$ this measure, so $\mathrm{SLE}_{\kappa}^{1}=\mathfrak{L}_{*}\mathbb{P}$ where $\mathbb{P}$ is the law of $\sqrt{\kappa}$ times standard real Brownian motion on $[0,1]$, a Borel probability measure on $\mathcal{C}^{0}([0,1])$. It is well-known that for $\kappa\in(0,4]$, the solution to the Loewner equation is almost surely a simple Jordan slit.
Proposition B.2.
For all $T>0$, $\mathrm{SLE}_{\kappa}^{T}$ is a Borel measure on $\mathcal{K}$. Moreover, $\mathrm{SLE}_{\kappa}^{T}$ converges weakly to a (Borel) probability measure on $\overline{\mathcal{K}}$ giving full mass to Jordan curves from 0 to $\infty$ in $\mathbb{H}$.
Proof.
We start by showing that $\mathrm{SLE}^{1}_{\kappa}$ is Borel.
For each $z\in\mathbb{H}$ and $r>0$, define the open set $E_{z,r}:=\{\eta\text{ visits }D(z,r)\}\subset\mathcal{K}$. This event is $\mathrm{SLE}_{\kappa}^{1}$-measurable since we can write
$E_{z,r}=\{\exists q\in\mathbb{Q}^{2}\cap D(z,r)|\,\mathrm{Im}(f_{1}(q))=0\},$
and $\{\mathrm{Im}(f_{1}(q))=0\}$ is measurable with respect to the Brownian motion driving the Loewner equation.
Now, let $\eta_{0},\eta\in\mathcal{K}$ and $r>0$. From the previous paragraph, for all $z\in\eta$, the event $\{d(z,\eta_{0})<r\}$ is measurable (with $d$ the usual distance). Hence, the event $\{\sup_{z\in\eta}\,d(z,\eta_{0})<r\}$ is also measurable. The same is true when exchanging the roles of $\eta_{0}$ and $\eta$, so by definition of the Hausdorff distance $d_{\mathcal{H}}$, the event $\{d_{\mathcal{H}}(\eta_{0},\eta)<r\}$ is measurable. This shows that metric balls in $\mathcal{K}$ are measurable, so $\mathrm{SLE}_{\kappa}^{1}$ is Borel.
By scaling, we easily deduce that $\mathrm{SLE}_{\kappa}^{T}$ is a Borel measure on $\mathcal{K}$ (or $\overline{\mathcal{K}}$) for all $T>0$. Chordal SLEκ from 0 to $\infty$ is then Borel on $\overline{\mathcal{K}}$ as a weak limit of Borel measures. ∎
Proof of Proposition B.1.
[…]